### v1

# Lucas Resende CREST, ENSAE, Institut Polytechnique de Paris, France Guillaume Lecué ESSEC Business School, France Lionel Wilner CREST, ENSAE, Institut Polytechnique de Paris, France Philippe Choné CREST, ENSAE, Institut Polytechnique de Paris, France

Abstract
Maximum likelihood estimators, such as the Poisson Pseudo-Maximum Likelihood (PPML), suffer from the incidental parameter problem: a bias in the estimation of structural parameters that arises from the joint estimation of structural and nuisance parameters. To address this issue in multi-way gravity models, we propose a novel, asymptotically unbiased estimator. Our method reframes the estimation as a series of classification tasks and is agnostic to both the number and structure of fixed effects. In sparse data environments — common in the network formation literature — it is also computationally faster than PPML. We provide empirical evidence that our estimator yields more accurate point estimates and confidence intervals than PPML and its bias-correction strategies. These improvements hold even under model misspecification and are more pronounced in sparse settings. While PPML remains competitive in dense, low-dimensional data, our approach offers a robust alternative for multi-way models that scales efficiently with sparsity. The method is applied to estimate the effect of a policy reform on spatial accessibility to health care in France.
Keywords weighted networks $\cdot$ incidental parameter problem $\cdot$ gravity models $\cdot$ sparse data $\cdot$ polyad estimator
JEL codes C13 C31 C33 C55
1 Introduction
Gravity-type models are extensively used in social sciences to study flows or interactions between entities of different nature. Economists have relied on gravity equations to study bilateral trade flows at least since Timb_1962, see the survey of head2014gravity. Gravity models are relevant in many other fields than trade economics. For instance, they have been applied in labor and development economics to study migration flows between regions or countries, beine2016practitioners; in international finance to model foreign direct investments, see di2005drives; in innovation economics to measure knowledge flows (citations, patents, etc.), peri2005determinants; in health economics to measure interregional patient flows, dingel2023market, as well as spatial accessibility to care, luo2003measures.
Since the important contribution of silva2006log, it is known that the traditional log-linear OLS estimation suffers from bias under heteroskedasticity, particularly when many flows are zero. [^1] Recently, chen2024logs argue that log-like transformations can also distort the interpretation of coefficients as percentage effects, since they depend on the units of the outcome. This observation greatly motivated the adoption of Poisson models for gravity. The properties of Poisson pseudo-maximum likelihood estimators have first been investigated by gourieroux1984pseudo in the absence of fixed effects, with one-way fixed effect panel data applications being pioneered by HHG. To uncover fundamental parameters at disaggregated levels and to control for unobserved heterogeneity along various dimensions, recent applied research has gradually considered models with more levels of fixed effects, introducing higher-dimensional interactions in panel or multi-way gravity models. [^2] Three-way gravity models, with exporter-year, importer-year, and exporter-importer fixed effects, are common in the modern trade literature. Recent studies recognize the economic importance of the sector or product levels, potentially leading to even richer structures of fixed effects, breinlich2024trade and Delb_Dina_WD_2020. As the structure of fixed effects becomes more complex, fernandez2016individual and weidner2021bias show that direct MLE under the Poisson model may be plagued by incidental parameter problems. Specifically, as the sample size grows, so too does the number of nuisance parameters representing the fixed effects, possibly creating bias in standard maximum likelihood estimation of the parameters of interest, as first described by Neym_Scot_Ecma_1948.
In this paper, we tackle the incidental parameter problem in multi-way gravity models by proposing a novel, asymptotically unbiased estimator. In the spirit of graham2017econometric, we regard data through the lens of graph theory. The key insight is that certain configurations of count outcomes within subgroups of observations, which we call polyads, have identical sufficient statistics for the fixed effects, making their relative likelihoods independent of the fixed effects. Our framework differs from graham2017econometric in two important dimensions. First, while Graham models undirected graphs, we consider multipartite graphs. Second, Graham’s network formation model considers only the extensive margin, i.e., the probability that potential links are realized. By contrast, we also consider the intensive margin, i.e., the strength of connections.
Our method compares favorably with existing approaches for gravity models along several dimensions. First, it accommodates multi-way models with an arbitrary number of node groups, in contrast to debiasing methods such as fernandez2016individual and weidner2021bias, which are restricted to two- and three-way structures, and to two-way approaches like graham2017econometric, charbonneau2012multiple, and jochmans2017two. Second, unlike PPML estimation, our estimator is bias-free by construction. Even relative to bias-corrected PPML procedures, our approach remains advantageous: the corrections may themselves be biased in finite samples, as weidner2021bias argue, or computationally prohibitive in large datasets, as in zylkin2024bootstrap. Third, compared with approaches such as charbonneau2012multiple, our estimator better exploits the available variability while remaining computationally feasible, and the convexity of our loss function delivers strong numerical performance relative to more general semiparametric method-of-moments procedures (e.g., jochmans2017two; yang2023three).
Our study is also connected to the literature on discrete choice models for panel and network data (Rasc_1960, Ande_1973,Cham_Hand_1984, Hono_Kyri_Ecma_2000, Magn_Ecma_2004). Presenting the conditional likelihood methods used in these settings, the recent review of Dano_Hono_Weid_WP_2025 highlights how identification strategies relate to difference-in-differences approaches. Specifically, they provide the differencing vectors that are valid to identify the parameters of interest. [^3] Muris_Pakel_WP_2025 follow this approach to study the formation of triadic networks. They introduce a hexad logit estimator that extends the tetrad logit estimator of graham2017econometric. We proceed the same way for count data and gravity models. The polyad estimator can be thought of as a nonlinear version of a difference-in-differences estimator. Contrary to the above cited literature, the polyad method handles count data and recovers both the existence and intensity of relationships. In the special case where the count data is right-censored at one, i.e., where the intensive margin is ignored, the polyad estimator maximizes the same conditional likelihood as graham2017econometric and the studies mentioned in this paragraph .
For researchers working with sparse networks — i.e., where most potential connections are not realized —, our approach offers a distinct computational advantage: polyads can be constructed by looping over pairs of edges with strictly positive counts (i.e., realized connections), allowing the estimation procedure to scale with the number of observed relationships rather than the number of potential relationships. This is a significant step forward, since the usage of tetrad-based methods has been limited by its computational cost: the available methods for computing tetrad-based statistics, which work only for two- or three-way models, either (i) require looping over all pairs of edges, including unrealized connections (e.g., graham2017econometric; Muris_Pakel_WP_2025), or (ii) rely on matrix multiplications that remain slow for sparse networks (e.g., jochmans2017two). In particular, as exemplified by our experiments, our computational implementation enables the use of the polyads method on large administrative datasets. See Section 5.5 for a more detailed discussion.
We establish consistency and asymptotic normality under mild assumptions, extending the current theoretical framework in two main directions: relaxing sampling assumptions and better exploiting convexity. First, we do not require the existence of a distribution to sample the nodes, thus treating the fixed effects as true parameters and revealing fundamental geometrical properties that a graph should satisfy for consistency and asymptotic normality (see Assumptions 2 and 4). Second, we obtain consistency without boundedness assumptions by modifying classical results from newey1994large under the light of convex analysis tools from rockafellar-1970a and asymptotic statistics results from Andersen1982. We also avoid the existence of a limiting risk for asymptotic normality exploring results from victor, which are based on MR1026303; MR1186263.
The practical limitations of our method merit clear statement. First, we do not consider interdependencies between observations beyond those captured by fixed effects and observed covariates. Second, our approach is designed specifically for count data, rather than continuous weights. Third, the method becomes computationally inefficient in dense networks where most potential relationships are realized. Within these constraints, however, our estimator provides a powerful tool for estimating gravity models with high-dimensional fixed effects structures.
The remainder of the paper proceeds as follows. Section 2 introduces the Poisson model and the structure of fixed effects. Section 3 introduces the polyad estimator. Section 4 establishes its theoretical properties, demonstrating consistency and asymptotic normality. Section 5 develops the computational implementation, emphasizing how the algorithm efficiently exploits sparsity. Section 6 documents the finite-sample properties of the estimator using artificial data and healthcare claims data.
2 Model assumptions
We consider a count variable $Y_{i_{1}i_{2}\cdots i_{D}}\in\mathbb{N}$ that is indexed by $(i_{1},\dots,i_{D})\in\mathcal{I}=[n_{1}]\times\cdots\times[n_{D}]$. Using the $D$-dimensional index ${\bf i}=(i_{1},\dots,i_{D})\in\mathcal{I}$, we represent the variable as $Y_{\bf i}$.
Throughout the paper, we think of $Y=(Y_{\bf i})_{\bf i}\in\mathbb{N}^{\mathcal{I}}$ as a random $D$-partite graph, with the sets $n_{1}],[n_{2}],\dots[n_{D}]$ representing the nodes of each category, the multidimensional index $\bf i$} representing a potential (hyper-)edge of a the graph, and $Y_{{\bf i}}$ being the number of connections along edge $\bf i$}, see the concrete examples below. We denote by $E$ the set of positive edges, i.e., the set of $D$-dimensional indices $\bf i$} such that $Y_{{\bf i}}>0$. The graph is sparse when the data contains many zeros, a case where our method delivers especially good results.
We assume that the dependent variable $Y_{{\bf i}}$ depends on a set of $p$ explanatory variables $X_{\bf i}\in\mathbb{R}^{p}$ and a set of fixed effects. A level of fixed effect is represented by a proper subset $g$ of $[D]$. By abuse of notation, we set $g({\bf i})=(i_{d}:d\in g)$ and the fixed effects for level $g$ are denoted as $\theta_{g({\bf i})}^{g}\in\mathbb{R}$. The structure of the fixed effects in the model is represented by a collection $\mathcal{G}$ of fixed effects levels. Of particular interest to us is the structure $\mathcal{G}^{{\rm max}}$ consisting of the $D$ subsets of $[D]$ of cardinal D1D-1; in this particular case, each level of fixed effect absorbs the variations of $Y_{\bf i}$ in all but one dimension of $\bf i$}. The set of all fixed effects is denoted by $\theta^{\mathcal{G}}=\{\theta^{g}_{\rho}:g\in\mathcal{G},\rho\in g({\bf i})\}$.
Assumption 1.
Let $\beta_{\star}\in\mathbb{R}^{p}$ be the parameter of interest. The distribution of $Y=(Y_{\bf i}\in\mathbb{N}:{\bf i}\in\mathcal{I})$ conditionally on $X=(X_{\bf i}\in\mathbb{N}:{\bf i}\in\mathcal{I})$ is $\mathbb{P}_{\beta_{\star},\theta^{\mathcal{G}}}^{Y|X}=\bigotimes_{{\bf i}\in\mathcal{I}}\mathcal{P}(\lambda_{\bf i})$ , where $\mathcal{P}(\lambda_{\bf i})$ is the Poisson distribution with intensity $\lambda_{\bf i}>0$ given by
$\ln\lambda_{\bf i}=\beta_{\star}^{\top}X_{\bf i}+\sum_{g\in\mathcal{G}}\theta_{g({\bf i})}^{g}.$ (1)
It follows that the residual $\varepsilon_{{\bf i}}=Y_{\bf i}-\lambda_{\bf i}$ satisfies $\mathbb{E}(\varepsilon_{\bf i}|X)=\mathbb{E}(\varepsilon_{\bf i}|X_{\bf i})=0$. In other words, the explanatory variables $X_{\bf i}$ are assumed to be strongly exogenous.
The log-likelihood of $(\beta,\theta^{\cal G})$ in the model (1) at the observed graph $y$ is
$\ln\mathbb{P}_{\beta,\theta^{\mathcal{G}}}\left(Y=y|X\right)=\sum_{{\bf i}\in\mathcal{I}}-\lambda_{\bf i}-\ln y_{\bf i}!+y_{\bf i}\left\{\beta^{\top}X_{\bf i}+\sum_{g\in\mathcal{G}}\theta_{g({\bf i})}^{g}\right\}$ (2)
involves a potentially high number of fixed effects. The MLE estimator of the parameter of interest $\beta_{\star}$ has been shown to be asymptotically biased for $D\geq 3$, see weidner2021bias as well as the experiments presented in Section 6.
The following examples show how our framework encompasses gravity models and other classical econometric models.
Example 1 (One-way model in panel data).
Taking $D=2$ and $\mathcal{G}=\{\{1\}\}$ yields the structure of the classical model studied by HHG
$\ln\lambda_{i_{1}i_{2}}=\beta_{\star}^{\top}X_{i_{1}i_{2}}+\theta_{i_{1}}^{1}$
Example 2 (Two-way model).
Taking $D=2$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1\},\{2\}\}$ yields the structure of the standard gravity model studied by silva2006log , i.e., $\ln\lambda_{i_{1}i_{2}}=\beta_{\star}^{\top}X_{i_{1}i_{2}}+\theta_{i_{1}}^{1}+\theta_{i_{2}}^{2}$. The usual econometric model
$\ln\lambda_{ij}=\beta_{\star}^{\top}X_{ij}+u_{i}+v_{j}$
obtains when relabeling the two-dimensional indices $(i_{1},i_{2})$ as $(i,j)$ and the fixed effects $\theta_{i_{1}}^{1}$ and $\theta_{i_{2}}^{2}$ as $u_{i}$ and $v_{j}$ respectively. In the trade literature, $i$ is an exporter, $j$ is an importer, $X_{i_{1}i_{2}}$ is a feature of the dyad (e.g., sharing borders or same language, having a free trade agreements in force).
Example 3 (Three-way model).
Taking $D=3$ and $\mathcal{G}=\mathcal{G}^{{\rm max}}=\{\{1,2\},\{1,3\},\{2,3\}\}$ yields the structure of the model studied by weidner2021bias, i.e., $\ln\lambda_{i_{1}i_{2}i_{3}}=\beta_{\star}^{\top}X_{i_{1}i_{2}i_{3}}+\theta_{i_{1}i_{2}}^{1,2}+\theta_{i_{1}i_{3}}^{1,3}+\theta_{i_{2}i_{3}}^{2,3}$. Relabeling the edges $(i_{1},i_{2},i_{3})$ as $(i,j,t)$ and the fixed effects $(\theta_{i_{1}i_{2}}^{1,2},\theta_{i_{1}i_{3}}^{1,3},\theta_{i_{2}i_{3}}^{2,3})$ as $(u_{ij},v_{jt},w_{it})$, we obtain the usual econometric formulation
$\ln\lambda_{ijt}=\beta_{\star}^{\top}X_{ijt}+u_{ij}+v_{jt}+w_{it}.$
This model is used in the trade literature in the presence of a time dimension, where $i_{1}$ is an exporter, $i_{2}$ is an importer, and $i_{3}$ is the time. Denoting the indices $i,j,t$ instead of $i_{1},i_{2},i_{3}$ yields a notation familiar to economists $\ln\lambda_{ijt}=\beta_{\star}^{\top}X_{ijt}+u_{ij}+v_{jt}+w_{it}$.
3 Estimation of the homophily parameter $\beta_{\star}$ via polyads
To avoid the incidental parameter problem mentioned above, we first condition the likelihood on a sufficient statistics for the fixed effects given by the degrees of the nodes, see Subsection 3.1. This approach has been followed in two-way contexts by charbonneau2012multiple, graham2017econometric and jochmans2018semiparametric. We construct a loss function based on a set of ‘directions’ that generate all the variability in the data for given degrees.
3.1 Generalized degrees and polyad transformations
Consider a level of fixed effect $g\in\mathcal{G}$. For $\rho\in g(\mathcal{I})$, the fixed effect $\theta_{\rho}^{g}$ enters the log-likelihood (2) only through the quantity
$\delta^{g}_{\rho}(y)=\sum_{{\bf i}\in\mathcal{I}:g({\bf i})=\rho}y_{\bf i},$
which we call the degree of $\rho$ relative to the fixed effect level $g$ in the graph $y$. This quantity generalizes the notion of degree used in graham2017econometric, capturing the total number of connections between edges $\bf i$} for which $g({\bf i})=\rho$. [^4] Consider for instance Example 3, where the index $i_{1},i_{2},i_{3}$ are denoted $i,j,t$ as in many gravity models. The degree of (4,5) relative to the fixed effect level $u_{ij}$ is $\delta^{1,2}_{4,5}(y)=\sum_{t}y_{45t}$. The set of degrees for the fixed effects level $g\in\mathcal{G}$ is denoted by $\delta^{g}(y)=\left[\delta^{g}_{\rho}(y)\right]_{\rho\in g(\mathcal{I})}$. The set of all degrees is denoted by $\delta(y)=\left[\delta^{g}(y)\right]_{g\in\mathcal{G}}$.
As announced above, we condition the likelyhood on the set of degrees $\delta(Y)$:
$\mathbb{P}_{\beta,\theta^{\mathcal{G}}}\left(Y=y\,|\,X,\delta(Y)=\delta(y)\right)=\frac{\displaystyle\prod_{\bf i}e^{-\lambda_{\bf i}}\lambda_{{\bf i}}^{y_{\bf i}}/y_{\bf i}!}{\displaystyle\sum_{z|\delta(z)=\delta(y)}\prod_{\bf i}e^{-\lambda_{\bf i}}\lambda_{{\bf i}}^{z_{\bf i}}/z_{\bf i}!}.$ (3)
Rewriting the conditional likelihood as
$\mathbb{P}_{\beta,\theta^{\mathcal{G}}}\left(Y=y\,|\,X,\delta(Y)=\delta(y)\right)$ $=$ $\frac{\exp\left\{\sum_{\bf i}y_{\bf i}\big(\sum_{g\in\mathcal{G}}\theta^{g}_{g({\bf i})}\big)+y_{\bf i}\beta^{\top}X_{\bf i}-\ln y_{\bf i}!\right\}}{\sum_{z:\delta(z)=\delta(y)}\exp\left\{\sum_{\bf i}z_{\bf i}\big(\sum_{g\in\mathcal{G}}\theta^{g}_{g({\bf i})}\big)+z_{\bf i}\beta^{\top}X_{\bf i}-\ln z_{\bf i}!\right\}}$ (4)
$=$ $\frac{\exp\left\{\sum_{g\in\mathcal{G},\rho\in g(\mathcal{I})}\delta^{g}_{\rho}(y)\theta^{g}_{\rho}+\sum_{\bf i}\big(y_{\bf i}\beta^{\top}X_{\bf i}-\ln y_{\bf i}!\big)\right\}}{\sum_{z:\delta(z)=\delta(y)}\exp\left\{\sum_{g\in\mathcal{G},\rho\in g(\mathcal{I})}\delta^{g}_{\rho}(z)\theta^{g}_{\rho}+\sum_{\bf i}\big(z_{\bf i}\beta^{\top}X_{\bf i}-\ln z_{\bf i}!\big)\right\}}$
$=$ $\frac{\exp\left\{\sum_{\bf i}y_{\bf i}\beta^{\top}X_{\bf i}-\ln y_{\bf i}!\right\}}{\sum_{z:\delta(z)=\delta(y)}\exp\left\{\sum_{\bf i}z_{\bf i}\beta^{\top}X_{\bf i}-\ln z_{\bf i}!\right\}}$
shows that it does not depend on the fixed effects $\theta^{\mathcal{G}}$ regardless of their structure $\mathcal{G}$. In the next subsection, we characterize the support of the distribution of the $Y$ conditional on all degrees $\delta(Y)$.
Definition 1.
A polyad $\xi$ of ${\cal I}=[n_{1}]\times\cdots\times[n_{D}]$ is a $2\times D$ matrix
$\xi=\begin{pmatrix}j_{1}&j_{2}&\cdots&j_{D}\\ j_{1}^{\prime}&j_{2}^{\prime}&\cdots&j_{D}^{\prime}\\ \end{pmatrix}$
[…]
An edge of a polyad $\xi$ is an index ${\bf i}=(i_{d})_{d}\in{\cal I}$ such that $i_{d}\in\{j_{d},j_{d}^{\prime}\}$ for all $d\in[D]$. We denote by ${\cal E}(\xi)$ the set of all edges of $\xi$. Any polyad has $|{\cal E}|=2^{D}$ edges. In other words, a polyad $\xi$ induces a subgraph of $Y$ made of $2^{D}$ edges with weights $(y_{{\bf i}})_{{\bf i}\in{\cal E}(\xi)}$.
Polyads as introduced in Definition 1 are generalizations to the $D$-dimensional framework of tetrads from charbonneau2012multiple; graham2017econometric; jochmans2018semiparametric defined for $D=2$. The total number of polyads, i.e. the size of $\Xi$, is $\prod_{d=1}^{D}n_{d}(n_{d}-1)$. In the upper left corner of Figure 1 we exemplify a polyad on $D=2$,
Definition 2.
Let $\xi\in\Xi$ be a polyad. Let $\bf i$$\in{\cal I}$. The sign of $\bf i$} relative to $\xi$ is defined as
$s_{\xi}({\bf i})=\prod_{d=1}^{D}\left(\mathbf{1}\{i_{d}=j_{d}\}-\mathbf{1}\{i_{d}=j_{d}^{\prime}\}\right).$
In particular, $s_{\xi}({\bf i})=0$ when $\bf i$} is not an edge of $\xi$ and $s_{\xi}({\bf i})\in\{\pm 1\}$ otherwise. The main purpose of the sign function $s_{\xi}$ is that it gives the signs of the diff-in-diff property stated below. The sign function is represented in the upper right corner of Figure 1, each edge has a sign associated to it, notice that the signs sum zero for all axis and that the sign of $\bf j$} is always 1.
We now define a class of transformations indexed by polyads. These transformations act on $D$-partite graphs with integer edge weights, i.e., on the set $\mathbb{Z}^{\mathcal{I}}$. Recall that we see the dependent variable $Y\in\mathbb{N}^{\mathcal{I}}$ as a graph with nonnegative edge weights. This discrepancy plays an important role in the analysis developed below.
Definition 3.
Let $\xi\in\Xi$ be a polyad. The polyad transformation $T_{\xi}:\mathbb{Z}^{\cal I}\rightarrow\mathbb{Z}^{\cal I}$ is defined by $T_{\xi}(y)=y+s_{\xi}$ where $s_{\xi}=(s_{\xi}({\bf i}):{\bf i}\in{\cal I})$ i.e. $T_{\xi}(y)$ is a graph with weights given for all $\bf i$$\in{\cal I}$ by
$T_{\xi}(y)_{\bf i}=y_{\bf i}+s_{\xi}({\bf i}).$ (5)
For all $r\in\mathbb{Z}$, $T_{\xi}^{r}(y)=y+rs_{\xi}$.
For any polyad $\xi$, the transformation $T_{\xi}$ alters only the subgraph of $y$ induced by the polyad $\xi$. In other words, $y^{\prime}_{\bf i}=y_{\bf i}$ for all ${\bf i}\notin{\cal E}(\xi)$. The proposition below states the polyad transformations preserve degrees and, conversely, that they allow to generate all graphs sharing the same degrees as a given graph. [^5] The converse result is stated in Proposition 1 only for $\mathcal{G}=\mathcal{G}^{{\rm max}}$. In the appendix, we consider any fixed effect structure $\mathcal{G}$.
Proposition 1 (Characterization of degree-preserving transformations).
Consider any graph $y\in\mathbb{Z}^{\cal I}$ and any polyad $\xi\in\Xi$. If $y^{\prime}=T_{\xi}(y)$, then we have
$\delta(y^{\prime})=\delta(y).$ (6)
Conversely, take two graphs $y$ and $y^{\prime}$ in $\mathbb{Z}^{\cal I}$ having the same degrees, i.e., such that $\delta(y^{\prime})=\delta(y)$. Suppose furthermore that the structure of fixed effect is $\mathcal{G}^{{\rm max}}$. Then there exists a finite sequence of integers $r_{1},\dots,r_{m}\in\mathbb{Z}$ and finite sequence of polyads such that
$y^{\prime}=T_{\xi_{1}}^{r_{1}}\circ\cdots\circ T_{\xi_{m}}^{r_{m}}(y).$ (7)
Proof.
Pick a polyad $\xi$, a fixed effect level $g\in\mathcal{G}$, $\rho\in g(\mathcal{I})$, and $y^{\prime}=T_{\xi}(y)$. We observe that the equality
$\sum_{{\bf i}\in\mathcal{I}:g({\bf i})=\rho}s_{\xi}({\bf i})=0$ (8)
immediately implies
$\delta^{g}_{\rho}(y^{\prime})=\sum_{{\bf i}\in\mathcal{I}:g({\bf i})=\rho}y^{\prime}_{{\bf i}}=\sum_{{\bf i}\in\mathcal{I}:g({\bf i})=\rho}y_{{\bf i}}+s_{\xi}({\bf i})=\sum_{{\bf i}\in\mathcal{I}:g({\bf i})=\rho}y_{{\bf i}}=\delta^{g}_{\rho}(y),$
and hence the direct part of the proposition. To prove (8), we consider an edge $\bf i$} such that $g({\bf i})=\rho$. If $\bf i$} is not an edge of the polyad-induced subgraph, i.e., if ${\bf i}\notin\mathcal{E}(\xi)$, the sign $s_{\xi}({\bf i})=0$. Consider now the edges that belong to $\mathcal{E}(\xi)$. There exist $2^{D-|g|}\geq 2$ edges ${\bf i}\in\mathcal{E}(\xi)$ such that $g({\bf i})=\rho$, with half of them having $s_{\xi}({\bf i})=1$ and the other half having $s_{\xi}({\bf i})=-1$, which yields (8). The converse part is proved in the Appendix C.1. ∎
According to Proposition 1, the conditioning set in the likelihood (3) can be written as
$\left\{\,Y\,|\,\delta(Y)=\delta(y)\,\right\}=\left\{\,Y=T_{\xi_{1}}^{r_{1}}\circ\cdots\circ T_{\xi_{m}}^{r_{m}}(y):m\geq 0,(\xi_{1},\dots,\xi_{m})\in\Xi^{m},(r_{1},\dots,r_{m})\in\mathbb{Z}^{m}\,\right\},$
which can be represented only at a prohibitively high computational cost.
3.2 A classification problem on polyads and the associated estimator of $\beta_{\star}$
Rather than attempting to exploit the data variations within this whole set, we propose to exploit variations in the directions induced by each polyad separately, i.e., to restrict attention to sequences of polyads of length $m=1$. Because the count variable $Y_{\bf i}$ takes nonnegative values, the transformations of the graph $T_{\xi}^{r}(Y)$ that do not belong to $\mathbb{N}^{\mathcal{I}}$ are irrelevant because they occur with zero probability. We thus further restrict the conditioning set. For any polyad $\xi\in\Xi$, we introduce the nonnegative integers $m_{\xi}(y)$ and $M_{\xi}(y)$ given by
$m_{\xi}(y)=\bigwedge_{{\bf i}:s_{\xi}({\bf i})=1}y_{\bf i}\ \ \ \text{ and }\ \ \ M_{\xi}(y)=\bigwedge_{{\bf i}:s_{\xi}({\bf i})=-1}y_{\bf i}.$
The range of integers $r$ such that $T_{\xi}^{r}(y)\in\mathbb{N}^{\mathcal{I}}$ is $\{-m_{\xi}(y),\dots,M_{\xi}(y)\}$. Accordingly, we define the orbit $\mathcal{O}_{\xi}(y)$ of a graph $y\in\mathbb{N}^{\mathcal{I}}$ with respect to $\xi$ as
$\mathcal{O}_{\xi}(y):=\left\{\,T_{\xi}^{r}(y):-m_{\xi}(y)\leq r\leq M_{\xi}(y)\,\right\}.$ (9)
All graphs $y$ in the orbit $\mathcal{O}_{\xi}(Y)$ of the observed graph have the same degrees as $Y$, $\delta(y)=\delta(Y)$, and coincide with $Y$ except in the subgraph induced by $\xi$, $Y_{\bf i}=y_{\bf i}$ for ${\bf i}\notin\mathcal{E}(\xi)$.
Figure 1: A numerical example illustrating the definitions introduced in this section and their interrelations is provided under the two-way setting (Example 2). The figure depicts a polyad $\xi$, the sign $s_{\xi}({\bf i})$ of an edge $\bf i$} with respect to the polyad, the quantities $m_{\xi}(y)$ and $M_{\xi}(y)$, as well as the classes $\mathcal{O}_{\xi}(y)$. Notice that taking $r$ below $m_{\xi}(y)$ or over $M_{\xi}(y)$ yields negative — and therefore invalid — edges.
We introduce the loss function at the level of the polyad $\xi$:
$\ell_{\xi}(y|X,\beta)=-\ln\mathbb{P}_{\beta}\left(\,Y=y\,|\,X,Y\in\mathcal{O}_{\xi}(y)\,\right).$ (10)
A polyad $\xi$ is not informative if the orbit $\mathcal{O}_{\xi}(Y)$ is a singleton, i.e., if both $m_{\xi}(Y)$ and $M_{\xi}(Y)$ are zero. For non-informative polyads, we have $\ell_{\xi}(y|X,\beta)=0$ for all $\beta\in\mathbb{R}^{p}$. We can thus restrict attention to informative (or “active”) polyads $\xi$ for which the corresponding orbit $\mathcal{O}_{\xi}(Y)$ has room for potential variation in the data, i.e., contains at least two distinct elements. Formally, active polyads satisfy: $|\mathcal{O}_{\xi}(Y)|=m_{\xi}(Y)+M_{\xi}(Y)+1\geq 2$. To simplify notations, we denote the transformed graph $T^{r}_{\xi}(y)$ by $y^{r}$, hence $y^{r}_{\bf i}=y_{\bf i}+rs_{\xi}({\bf i})$. The same computation as in (4) yields
$\mathbb{P}_{\beta}\left(\,Y=y\,|\,X,Y\in\mathcal{O}_{\xi}(y)\,\right)$ $=$ $\frac{\displaystyle\exp\left\{\sum_{\bf i}y_{\bf i}\beta^{\top}X_{\bf i}-\ln y_{\bf i}!\right\}}{\displaystyle\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}\exp\left\{\sum_{\bf i}y_{\bf i}^{r}\beta^{\top}X_{\bf i}-\ln y_{\bf i}^{r}!\right\}}$
$=$ $\left[\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}\exp\left\{\sum_{\bf i}rs_{\xi}({\bf i})\beta^{\top}X_{\bf i}+\ln\frac{y_{\bf i}!}{y_{\bf i}^{r}!}\right\}\right]^{-1}.$
As already explained, all graphs in the orbit $\mathcal{O}_{\xi}(Y)$ coincide with $Y$ except in the subgraph induced by the polyad $\xi$. Formally, for ${\bf i}\notin\mathcal{E}(\xi)$, we have $s_{\xi}({\bf i})=0$ and hence $y_{\bf i}^{r}=y_{\bf i}$ for all $r$ between $m_{\xi}$ and $M_{\xi}$. In the above sums over edges $\bf i$}, we can thus restrict attention to edges ${\bf i}\in\mathcal{E}(\xi)$. From (10), the loss function associated with the polyad $\xi$ is
$\ell_{\xi}(y|X,\beta)=\ln\left(\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}\exp\left\{r\beta^{\top}\widetilde{X}_{\xi}+\sum_{{\bf i}\in\mathcal{E}(\xi)}\ln\frac{y_{\bf i}!}{y^{r}_{\bf i}!}\right\}\right),$ (11)
where the generalized “difference-in-differences” (DiD) operator $\widetilde{X}_{\xi}$ is defined as
$\widetilde{X}_{\xi}=\sum_{{\bf i}\in\mathcal{I}}s_{\xi}({\bf i})X_{\bf i}=\sum_{{\bf i}\in\mathcal{E}(\xi)}s_{\xi}({\bf i})X_{\bf i}.$ (12)
In Example 2, consider the polyad
$\xi=\begin{pmatrix}1&1\\ 2&2\\ \end{pmatrix}$.
In this two-way example, we get the usual DiD formula:
$\widetilde{X}_{\xi}=X_{22}-X_{21}-(X_{12}-X_{11})=X_{22}-X_{21}-X_{12}+X_{11}.$
In Example 3, consider the polyad
$\xi=\begin{pmatrix}1&1&1\\ 2&2&2\\ \end{pmatrix}$.
In this three-way example, we get (the opposite of) a “triple difference” formula:
$-\widetilde{X}_{\xi}$ $=$ $[X_{222}-X_{221}-(X_{212}-X_{211})]-[X_{122}-X_{121}-(X_{112}-X_{111})]$
$=$ $X_{222}-X_{221}-X_{212}+X_{211}-X_{122}+X_{121}+X_{112}-X_{111}.$
Remark 1 (Difference-in-differences (DiD)).
For a given polyad $\xi$, we may look at the sign $s_{\xi}=(s_{\xi}({\bf i}))_{\bf i}$ as a DiD operator acting on the observed graph $y$ and the observed feature vector $(X_{\bf i})_{\bf i}$ either in an additive or a multiplicative way leading to the polyad transformation $T_{\xi}(y)=y+s_{\xi}$ and the polyad feature $\widetilde{X}_{\xi}=\bigl<s_{\xi},(X_{\bf i})_{\bf i}\bigr$>. The tensor $s_{\xi}=(s_{\xi}({\bf i}))_{\bf i}$ provides the sign in a DiD approach.
Finally, we call $\Xi_{a}$ the set of informative (or active) polyads and form the loss function taking $y=Y$:
$\widehat{L}_{\Xi}(Y|X,\beta)=\sum_{\xi\in\Xi}\ell_{\xi}(Y|X,\beta)=\sum_{\xi\in\Xi_{a}}\ell_{\xi}(Y|X,\beta).$ (13)
Two immediate observations will play an important role in our analysis. First, because the LogSumExp function is convex, [^6] Recall that the LogSumExp function $\mathbb{R}_{+*}^{S}\rightarrow\mathbb{R}$ is given by $\mbox{LSE}(U_{0},\cdots,U_{S-1};S)=\ln\left(\sum_{s=0}^{S-1}e^{U_{s}}\right)$. the function $\ell_{\xi}(y|X,\beta)$ is convex in $\beta$ for any polyad $\xi$ and hence the loss function is convex. Second, given two active polyads $\xi$ and $\xi^{\prime}$, the terms $\ell_{\xi}(Y|X,\beta)$ and $\ell_{\xi^{\prime}}(Y|X,\beta)$ in the above some are not independent if the two polyads share at least one edge, i.e., if $|\mathcal{E}(\xi)\cap\mathcal{E}(\xi^{\prime})|\geq 1$.
We are now in a position to define our loss function and the associated estimator.
Definition 4.
Given the observed graph $Y$, our loss function is $\beta\to\widehat{L}_{\Xi}(Y|X,\beta)$ and the polyads estimator of the parameter of interest $\beta_{\star}$ is given by
$\widehat{\beta}_{\Xi}=\operatorname*{arg\,min}_{\beta}\widehat{L}_{\Xi}(Y|X,\beta).$ (14)
The above polyad estimator can be interpreted in relation to a logit classification problem. Specifically, given a polyad $\xi$ and the observed orbit $\mathcal{O}_{\xi}(Y)$, the minimal rank $m_{\xi}(Y)$ is distributed according to a conditional Logit model, see McFa_Chap_1974. To see this, consider the transformed graph $\underline{y}=T^{-m_{\xi}(Y)}(Y)$, which can be thought of as the “minimal” graph in the orbit of the observed graph $y$. In the example of Figure 1, the graph $\underline{y}$ is the second graph among the six graphs shown on the bottom line (that is for $r=-m_{\xi}(y)=-1$). The observed orbit $\mathcal{O}_{\xi}(Y)$ can thus be represented as
$\mathcal{O}_{\xi}(Y):=\left\{\,T_{\xi}^{m}(\underline{y}):0\leq m\leq|\mathcal{O}_{\xi}(Y)|-1\,\right\}.$
If the polyad $\xi$ is active, the size of the orbit $|\mathcal{O}_{\xi}(Y)|=m_{\xi}(Y)+M_{\xi}(Y)+1$ is greater than 2. Changing the indices $m=r+m_{\xi}(Y)$ in (3.2) yields the conditional logit structure for the distribution of $m_{\xi}(Y)$
$\mathbb{P}_{\beta}(m_{\xi}(Y)=m_{\xi}(y)\,|\,X,Y\in\mathcal{O}_{\xi}(y))=\frac{e^{v(m_{\xi}(y),\xi;\beta)}}{\sum_{m=0}^{|\mathcal{O}_{\xi}(Y)|-1}e^{v(m,\xi;\beta)}}\ \ \mbox{ with }\ \ v(m,\xi;\beta)=m\beta^{\top}\widetilde{X}_{\xi}-\sum_{{\bf i}\in\mathcal{E}(\xi)}\ln\underline{y}_{\bf i}^{m}!$ (15)
Remark 2.
(Analogy with conditional likelihood methods used in panel and network data) For any active polyads $\xi\in\Xi_{a}$, the tensor $(s_{\xi}({\bf i}))_{\bf i}$, whose entries belong to $\{-1,0,1\}$, is a “differencing vector” in the sense of Dano_Hono_Weid_WP_2025. It applies linearly to the independent features $X$ to deliver the generalized difference-in-differences term $\widetilde{X}_{\xi}$. The corresponding transformation $T_{\xi}$ plays the same role for the count data $Y$. Conditionally on the observed graph $y$ belonging to the orbit $\mathcal{O}_{\xi}(Y)$, we thus obtain a conditional logit classification problem, with the number of alternatives, $|\mathcal{O}_{\xi}(Y)|-1$, being polyad-specific.
The conditional logit distribution of $m_{\xi}(Y)$ yields the following result.
Lemma 1.
For every polyad $\xi$ and $y=(y_{\bf i})_{{\bf i}}$, the function $\beta\to\ell_{\xi}(y|X,\beta)$ has gradient
$\nabla_{\beta}\ell_{\xi}(y|X,\beta)=\left(\operatorname{\mathbb{E}}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]-m_{\xi}(y)\right)\widetilde{X}_{\xi}$ (16)
and its Hessian is non-negative since it is given by
$\nabla_{\beta}^{2}\ell_{\xi}(y|X,\beta)=\mathbb{V}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}$ (17)
Proof.
Denote by $p_{\xi}(m;\beta)$ the conditional probability that $m_{\xi}(Y)=m$ given by (15). Notice that
$\nabla_{\beta}p_{\xi}(m;\beta)$ $=$ $mp_{\xi}(m;\beta)\widetilde{X}_{\xi}-\sum_{m^{\prime}=0}^{|\mathcal{O}_{\xi}(y)|-1}m^{\prime}p_{\xi}(m;\beta)p_{\xi}(m^{\prime};\beta)\widetilde{X}_{\xi}$ (18)
$=$ $p_{\xi}(m;\beta)\left\{m-\operatorname{\mathbb{E}}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]\right\}\widetilde{X}_{\xi}.$
Using $\ell_{\xi}(y|X,\beta)=-\ln p_{\xi}(m_{\xi}(y);\beta)$, we get
$\nabla_{\beta}\ell_{\xi}(y|X,\beta)=\widetilde{X}_{\xi}\sum_{m=0}^{|\mathcal{O}_{\xi}(y)|-1}p_{\xi}(m;\beta)\,(m-m_{\xi}(y)),$
which gives (16). Differentiating the above equality and using (18) yields
$\nabla_{\beta}^{2}\ell_{\xi}(y|X,\beta)=\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}\sum_{m=0}^{|\mathcal{O}_{\xi}(y)|-1}p_{\xi}(m;\beta)\,(m-m_{\xi}(y))\left\{m-\operatorname{\mathbb{E}}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]\right\},$
and hence (17). ∎
As mentioned above, the loss function $L(y|X,\beta)$ defined by (13) is convex in $\beta$. Thanks to Lemma 1, we can now compute its Hessian
$\nabla^{2}_{\beta}\widehat{L}_{\Xi}(y|X,\beta)=\sum_{\xi\in\Xi_{a}}\mathbb{V}_{\beta}\left[m_{\xi}(Y)|Y\in\mathcal{O}_{\xi}(y)\right]\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}$
and discuss strict convexity.
Lemma 2.
The loss function $\widehat{L}_{\Xi}(y|X,\beta)$ defined by (13) is strictly convex if and only if $\sum_{\xi\in\Xi_{a}}\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}\succ 0$.
In other words, $\widehat{L}_{\Xi}(y|X,\beta)$ is strictly convex in $\beta$ if the DiD-features vary in the data.
Proof.
Since $|\Xi_{a}|<\infty$ for any finite sample, $|\mathcal{O}_{\xi}(y)|>1$ for all $\xi\in\Xi_{a}$ and the shape of the distribution of $m_{\xi}(Y)$ in (15), the minimum $\underline{\sigma}^{2}(\beta):=\min_{\xi\in\Xi_{a}}\mathbb{V}_{\beta}\left[m_{\xi}(Y)|Y\in\mathcal{O}_{\xi}(y)\right]$ is strictly positive for all $\beta\in\mathbb{R}^{p}$. Therefore, for all $\beta\in\mathbb{R}^{p}$,
$\nabla^{2}_{\beta}L(y|X,\beta)\succeq\underline{\sigma}^{2}(\beta)\sum_{\xi\in\Xi_{a}}\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}.$
4 Large sample properties of the polyads estimator
In this section we establish consistency and asymptotic normality of the polyads estimator introduced in Section 3. The concept of an active polyad is central to both our theoretical results and the implementation of the method (see Section 5). Recall that a polyad $\xi$ is active when its orbit satisfies $|\mathcal{O}_{\xi}(Y)|>1$, meaning it exhibits variation from which the parameters can be identified. This observation motivates normalizing the loss function by the expected number of active polyads and suggests deriving limiting behavior as the number of active polyads increases.
Let $\widehat{N}_{a}=|\Xi_{a}|$ denote the observed number of active polyads and let $N_{a}$ denote its conditional expectation given $X$ under Assumption 1. Formally,
$\widehat{N}_{a}:=\sum_{\xi\in\Xi}\mathbb{I}\{\xi\text{ is active}\}\text{ and }N_{a}:={\mathbb{E}}_{\beta_{\star}}[\widehat{N}_{a}|X]=\sum_{\xi\in\Xi}\mathbb{P}_{\beta_{\star}}(\xi\text{ is active}|X).$ (19)
As mentioned in Remark 2, $N_{a}$ plays the role of the average of number of data, we therefore normalize the loss function by $N_{a}$, defining
$\widehat{Q}_{\Xi}(\beta)=\frac{1}{N_{a}}\widehat{L}_{\Xi}(Y|X,\beta)\text{ and }Q_{\Xi}(\beta)={\mathbb{E}}_{\beta_{\star}}\left[\widehat{Q}_{\Xi}(\beta)\mid X\right].$ (20)
In machine learning $\beta\to Q_{\Xi}(\beta)$ is refered as the risk function.
Recall that $n=\prod_{d=1}^{D}n_{d}$ is the size of the graph indexed by $\mathcal{I}=[n_{1}]\times[n_{2}]\times\dots\times[n_{D}]$ and that if $\Xi$ is the set of polyads given by the indices $\mathcal{I}$, then $|\Xi|$ has order $n^{2}$. Through this section, one is given a sequence $\left(\Xi^{(1)},\Xi^{(2)},\dots\right)$ of families of polyads, defined over growing sets of indices. Let $n_{d}^{(k)}$ be the size of the $d$-th dimension associated with $\Xi^{(k)}$. We impose no restrictions on how each dimension $n_{d}^{(k)}$ grows, requiring only that $n^{(k)}=\prod_{d=1}^{D}n_{d}^{(k)}\to\infty$ — or, equivalently, $|\Xi^{(k)}|\to\infty$. Notably, our results accommodate short panels where one dimension remains bounded while others diverge. For three-way models, this includes settings where the PPML estimator suffers from the incidental parameter problem (weidner2021bias). We also do not assume the existence of a limiting function $Q_{\infty}$, thus avoiding restrictive assumptions on the asymptotic behavior of fixed effects and covariates. In particular, we do not impose that fixed effects or covariates are drawn from any probability distribution. To circumvent a convoluted notation we use $(k)$ as index in place of $\Xi^{(k)}$, for instance, $\widehat{\beta}_{\Xi^{(k)}}$ is written as $\widehat{\beta}^{(k)}$. Also, notice that associated with each $(k)$ we also have covariates $X^{(k)}$, fixed effects $\theta^{(k)}$ and random variables $Y^{(k)}$.
Finally, unlike graham2017econometric; jochmans2018semiparametric, we do not require the components of $\left[n_{d}^{(k)}\right]$ to arise from random sampling. This permits a more general interpretation of the data generating process. Consider a two-way model of doctor-patient interactions. Under the framework of graham2017econometric; jochmans2018semiparametric, one assumes the existence of a large population graph containing all doctors and patients, from which patients and doctors are randomly sampled, inducing distributions on both fixed effects and covariates. Our approach is more constructive: for each $(k)$, the distribution of fixed effects and covariates may differ entirely. In the doctor-patient example, our framework accommodates data collection that expands geographically. For instance, initially observing patients and doctors in Paris only, then adding doctors from Marseilles, then adding also patients from Marseilles, and so forth. Imposing a random sampling structure would limit this possibility, as any finite sample could contain doctors from Marseilles with positive probability.
4.1 Consistency
The polyads estimator (14) is defined as the minimizer of a loss function that is itself the sum of possibly non-independent terms. Our consistency result proceeds in two steps. First, we show that the normalized empirical loss $\widehat{Q}^{(k)}$ converges to its expectation $Q^{(k)}$ as $k\to\infty$. Second, we establish regularity of $Q^{(k)}$ around its minimum $\beta_{\star}$. Note that at no point do we require the existence of a limit risk function $Q_{\infty}$.
Assumption 2 controls the approximation between the empirical and expected versions of $\widehat{Q}^{(k)}$. Observe that $\ell_{\xi}$ and $\ell_{\xi^{\prime}}$ are independent as long as $\xi$ and $\xi^{\prime}$ share no edges. Moreover, a polyad $\xi$ contributes to the loss only when it is active. Thus, Assumption 2 controls the amount of dependence across the terms of $\widehat{Q}^{(k)}$, enabling a law of large numbers for the normalized losses.
Assumption 2.
We assume that, as $k\to\infty$,
[…]
Assumption 2 is not restrictive: as $n^{(k)}$ grows, the number of positive entries $Y_{\bf i}^{(k)}$ with disjoint indices should also increase, and therefore, the number of pairs of active polyads sharing no edge should outnumber those sharing at least one edge. Notice that if one assumes a distribution on the fixed effects and on the covariates this assumption is automatically true, as in this case the probability that $\xi$ and $\xi^{\prime}$ are both active is always upper and lower bounded by a constant, thus it is just a matter of comparing the number of pairs of polyads – which has order $(n^{(k)})^{4}$ – and the number of pairs of polyads sharing an edge – which has order $(n^{(k)})^{3}$. This last approach is used by graham2017econometric and jochmans2018semiparametric.
Next, Assumption 3 ensures enough regularity of $Q^{(k)}$ around the minimum $\beta_{\star}$ so that it can be identified.
Assumption 3.
We assume that as $k\to\infty$, the polyads estimator $\widehat{\beta}^{(k)}$ defined in (14) is the unique minimizer of $\widehat{Q}^{(k)}(\beta)$. We assume that there exists $\varepsilon_{0}>0$ such that
  • there exists $L_{\star}\in\mathbb{R}$ such that $\inf_{k}Q^{(k)}(\beta_{\star})\geq L_{\star}$ and for all $\beta\in\beta_{\star}+3\varepsilon_{0}B_{2}$, there exists $L$ satisfying $\sup_{k}Q^{(k)}(\beta)\leq L$;
  • for all $0<\varepsilon\leq\varepsilon_{0}$ there exists $\eta>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and all $\beta\in\mathbb{R}^{p}$, if $\left\|\beta-\beta_{\star}\right\|_{2}=\varepsilon$ then $Q^{(k)}(\beta)-Q^{(k)}(\beta_{\star})\geq\eta$.
We are now in position to state the following consistency result:
Theorem 1 (Consistency of the polyads estimator).
Grant Assumptions 1, 2 and 3. Then, for almost all $\left(X^{(k)}\right)_{k}$ and for all $\varepsilon>0$, if $n^{(k)}\to\infty$ as $k\to\infty$, then
$\mathbb{P}_{\beta_{\star},\theta^{(k)}}^{\left.Y^{(k)}\right|X^{(k)}}\left[\left\|\widehat{\beta}^{(k)}-\beta_{\star}\right\|_{2}\geq\varepsilon\right]\to 0.$
Proof.
The proof can be found in Section C.3. The proof uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of newey1994large that requires revisiting classical results in convex analysis (such as those from Chapter 10 in rockafellar-1970a), as well as some results in asymptotic statistics from newey1994large and Andersen1982. ∎
4.2 Asymptotic normality
Before introducing the assumptions and main result on the asymptotic normality of the polyads estimator we first investigate the error $\beta^{(k)}-\beta_{\star}$. A Taylor approximation yields, for some $\overline{\beta}^{(k)}$ between $\beta_{\star}$ and $\widehat{\beta}^{(k)}$,
$\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)-\nabla\widehat{Q}^{(k)}\left(\widehat{\beta}^{(k)}\right)=\nabla^{2}\widehat{Q}^{(k)}\left(\overline{\beta}^{(k)}\right)\left(\beta_{\star}-\widehat{\beta}^{(k)}\right),$
which implies
$\widehat{\beta}^{(k)}-\beta_{\star}=-\left(\nabla^{2}\widehat{Q}^{(k)}\left(\overline{\beta}^{(k)}\right)\right)^{-1}\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right).$ (21)
If $\widehat{\beta}^{(k)}$ is close to $\beta_{\star}$, $\widehat{Q}^{(k)}$ is a good approximation of $Q^{(k)}$ and $\nabla^{2}Q^{(k)}(\beta_{\star})\to\Gamma$ for some invertible $\Gamma$ we may approximate
$\widehat{\beta}^{(k)}-\beta_{\star}\approx-\Gamma^{-1}\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right).$
The last approximation shows that to control the fluctuations of $\widehat{\beta}^{(k)}-\beta_{\star}$ we need to control the fluctuations of $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$, which is not a sum of independent random variables. Following graham2017econometric; jochmans2018semiparametric, we use Hájek projections to approximate $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$ by a sum of independent random variables for which classical central limit theorems hold. The following assumption is useful to control the error of this approximation:
Assumption 4.
Given $c\in\mathbb{R}^{p}$, define $w_{\xi}=\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr$>. We assume that, as $k\to\infty$,
$\operatorname{\mathbb{E}}_{\beta_{\star},\theta^{(k)}}\left[\left.\sum_{\xi,\xi^{\prime}\in\Xi^{(k)}}w_{\xi}w_{\xi^{\prime}}\mathbf{1}_{|\mathcal{E}(\xi)\cap\mathcal{E}(\xi^{\prime})|=1}\right|X^{(k)}\right]\gg\operatorname{\mathbb{E}}_{\beta_{\star},\theta^{(k)}}\left[\left.\sum_{\xi,\xi^{\prime}\in\Xi^{(k)}}|w_{\xi}w_{\xi^{\prime}}|\mathbf{1}_{|\mathcal{E}(\xi)\cap\mathcal{E}(\xi^{\prime})|\geq 2}\right|X^{(k)}\right].$
Assumption 4 bears resemblance to Assumption 2. Assumption 2 asks for the active polyads to be distributed among the edges of $Y_{\bf i}$ in a way that no edge $\bf i$} has a substantial amount of active polyads $\xi$ such that ${\bf i}\in\mathcal{E}(\xi)$. Meanwhile and up to the weights $w_{\xi}$, which are zero when $\xi$ is not active, Assumption 4 looks at the cases where two polyads share one edge and requires that, up to some weights, in most of these cases only one edge is being shared.
Assumption 5 below is a technical assumption to control the convergence of the Hessian:
Assumption 5.
We assume that
  • $\left(Q^{(k)}\right)_{k}$ is equicontinuous, i.e., for all $\beta_{0}\in\mathbb{R}^{p}$, for all $\varepsilon>0$, there exists $k_{1}$ and $\eta>0$ such that for all $k\geq k_{1}$ and all $\beta_{1}\in\mathbb{R}^{p}$, if $\left\|\beta_{1}-\beta_{0}\right\|_{2}\leq\eta$ then $|Q^{(k)}(\beta_{0})-Q^{(k)}(\beta_{1})|\leq\varepsilon$;
  • $\nabla^{2}Q^{(k)}(\beta_{\star})\to\Gamma$ is invertible;
  • the sequence of Hessian of $\left(Q^{(k)}\right)_{k}$ is equicontinuous over $B_{2}(\beta_{\star},1)$.
Moreover, we assume that the fixed effects and the covariates are uniformly bounded as $k\to\infty$.
In order to present the main result we introduce the quantity $\Sigma^{(k)}$, which is the covariance of the Hájek projection of $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$ and is given by
$\Sigma^{(k)}=\left(N_{a}^{(k)}\Gamma\right)^{-1}\left(\sum_{{\bf i}\in{\cal I}}{\mathbb{E}}_{\beta_{\star}}\left[\left.\bar{s}^{(k)}_{{\bf i}}\bar{s}^{(k)\top}_{{\bf i}}\right|X^{(k)}\right]\right)\left(N_{a}^{(k)}\Gamma\right)^{-1},$ (22)
where
[…]
Under Assumption 4, $\Sigma^{(k)}$ will play the role of the asymptotic variance of the polyads estimator if it does not vanish to zero as $k\to+\infty$ and if a third moment condition holds. We gather these last conditions in Assumption 6 below:
Assumption 6.
Given $c\in\mathbb{R}^{p}$. We assume that
$N_{a}^{-1}\operatorname{\mathbb{E}}_{\beta_{\star},\theta^{(k)}}\left[\left.\sum_{\xi,\xi^{\prime}\in\Xi^{(k)}}\mathbf{1}_{\xi\text{ and }\xi^{\prime}\text{ are active}}\right|X^{(k)}\right]\not\gg\operatorname{\mathbb{E}}_{\beta_{\star},\theta^{(k)}}\left[\left.\sum_{\xi,\xi^{\prime}\in\Xi^{(k)}}w_{\xi}w_{\xi^{\prime}}\mathbf{1}_{|\mathcal{E}(\xi)\cap\mathcal{E}(\xi^{\prime})|=1}\right|X^{(k)}\right].$ (23)
[…]
$S_{3}^{(k)}+a^{(k)}\sqrt{V_{3}^{(k)}}\ll\left(S_{2}^{(k)}-a^{(k)}\sqrt{V_{2}^{(k)}}\right)^{3/2}.$ (24)
Assumption (24) is useful to bound the third moment of the Hájek projection of $\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)$ in a central limit theorem and is not restrictive. It is satisfied as long as the quantities $\bar{s}_{\bf i}$ (and their variances) are not concentrated on a limited number of edges $\bf i$} as $k\to\infty$.
Theorem 2 (Asymptotic normality of the polyads estimator).
Grant Assumptions 1 to 6 for some $c\in\mathbb{R}^{p}$, $c\neq 0$. It holds that, for almost all $\left(X^{(k)}\right)_{k}$, conditionally on $\left(X^{(k)}\right)_{k}$, if $n^{(k)}\to\infty$ as $k\to\infty$, then
$\frac{\bigl<\widehat{\beta}^{(k)}-\beta_{\star},c\bigr>}{\sqrt{c^{\top}\Sigma^{(k)}c}}\overset{d}{\to}{\cal N}(0,1).$
Proof.
The proof of Theorem 2, given in Section B.2, follows from a general lemma presented in Section B.1, adapted from Brunel’s lecture notes victor, which reference MR1026303 and MR1186263. Unlike the classical setup, where the loss is a sum of independent terms, we handle dependent polyads and do not assume the existence of a limiting $Q_{\infty}$, only a limiting covariance matrix at $\beta_{\star}$. Hence, we extend the standard asymptotic normality proof under convexity to this dependent setting. ∎
In Theorem 2, the asymptotic normality is written in terms of $\Sigma^{(k)}$, which cannot be directly evaluated from the sample. In Section 5.4 we discuss two alternatives to approximate $\Sigma^{(k)}$ from the sample; in Section 6 we provide empirical validation that the confidence intervals obtained by each approach are accurate.
5 Computational implementation
Lemma 1 provides the tools to solve the optimization problem (14) defining the polyads estimator $\widehat{\beta}_{\Xi}$. Since each term in the loss function is convex and we have expressions for both their gradients and their Hessians, we can efficiently solve the minimization problem using Newton’s method. However, computing the gradient and Hessian involves summing over all polyads $\xi\in\Xi$, which naively requires looping over $\prod_{d=1}^{D}n_{d}(n_{d}-1)$ terms, an operation that quickly becomes computationally expensive. The main goal of this section is to reduce this complexity by avoiding unnecessary iterations, which is done characterizing the set of active polyads $\Xi_{a}$. Besides that, we also discuss the implementation of two approximations of the variance. These approximations are essential to construct confidence intervals. The methods presented in this section are of special interest when $Y$ is sparse. In particular, the computational complexity of our methods outperforms the PPML alternative correia2019ppmlhdfe when the size of $E=\{{\bf i}:Y_{\bf i}>0\}$ is of order smaller than $\sqrt{n}$.
5.1 Permutating polyads
A key tool that we explore to obtain efficient computational implementations of the polyads estimator is their invariance to permutations. A permutation of a polyad defined by $({\bf i},{\bf i}^{\prime})$ is obtained by flipping some indices between $\bf i$} and $\bf i$$^{\prime}$. For instance,
[…]
We say that a permutation is odd when an odd number of indices are flipped and even when an even number of indices are flipped. Notice that each polyad $\xi$ has a total of $2^{D}$ unique permutations, including itself. Lemma 3 below collects the main tools that will be necessary in this section:
Lemma 3.
It holds that:
  1. (i)
    If ${\bf i}\in\mathcal{E}(\xi)$ for some $\xi$, then $\bf i$} belongs to all $2^{D}$ permutations of $\xi$ and exists exactly one permutation that can be written as $({\bf i},{\bf i}^{\prime})$ for some $\bf i$$^{\prime}\in\mathcal{I}$.
  2. (ii)
    Let $\xi^{\prime}$ be any permutation of $\xi$. Then $s_{\xi^{\prime}}({\bf i})=s_{\xi}({\bf i}),\forall{\bf i}$ iff the permutation is even and $s_{\xi^{\prime}}({\bf i})=-s_{\xi}({\bf i}),\forall{\bf i}$ iff the permutation is odd. In particular, for odd permutations $m_{\xi^{\prime}}(y)=M_{\xi}(y)$ and $M_{\xi^{\prime}}(y)=m_{\xi}(y)$ for all $y\in\mathbb{Z}^{\mathcal{I}}$ and for even permutations $m_{\xi^{\prime}}(y)=m_{\xi}(y)$ and $M_{\xi^{\prime}}(y)=M_{\xi}(y)$ for all $y\in\mathbb{Z}^{\mathcal{I}}$.
  3. (iii)
    If $\xi^{\prime}$ is a permutation of $\xi$, then $\ell_{\xi}(y|X,\beta)=\ell_{\xi^{\prime}}(y|X,\beta)$ for all $\beta\in\mathbb{R}^{p}$.
Proof.
To see (i), notice that flipping the $d$-th index of $\bf i$} with the $d$-th index of $\bf i$$^{\prime}$ produces a new edge that still belongs to $\mathcal{E}(\xi)$. Repeating this operation for all possible combinations of indices produces all $2^{D}$ permutations. Also, given ${\bf i}\in\mathcal{E}(\xi)$, to find the unique $\bf i$$^{\prime}$ such that $({\bf i},{\bf i}^{\prime})$ is a permutation of $\xi$ one just needs to flip the $d$-th index of $\bf i$} with the $d$-th index of $\bf i$$^{\prime}$ if and only if $i_{d}\neq i_{d}^{\prime}$.
To see (ii), notice that flipping one index changes the sign of $s_{\xi}({\bf i})$, thus flipping an even number of indices preserves the sign while flipping an odd number of indices changes it. The expressions for $m_{\xi^{\prime}}(y)$ and $M_{\xi^{\prime}}(y)$ follow directly from the definition.
Finally, (iii) follows observing that $\mathcal{O}_{\xi}(y)=\mathcal{O}_{\xi^{\prime}}(y)$ and so
[…]
Since the event $M_{\xi^{\prime}}(Y)=M_{\xi^{\prime}}(y)$ is equivalent to $m_{\xi^{\prime}}(Y)=m_{\xi^{\prime}}(y)$ we are done. ∎
5.2 Efficiently gathering all active polyads
A direct conclusion of Lemma 3 is that if $\xi\in\Xi_{a}$, then exists a permutation $\xi^{\prime}$ of $\xi$ such that $i_{d}<i_{d}^{\prime}$ for all $d=2,\dots,D$ and $m_{\xi^{\prime}}(y)>0$. If $M_{\xi^{\prime}}=0$ then this permutation is unique, but if $M_{\xi^{\prime}}>0$ then there are two such permutations, one with $i_{1}<i_{1}^{\prime}$ and another with $i_{1}>i_{1}^{\prime}$. Based on this observation we define the following set:
$\Xi_{a}^{\star}=\left\{\xi=({\bf i},{\bf i}^{\prime}):m_{\xi}(y)>0\text{ and }i_{d}<i_{d}^{\prime}\,\forall d=2,\dots,D\text{ and }(M_{\xi}(y)=0\text{ or }i_{1}<i_{1}^{\prime})\right\}.$
The set $\Xi_{a}^{\star}$ contains exactly one permutation of each active polyad $\xi\in\Xi_{a}$ and, by part (iii) of Lemma 3,
$\widehat{L}_{\Xi}(y|X,\beta)=2^{D}\sum_{\xi\in\Xi_{a}^{\star}}\ell_{\xi}(y|X,\beta)\text{ for all }\beta\in\mathbb{R}^{p}.$ (25)
Thus, to solve (9) it suffices to look at all polyads in $\Xi_{a}^{\star}$. The definition of $\Xi_{a}^{\star}$ also leads to an efficient method to construct it. Notice that $m_{\xi}(y)$ it is positive if and only if $y_{\bf i}>0$ for all $\bf i$} with $s_{\xi}({\bf i})=1$. In particular, we need at least $y_{\bf i}$ to be positive to have $m_{\xi}(y)$ positive. This suggests looping over pairs ${\bf i},{\bf i}^{\prime}\in E$. The procedure to do it differs slightly depending on the parity of $D$.
First, take $D=2$ and let $i_{1}\neq i_{1}^{\prime}$ be given. To have $({\bf i},{\bf i}^{\prime})\in\Xi_{a}^{\star}$ we need to find $i_{2}\neq i_{2}^{\prime}$ such that $y_{i_{1}i_{2}}\wedge y_{i_{1}^{\prime}i_{2}^{\prime}}>0$. Now let $D=3$ and $i_{1}\neq i_{1}^{\prime}$ be given. We search for $i_{2}\neq i_{2}^{\prime}$ and $i_{3}\neq i_{3}^{\prime}$ satisfying $y_{i_{1}i_{2}i_{3}}\wedge y_{i_{1}i_{2}^{\prime}i_{3}^{\prime}}\wedge y_{i_{1}^{\prime}i_{2}i_{3}^{\prime}}\wedge y_{i_{1}^{\prime}i_{2}^{\prime}i_{3}}>0$, in particular, $y_{i_{1}i_{2}i_{3}}\wedge y_{i_{1}i_{2}^{\prime}i_{3}^{\prime}}>0$. More generally, given $i_{1}$ we let
$E_{i_{1}}=\{(j_{2},\dots,j_{D}):y_{i_{1}j_{2}\dots j_{D}}>0\},$
it holds that if $({\bf i},{\bf i}^{\prime})\in\Xi_{a}^{\star}$, then (i) $(i_{2},\dots,i_{D})\in E_{i_{1}}$ and $(i_{2}^{\prime},\dots,i_{D}^{\prime})\in E_{i_{1}^{\prime}}$ when $D$ is even; or (ii) $(i_{2},\dots,i_{D}),(i_{2}^{\prime},\dots,i_{D}^{\prime})\in E_{i_{1}}$ when $D$ is odd. Thus, one only needs to loop over the pairs $i_{1}\neq i_{1}^{\prime}$ and, for each of these pairs, over $((i_{2},\dots,i_{D}),(i_{2}^{\prime},\dots,i_{D}^{\prime}))\in E_{i_{1}}\times E_{i_{1}^{\prime}}$ (if $D$ is even) or $((i_{2},\dots,i_{D}),(i_{2}^{\prime},\dots,i_{D}^{\prime}))\in E_{i_{1}}\times E_{i_{1}}$ (if $D$ is odd). Then, one simply verifies the remaining conditions for $({\bf i},{\bf i}^{\prime})\in\Xi_{a}^{\star}$. This procedure is summarized in Algorithm 1.
Input: $\{E_{i_{1}}\}_{i_{1}\in[n_{1}]}$
Output: $\Xi_{a}^{\star}$
1
2for $i_{1}\in[n_{1}]$ do
3 for $i_{1}^{\prime}\in[n_{1}]$, $i_{1}^{\prime}\neq i_{1}$ do
4    for $(i_{2},\dots,i_{D})\in E_{i_{1}}$ do
5       $E^{\prime}\leftarrow E_{i_{1}^{\prime}}$ if $D\%2=0$ else $E_{i_{1}}$;
6
7       for $(i_{2}^{\prime},\dots,i_{D}^{\prime})\in E^{\prime}$ such that $i_{d}<i_{d}^{\prime}$ for all $d=2,\dots,D$ do
8          $\xi=((i_{1},\cdots,i_{D}),(i_{1}^{\prime},\cdots,i_{D}^{\prime}))$;
9
10          if $m_{\xi}(y)>0$ and ($M_{\xi}(y)=0$ or $i_{1}<i_{1}^{\prime}$) then
11             $\Xi_{a}^{\star}\leftarrow\Xi_{a}^{\star}\cup\{\xi\}$;
12
13
14
15
16
Algorithm 1 Construct $\Xi_{a}^{\star}$
The next theorem establishes the computational complexity of constructing $\Xi_{a}^{\star}$ using Algorithm 1.
Theorem 3.
When $D$ is odd assume there exists $c\geq 1$ that $|E_{i_{1}}|<c\frac{|E|}{n_{1}}$ for all $i_{d}\in[n_{d}]$. Make no assumption if $D$ is even. The set $\Xi_{a}^{\star}$ can be computed in $O(|E|^{2})$ using Algorithm 1.
Proof.
First, notice that checking for $m_{\xi}(y)>0$ and $M_{\xi}(y)=0$ requires checking the values of all $2^{D}$ edges in $\mathcal{E}(\xi)$. Implementing the sets $E_{i_{1}}$ as hash tables allows us to check these values in constant time. Thus, we just need to count the number of times the innermost loop is executed.
If $D$ is even, the innermost loop is executed $\sum_{i_{1}\neq i_{1}^{\prime}}|E_{i_{1}}||E_{i_{1}^{\prime}}|=\left(\sum_{i_{1}}|E_{i_{1}}|\right)^{2}-\sum_{i_{1}}|E_{i_{1}}|^{2}\leq|E|^{2}$ times. If $D$ is odd, the innermost loop is executed $\sum_{i_{1}\neq i_{1}^{\prime}}|E_{i_{1}}|^{2}\leq c\frac{|E|}{n_{1}}\sum_{i_{1}\neq i_{1}^{\prime}}|E_{i_{1}}|\leq c|E|^{2}$ times. Thus, in both cases the total complexity is $O(|E|^{2})$. ∎
Remark 3.
In practice, one not only keep track of the polyads in $\Xi_{a}^{\star}$ but also of their corresponding edge values $\{y_{\bf i}:{\bf i}\in\mathcal{E}(\xi)\}$ and of $\widetilde{X}_{\xi}$. This precomputation allows us to avoid recomputing these quantities at each iteration of the optimization algorithm. Besides that, implementing $E_{i_{1}}$ as an ordered list allows the usage of binary search, which although theoretically slower than a hash table, tends to be faster in practice.
Another key point is that we do not require all features $X_{\bf i}$ to be precomputed. All that suffices is a function that can map $\bf i$} into $X_{\bf i}$, this function will be called $2^{D}$ times for each active polyad to obtain $\widetilde{X}_{\xi}$. This is essential to get an efficient implementation of our method, otherwise the computational cost would be at least the cost of computing all features, which is $O(n)$.
5.3 Solving the optimization problem
Once the set of polyads $\Xi_{a}^{\star}$ is computed, we minimize the loss (25) using Newton’s method. Lemma 1 provides closed-form expressions for the gradient and Hessian of each $\ell_{\xi}$, so a Newton step can be computed exactly. When the loss is strictly convex (see Lemma 2), Newton’s method converges from any initial value $\beta^{0}$. Algorithm 2 displays one update step from $\beta^{t}$ to $\beta^{t+1}$ for $t\geq 0$.
Input: $\Xi_{a}^{\star},\beta^{t}$
Output: $\beta^{t+1}$
1
[…]
3 $H\leftarrow 0$;
4
5for $\xi\in\Xi_{a}^{\star}$ do
6 $(\mu,\sigma^{2})\leftarrow\texttt{EvaluateMoments}(\xi,\beta^{t})$;
7 $g\leftarrow g+(\mu-m_{\xi}(y))\widetilde{X}_{\xi}$;
8 $H\leftarrow H+\sigma^{2}\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}$;
9
10
[…]
Algorithm 2 Newton update for $\beta^{t}$
The function EvaluateMoments in Algorithm 2 must return the expectation and variance of $m_{\xi}(Y)$ conditioned on $Y\in\mathcal{O}_{\xi}(y)$ when $Y$ has law parametrized by $\beta=\beta^{t}$. Notice that from (15), for all $\beta$,
$\mathbb{P}_{\beta}\!\left(m_{\xi}(Y)=m\,\middle|\,X,Y\in\mathcal{O}_{\xi}(y)\right)$ $=\frac{\exp\bigl(v(m,\xi;\beta)\bigr)}{\sum_{m^{\prime}=0}^{m_{\xi}(y)+M_{\xi}(y)}\exp\bigl(v(m^{\prime},\xi;\beta)\bigr)}$
$=\frac{\exp\bigl(v(m,\xi;\beta)-v(0,\xi;\beta)\bigr)}{\sum_{m^{\prime}=0}^{m_{\xi}(y)+M_{\xi}(y)}\exp\bigl(v(m^{\prime},\xi;\beta)-v(0,\xi;\beta)\bigr)},$
where
$v(m,\xi;\beta)=m\,\beta^{\top}\widetilde{X}_{\xi}\;-\;\sum_{{\bf i}\in\mathcal{E}(\xi)}\ln\bigl(\underline{y}_{\bf i}^{\,m}!\bigr),$
and $\underline{y}_{\bf i}^{\,m}=y_{\bf i}+(m-m_{\xi}(y))s_{\xi}({\bf i})$ . Directly evaluating $v(m,\xi;\beta)$ for each $m$ would require computing log factorials, we avoid this computation by noticing that
$v(m,\xi;\beta)-v(m-1,\xi;\beta)=\beta^{\top}\widetilde{X}_{\xi}+\sum_{{\bf i}:s_{\xi}({\bf i})=-1}\ln\!\bigl(y_{\bf i}-(m-1-m_{\xi}(y))\bigr)-\sum_{{\bf i}:s_{\xi}({\bf i})=1}\ln\!\bigl(y_{\bf i}+(m-m_{\xi}(y))\bigr),$
so the values $v(m,\xi;\beta)-v(0,\xi;\beta)$ can be computed sequentially by cumulative summation without a log factorial.
Input: $\xi,\beta$
Output: $\mu,\sigma^{2}$
1
[…]
$Z\leftarrow 1$ ;
// normalizing constant (unscaled)
3 for $m=1$ to $m_{\xi}(y)+M_{\xi}(y)$ do
4 $v_{m}\leftarrow v_{m-1}+\beta^{\top}\widetilde{X}_{\xi}+\sum_{{\bf i}:s_{\xi}({\bf i})=-1}\ln\!\bigl(y_{\bf i}-(m-1-m_{\xi}(y))\bigr)-\sum_{{\bf i}:s_{\xi}({\bf i})=1}\ln\!\bigl(y_{\bf i}+(m-m_{\xi}(y))\bigr)$;
5 $Z\leftarrow Z+e^{v_{m}}$;
6
7
8for $m=0$ to $m_{\xi}(y)+M_{\xi}(y)$ do
9 $p_{m}\leftarrow e^{v_{m}}/Z$;
10
11
[…]
13 $\sigma^{2}\leftarrow\sum_{m}m^{2}p_{m}-\mu^{2}$;
Algorithm 3 Evaluate $\mathbb{E}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]$ and $\mathbb{V}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]$
The procedure EvaluateMoments, given in Algorithm 3, computes the moments $\mu,\sigma^{2}$ required in Algorithm 2 using approximately $2^{D}|\mathcal{O}_{\xi}(y)|$ operations. Since this cost scales linearly with the orbit size, evaluating all $Y\in\mathcal{O}_{\xi}(y)$ may become prohibitive when the orbit is large. In practice, whenever $|\mathcal{O}_{\xi}(y)|$ exceeds a predefined threshold $L$, we approximate the conditional distribution of $m_{\xi}(Y)$ by restricting the computation to the truncated set
$m\in\Bigl[m_{\xi}(y)-\bigl(L/2\wedge m_{\xi}(y)\bigr),\dots,m_{\xi}(y)+\bigl(L/2\wedge M_{\xi}(y)\bigr)\Bigr].$
This truncation has negligible numerical effects, because the distribution of $m_{\xi}(Y)$ is concentrated around $m_{\xi}(y)$, and extreme values contribute essentially nothing to the expectation or variance.
Theorem 4.
Algorithm 2 runs with $O(\widehat{N}_{a})$ operations.
Proof.
Follows from the discussion above and observing that Algorithm 2 requires exactly one loop over all $\xi\in\Xi_{a}^{\star}$. ∎
5.4 Evaluating the variance
We now discuss two approaches for evaluating the covariance matrix of the polyads estimator. To shorten notation let $\nabla\ell_{\xi}\left(Y|X,\widehat{\beta}_{\Xi}\right)$ be denoted by $\nabla\widehat{\ell}_{\xi}$ and define
$\widehat{\Gamma}=\nabla^{2}L_{\Xi}\left(Y|X,\widehat{\beta}_{\Xi}\right).$
In Theorem 2, the asymptotic normality is written in terms of $\Sigma^{(k)}$, which can not be directly evaluated from the sample. In practice, (22) suggest to approximate it by $\widehat{\Sigma}$ given by
$\widehat{\Sigma}=\widehat{\Gamma}^{-1}\widehat{\Omega}\widehat{\Gamma}^{-1},\text{ where }\widehat{\Omega}=\sum_{{\bf i}\in{\cal I}}\left(\sum_{\xi:{\bf i}\in{\cal E}(\xi)}\nabla\widehat{\ell}_{\xi}\right)\left(\sum_{\xi:{\bf i}\in{\cal E}(\xi)}\nabla\widehat{\ell}_{\xi}\right)^{\top}.$
We also implement and empirically verify the performance of another variance estimator. Equation (21) suggests approximating the covariance of $\left(\nabla^{2}\widehat{Q}^{(k)}\left(\overline{\beta}^{(k)}\right)\right)\left(\widehat{\beta}^{(k)}-\beta_{\star}\right)$ by the expectation of
$\left(\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)\right)\left(\nabla\widehat{Q}^{(k)}\left(\beta_{\star}\right)\right)^{\top}=\left(\widehat{N}_{a}^{(k)}\right)^{-2}\sum_{\xi,\xi^{\prime}\in\Xi^{(k)}}\left(\nabla\ell_{\xi}\left(Y^{(k)}|X^{(k)},\beta_{\star}\right)\right)\left(\nabla\ell_{\xi^{\prime}}\left(Y^{(k)}|X^{(k)},\beta_{\star}\right)\right)^{\top}.$
Notice that if $\xi$ and $\xi^{\prime}$ share no edges, the expectation of their corresponding term is zero since it is the product of independent quantities with zero mean. This suggests approximating $\Sigma^{(k)}$ by
$\widehat{\Sigma}^{\prime}=\widehat{\Gamma}^{-1}\widehat{\Omega}^{\prime}\widehat{\Gamma}^{-1},\text{ where }\widehat{\Omega}^{\prime}=\sum_{\xi,\xi^{\prime}\text{ sharing edges}}\left(\nabla\widehat{\ell}_{\xi}\right)\left(\nabla\widehat{\ell}_{\xi^{\prime}}\right)^{\top}.$
The difference between $\Omega$ and $\Omega^{\prime}$ is subtle. The next lemma illuminates this difference and provides computationally tractable expressions for $\Omega$ and $\Omega^{\prime}$.
Lemma 4.
Let
$\mathcal{I}_{a}=\{{\bf i}:\exists\xi\in\Xi_{a}\text{ such that }{\bf i}\in\mathcal{E}(\xi)\}$
be the set of all edges that belong to at least one active polyad. It holds that
$\widehat{\Omega}$ $=\sum_{{\bf i}\in{\cal I}_{a}}\left(\sum_{{\bf i}^{\prime}:({\bf i},{\bf i}^{\prime})\in\Xi_{a}}2^{D}\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime})}\right)\left(\sum_{{\bf i}^{\prime}:({\bf i},{\bf i}^{\prime})\in\Xi_{a}}2^{D}\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime})}\right)^{\top}$ (26)
$=\sum_{{\bf i}\in{\cal I}_{a}}\sum_{{\bf i}^{\prime}:({\bf i},{\bf i}^{\prime})\in\Xi_{a}}\sum_{{\bf i}^{\prime\prime}:({\bf i},{\bf i}^{\prime\prime})\in\Xi_{a}}\left(\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime})}\right)\left(\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime\prime})}\right)^{\top}2^{2D}$
and
$\widehat{\Omega}^{\prime}=\sum_{{\bf i}\in{\cal I}_{a}}\sum_{{\bf i}^{\prime}:({\bf i},{\bf i}^{\prime})\in\Xi_{a}}\sum_{{\bf i}^{\prime\prime}:({\bf i},{\bf i}^{\prime\prime})\in\Xi_{a}}\left(\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime})}\right)\left(\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime\prime})}\right)^{\top}2^{D+\sum_{d=1}^{D}\mathbf{1}\{i_{d}^{\prime}\neq i_{d}^{\prime\prime}\}}.$ (27)
Proof.
To see (26) we start noticing that if ${\bf i}\not\in\mathcal{I}_{a}$, then its corresponding term is zero. Thus, the first sum can be taken only over ${\bf i}\in\mathcal{I}_{a}$. Now recall from Lemma 3 that if ${\bf i}\in\mathcal{E}(\xi)$ for some $\xi$, then $\bf i$} belongs to all $2^{D}$ permutations of $\xi$ and there exists exactly one permutation that can be written as $({\bf i},{\bf i}^{\prime})$ for some $\bf i$$^{\prime}$. Since each $\ell_{\xi}$ is invariant by permutation we have
$\sum_{\xi:{\bf i}\in{\cal E}(\xi)}\nabla\widehat{\ell}_{\xi}=\sum_{{\bf i}^{\prime}:({\bf i},{\bf i}^{\prime})\in\Xi_{a}}2^{D}\nabla\widehat{\ell}_{({\bf i},{\bf i}^{\prime})}$
and so (26) is proved.
The proof of (27) is more intricate. Notice that if $\xi,\xi^{\prime}$ share one edge we can put this edge “in evidence” to obtain a permutation $({\bf i},{\bf i}^{\prime})$ of $\xi$ and a permutation $({\bf i},{\bf i}^{\prime})$ of $\xi^{\prime}$. If $\xi,\xi^{\prime}$ share exactly one edge this representation is unique and a total of $2^{D}\times 2^{D}$ permutations ($2^{D}$ for each polyad in the pair) will be represented by the pair $({\bf i},{\bf i}^{\prime}),({\bf i},{\bf i}^{\prime\prime})$. Now assume that the polyads share exactly $m$ edges, in this case the representation is not unique anymore, as there are $m$ possible choices of $\bf i$}. To avoid double counting we split the $2^{D}\times 2^{D}$ permutations evenly between all possible choices of $\bf i$}, making each one account for $\frac{2^{D}\times 2^{D}}{m}$ permutations. It remains to understand, for a given pair $({\bf i},{\bf i}^{\prime}),({\bf i},{\bf i}^{\prime\prime})$, how many edges they share. Notice that if $i_{d}^{\prime}=i_{d}^{\prime\prime}$, then we can simultaneously flip the $d$-th coordinate of $\bf i$} with the $d$-th coordinate of $\bf i$$^{\prime}$ and $\bf i$$^{\prime\prime}$ to obtain a new shared edge. It can not be done if $i_{d}^{\prime}\neq i_{d}^{\prime\prime}$. Thus, the number of shared edges is $m=2^{\sum_{d=1}^{D}\mathbf{1}\{i_{d}^{\prime}=i_{d}^{\prime\prime}\}}$, which yields
$\frac{2^{D}\times 2^{D}}{2^{\sum_{d=1}^{D}\mathbf{1}\{i_{d}^{\prime}=i_{d}^{\prime\prime}\}}}=2^{D+\sum_{d=1}^{D}\mathbf{1}\{i_{d}^{\prime}\neq i_{d}^{\prime\prime}\}}$
permutations counted for each pair $({\bf i},{\bf i}^{\prime}),({\bf i},{\bf i}^{\prime\prime})$.∎
Indeed, (26) and (27) make explicit the two main differences between $\widehat{\Omega}$ and $\widehat{\Omega}^{\prime}$. First, $\widehat{\Omega}$ contains duplicates of certain pairs of polyads. A closer inspection of the proof reveals that these duplicates arise precisely on pairs that share strictly more than one edge. This clarifies the role of Assumption 4 in Theorem 2: for the projection strategy to be valid, the covariance of the Hájek projection of $\nabla\widehat{Q}^{(k)}(\beta_{\star})$ — which is approximately $\widehat{N}_{a}^{-2}\mathbb{E}\,\widehat{\Omega}$ — and the true covariance — approximately $\widehat{N}_{a}^{-2}\mathbb{E}\,\widehat{\Omega}^{\prime}$ — must converge to each other. Second, computing $\widehat{\Omega}$ is less costly than computing $\widehat{\Omega}^{\prime}$. For each $\mathbf{i}\in\mathcal{I}_{a}$, evaluating $\widehat{\Omega}$ requires only a single pass over each $\mathbf{i}^{\prime}$ such that $(\mathbf{i},\mathbf{i}^{\prime})\in\Xi_{a}$. In contrast, computing $\widehat{\Omega}^{\prime}$ requires an extra loop over $\bf i$$^{\prime\prime}$ such that $(\mathbf{i},\mathbf{i}^{\prime\prime})\in\Xi_{a}$.
[…]
Theorem 5.
Computing $\widehat{\Sigma}$ requires $O(\widehat{N}_{a})$ operations and computing $\widehat{\Sigma}^{\prime}$ requires $O(\widehat{N}_{a}m_{a})$ operations, where $m_{a}$ is the maximum over all edges $\bf i$} of the number of active polyads $\xi$ such that ${\bf i}\in\mathcal{E}(\xi)$.
Proof.
[…]
5.5 Final cost analysis and practical considerations
As a consequence of this section’s discussion, we can provide a complete computational cost analysis of our method:
Theorem 6.
Assume that $m_{a}$ is bounded. When $D$ is odd, assume also that there exists $c\geq 1$ that $|E_{i_{1}}|<c\frac{|E|}{n_{1}}$ for all $i_{d}\in[n_{d}]$. Thus, evaluating $\widehat{\beta}_{\Xi}$ and estimating its variance requires $O(T|E|^{2})$ operations, where $T$ is the number of iterations of Newton’s method.
Proof.
First notice that constructing $\Xi_{a}^{\star}$ has cost $O(|E|^{2})$, thus the size of $\Xi_{a}$, i.e. $\widehat{N}_{a}$, must be of order at most $|E|^{2}$. Each Newton’s method update has cost $O(\widehat{N}_{a})$ and, under the assumption that $m_{a}$ is bounded, both variance estimates have cost $O(\widehat{N}_{a})$. Thus, the total cost is driven by the number of iterations of Newton’s method times $|E|^{2}$. ∎
In practice, Newton’s method converges in less than $10$ iterations, yielding $O(|E|^{2})$ operations. This quantity is to be compared with the fast implementation of PPML from correia2019ppmlhdfe, which is $O(n)$. Our analysis suggests that our method is faster when $|E|\ll\sqrt{n}$ and competitive when $|E|$ is of order $\sqrt{n}$. Although being the standard practice when reporting the complexity of algorithms, the big-O notation hides a constant that matters to practitioners. In the next section, we empirically verify that, as predicted by our cost analysis, our method outperforms PPML in terms of computational time when $|E|$ is smaller than $\sqrt{n}$ and remains competitive as $|E|$ grows. Indeed, in our computational setup (see Section 6) the running time of our method is shorter than that of PPML as long as $|E|\leq 15\sqrt{n}$. This is, for example, the case of a bipartite network with $n_{1}=n_{2}$ such that the average degree of a node is at most $15$.
Remark 4.
We also note that the improvements developed in this section, and particularly the construction of $\Xi_{a}^{\star}$, may be useful for other polyads-based methods. For example, jochmans2017two proposes a generalized method of moments for two-way models with $n_{1}=n_{2}=\sqrt{n}$. His tetrad-based estimator leverages matrix-multiplication tricks and achieves a computational complexity of order $O(n^{1.1877})$ when using the best available theoretical matrix-multiplication algorithm. In contrast, our $O(|E|^{2})$ complexity can yield substantial gains in sparse regimes.
6 Experiments
We provide experiments comparing our method with PPML and with the analytical debias proposed by zylkin2024bootstrap. We consider artificial and real data. With artificial data we investigate the impact of the incidental parameter problem while knowing the correct value of $\beta_{\star}$. With real data we display evidence of the incidental parameter bias and show how it may lead towards wrong conclusions in inference. We also discuss the computational time and the effect of sparsity for all methods.
Our polyads estimator is implemented as described in Section 6, and confidence intervals use the covariance approximation $\widehat{\Sigma}^{\prime}$ from Section 5.4. All reported running times correspond to executions of $25$ seeds in parallel on a server equipped with a 32vCPU Intel Xeon Gold 6444Y and 3 TB of RAM.
6.1 Artificial data
We conduct computational experiments using a three-way data-generating process inspired by weidner2021bias; see Example 3. The dimensions are $n_{1}=n_{2}$ (varied) and $n_{3}=5$ (fixed). The fixed effects $u_{ij},w_{it},v_{jt}$ are i.i.d. $\mathcal{N}(0,1/16)$, and $\beta_{\star}=1$. The covariates $X_{ijt}$ are correlated with the fixed effects and, along the third axis, with their own past values:
$X_{ijt}=\begin{cases}\tfrac{1}{2}X_{ij(t-1)}+w_{it}+v_{jt}+\tfrac{1}{4}\mathcal{N}(0,1),&t>1,\\[3.0pt] w_{it}+v_{jt}+\tfrac{1}{4}\mathcal{N}(0,1),&t=1.\end{cases}$
The mean of the ${\bf i}=(i,j,t)$ edge’s weight satisfies
$\mathbb{E}(Y_{ijt})=\exp\!\left(c+\beta_{\star}X_{ijt}+u_{ij}+w_{it}+v_{jt}\right),$
where the constant $c$ can be selected to control the density $|E|/n$ of the graph. We generate both Poisson data (satisfying Assumption 1 with intensity $\lambda_{\bf i}=\lambda_{ijt}$) and non-Poisson data. To obtain non-Poisson outcomes, we generate $Y_{{\bf i}}$ as Negative Binomial via a Gamma–Poisson mixture: the rate of the Gamma controls the variance, while its shape is scaled to match the desired mean $\lambda_{{\bf i}}$. Setting the rate to $\infty$ recovers the Poisson model; for experiments with overdispersion, we take the rate equal to $0.1$. We executed $600$ replications of each configuration.
We compare three estimators: PPML, PPML debiased, and our polyads estimator. For PPML we use the fast implementation of correia2019ppmlhdfe; for PPML debiased we use the analytical correction of weidner2021bias via their Stata package. Our first experiment varies the graph density with
$|E|\in\{0.02n,\ 0.03n,\ 0.04n,\ 0.05n,\ 0.1n\},\quad\text{and}\quad n_{1}=n_{2}\in\{50,100\}.$
Each configuration is replicated $600$ times. Figure 2 summarizes the results. The first row shows the distributions of the normalized errors $\sqrt{n}\,(\widehat{\beta}_{\Xi}-\beta_{\star})$ at densities 2%, 5%, and 10%. The second row reports, from left to right, the empirical coverage of the 95% confidence interval (which should be close to 95%), the convergence rate, and the running time. We observe:
  • Incidental parameter bias. PPML exhibits a clear incidental parameter problem (blue curves shifted to the right), especially at low densities. Our polyads estimator eliminate this bias and the debiased PPML, at larger sample sizes $n$, do not display bias.
  • Instability at low density. For small $n$ and sparse graphs, PPML and PPML debiased often fail to converge and produce poor coverage.
  • Coverage issues. Even at larger samples, PPML intervals remain unreliable (best coverage $\approx 90\%$ for $n_{1}=n_{2}=100$ and $|E|=0.1n$). The debiased method improves coverage but still undercovers at low densities and $n$, and overcovers at high densities.
  • Computational efficiency. Our estimator is substantially faster at low densities and remains competitive as density increases.
Figure 2: Comparison of PPML, PPML debiased, and the polyads estimator across different graph densities.
Figure 3: Sparse case ($|E|=4\sqrt{n}$, Poisson model): Polyads estimator maintains good performance in sparse settings.
We next consider a second experiment aimed at evaluating performance in the sparse regime. Here we let $n$ vary from $2\times 10^{5}$ to $32\times 10^{5}$ and choose the constant $c$ so that
$|E|\approx 4\sqrt{n},$
thus forcing the density $\frac{|E|}{n}$ to shrink towards zero as the sample size grows. This setting is substantially sparser than the fixed-density designs considered above. Figure 3 reports the results for the Poisson case and Figure 4 reports the results for the Negative Binomial case. The qualitative conclusions of both cases are similar and consistent with those found in the low-density examples from Figure 2, but become even more pronounced under sparsity:
  • Estimation error. The first panel presents boxplots of the non-scaled estimation error $\widehat{\beta}_{\Xi}-\beta_{\star}$ (conditional on convergence). PPML again displays a pronounced incidental parameter bias, with its median shifted upward across all sample sizes. The debiased PPML estimator reduces, but does not eliminate—this distortion. In contrast, the polyads estimator remains concentrated around $\beta_{\star}$ for every $n$. More interestingly, as $n$ increases the variance of the PPML debiased approach shrinks, but the incidental parameter bias remains relevant in a way that its coverage starts to decay.
  • Coverage of the 95% confidence interval. The second panel highlights the severity of the incidental parameter problem in sparse settings. PPML confidence intervals exhibit near-zero coverage throughout: the bias places the estimator far outside the nominal interval in almost all replications. The debiased PPML method improves coverage but still undercovers for smaller $n$, where sparsity is most acute. The polyads estimator, by contrast, maintains coverage close to the nominal 95% level uniformly across all sample sizes.
  • Running time. The third panel reports computation times. Both PPML and debiased PPML become increasingly expensive as $n$ grows, despite the sparsity of the graph. The polyads estimator directly leverages sparsity and is substantially faster, especially for moderate and large $n$.
Across bias, coverage, computation time, and convergence, the results in this sparse regime reinforce the findings from the low-density experiment: PPML exhibits severe incidental parameter bias and essentially zero inferential validity; the debiased PPML estimator improves upon PPML but continues to suffer from miscoverage under sparsity; the polyads estimator remains accurate, fast, and statistically reliable, even when sparsity increases with sample size.
Figure 4: Sparse case ($|E|=4\sqrt{n}$, Negative Binomial model): Polyads estimator remains robust under Negative Binomial noise.
Comparing Figures 3 and 4 we notice that the Negative Binomial case closely match the Poisson one: PPML remains biased, while both the debiased PPML and our polyads estimator remove the bias and achieve better coverage. Thus, although our theory assumes a Poisson model, these results indicate that the polyads estimator is empirically robust to overdispersion and model misspecification.
6.2 Real data
We study health insurance claims data that cover the universe of physician consultations in metropolitan France over the years 2016 to 2018.777More details on data are provided in Appendix D. In May 2017, the fees charged by general practitioners (GPs) belonging to the regulated sector [^8] The majority of French GPs are subject to fee regulation. In January 2016, only 8.6% of GPs can charge fees in excess of the regulated level. have increased by 8.7%. Using a difference-in-differences approach at the doctor level, we find that the stronger financial incentives have caused physician activity (as measured by number of visits) to rise by approximately 10% (see Appendix D for details about the considered control groups and specifications, as well as Table D.1).
Given the strong policy concern about care accessibility, it is important to understand how the reform has affected potential patients. Have certain categories of patients (patients with chronic diseases, women, elderly patients, patients living in medically undeserved areas, low-income patients, etc.) been disproportionately affected by the reform? Has the reform caused patients to seek care further away from their home? To answer these questions, we need to perform the analysis at the doctor-patient level rather than at the doctor level. Using a three-way model and controlling for dyads fixed effects allows to estimate how the reform has changed the doctor-patient connections. To illustrate the method, we highlight in this subsection the role of two variables: patient sex and geographic distance (spatial accessibility).
The outcome $Y_{ijt}$ is the number of visits by patient $i$ to doctor $j$ on month $t$. We aggregate data at the city-sex level. The high number of potential patients in each city-sex group makes to the Poisson assumption plausible. [^9] The aggregate number of consultations in each group is the sum of (possibly heterogeneous) Bernoulli distributions that represent the occurrence of a consultation for all potential patients in the group. This sum follows approximately a Poisson distribution under the conditions exposed in le1960approximation. The assumption that the individual occurrences of a consultation are independent across potential patients is relaxed in galambos1973general and serfling1978some. The index $i$ (resp. $j$) stands thus for the set of patients (resp. doctors) in a given municipality with same gender. The treatment $T_{j}$ is a binary variable equal to 1 for sector 1 GPs, and to 0 for direct access specialists. The reform has been implemented from May 2017 onward, hence the definition of $\text{Post}_{t}$, a dummy variable equal to 1 after that date. The model includes three features that account for any possible policy-induced change in homophily preferences as regards gender and spatial dimensions: the interactions between $\text{Post}_{t}\times T_{j}$ (as in any difference-in-differences approach) and (i) a dummy variable equals to 1 when patients and doctors have the same sex, (ii) a dummy variable that is equal to 1 when the medical office chosen is located in the same municipality as the patient’s home, and (iii) travel time (measured in minutes between the centroids of municipalities).
We thus consider the three-way Poisson model $Y_{ijt}\sim\mathcal{P}(\lambda_{ijt})$ with intensity given by
$\ln\lambda_{ijt}=\big(\beta_{\texttt{d}}d_{ij}+\beta_{\texttt{sc}}\mathds{1}\{\text{city}_{i}=\text{city}_{j}\}+\beta_{\texttt{ss}}\mathds{1}\{\text{sex}_{i}=\text{sex}_{j}\}\big)\times\text{Post}_{t}\times T_{j}+u_{ij}+v_{jt}+w_{it}.$
We compare three estimators: PPML, PPML debiased and our polyads estimator. Because the full dataset contains $n_{1}=69{,}265$ patient groups, $n_{2}=16{,}941$ doctors, and $n_{3}=34$ months — corresponding to roughly $n\approx 40$ billion edges and |E|=56,034,015|E|=56{,}034{,}015 — direct estimation on the full graph is computationally infeasible for all methods. We therefore adopt a subsampling strategy combined with meta-analysis.
For each configuration, we sample a proportion $s\in\{2\%,3\%,4\%\}$ of patient groups and the same proportion of doctors, while always retaining all $34$ months [^10] Notice that we sample the groups of patients, the doctors and then get all edges through the sampled nodes, keeping all times. This sampling strategy is consistent with Assumption 1 and also with the sampling assumptions in fernandez2016individual; weidner2021bias; graham2017econometric; jochmans2018semiparametric. One could also propose sampling directly the edges and not the nodes, but there is evidence that this procedure can lead to bias, see subsamplegraphs for details.. Each subsample consists of independent random draws of patient and doctor groups; doctors in the treatment and control groups are sampled independently to ensure comparability. We repeat the procedure independently to obtain $100$ subsamples. The final estimates are obtained via a random-effects meta-analysis using the default implementation in the statsmodels Python library, which is based on the iterated method of paule1982consensus, a refinement of the classical method of dersimonian1986meta.
Subsample proportion
Parameter Method 2% 3% 4%
PPML 1.94 (1.43, 2.44) 1.87 (1.56, 2.18) 1.66 (1.48, 1.84)
$\beta_{\texttt{d}}(\times 10^{4})$ PPML (debiased) 15.14 (-8.33, 38.61) 1.33 (0.57, 2.09) 1.44 (0.99, 1.89)
Polyads 1.23 (0.26, 2.20) 1.73 (1.18, 2.28) 1.45 (1.03, 1.86)
PPML -6.54 (-11.45, -1.63) -4.28 (-6.47, -2.08) -4.46 (-5.88, -3.04)
$\beta_{\texttt{sc}}(\times 10^{2})$ PPML (debiased) -62.17 (-174.73, 50.40) -0.98 (-5.90, 3.94) -3.17 (-6.13, -0.21)
Polyads -7.60 (-13.10, -2.10) -3.91 (-6.55, -1.27) -4.52 (-6.28, -2.75)
PPML -1.54 (-2.84, -0.23) -0.91 (-1.67, -0.15) -0.78 (-1.29, -0.27)
$\beta_{\texttt{ss}}(\times 10^{2})$ PPML (debiased) -29.78 (-60.67, 1.11) -0.58 (-1.90, 0.74) -2.07 (-4.61, 0.47)
Polyads -2.48 (-4.45, -0.50) -0.35 (-1.61, 0.92) -0.20 (-1.13, 0.73)
Table 1: Random effects meta-analysis estimates with 95% confidence intervals (in parentheses) by subsample proportion and method. $\beta_{\texttt{d}}$ estimates are scaled by $10^{4}$; $\beta_{\texttt{sc}}$ and $\beta_{\texttt{ss}}$ by $10^{2}$.
Time (min)
Subsample (%) $n$ $|E|$ $|\Xi_{a}^{\star}|$ PPML PPML (debiased) Polyads
2% 15,958,490 22,644 151,743 19 53 5
3% 35,906,602 50,617 739,271 126 199 30
4% 63,833,959 89,520 2,301,766 203 315 111
Table 2: Computational time: average number of edges, average number of observed positive edges, average number of active polyads (without counting permutations), and average running times (in minutes) for each method across $100$ replications.
Figure 5: Estimated coefficients by subsample proportion and estimation method (100 replications each).
The results are the following. First, Table 1 and Figure 5 show that debiased PPML has many outliers, mostly on ‘small’ samples (admittedly, not that small since $|E|$ amounts to about 20,000 with the 2% subsample size), which results in (very) large confidence intervals and point estimates. As expected, the variance decreases with the subsample size, regardless of the estimation method used.
Second, the running time of PPML and debiased PPML scales with $n$ while the polyads time scales with $|E|^{2}$ (Table 2).
Third, a closer look at the 4% subsample coefficients suggests that: (i) PPML estimates of $\beta_{d}$ yields confidence intervals that do not contain the point estimate of debiased PPML and of Polyads, which may be a sign of incidental parameter bias. This is to be compared with the artificial experiments where PPML yields small variance around a biased point, missing the point estimate. All methods point towards a positive coefficient, implying that after the reform some patients travel further away from home to get healthcare; (ii) as regards the preference coefficient for being treated in the same municipality $\beta_{sc}$, the Polyads estimator yields similar results to the PPML, while debiased PPML gives a different point estimate, although the latter confidence intervals are overall consistent with PPML and Polyads. All methods point towards a negative coefficient, meaning that after the reform patients are more likely to visit a doctor outside their own city ; (iii) as regards gender homophily, we see some difference up to the level of signs. The PPML would lead to estimate a negative impact on the reform on that homophily (this diagnosis holding true regardless of the subsample size considered here), while Polyads and debiased PPML both contain 0 in their confidence intervals in the $4\%$ sample. Note also that, PPML yields a confidence interval that is close to the lower half of Polyads confidence interval, this may be a sign of the incidental parameter problem as PPML may be giving a narrow variance around a biased point estimate and leading to possibly wrong conclusions (in this case, that gender homophily between patients and doctors decreased).
Appendix A Consistency of $\widehat{\beta}_{\Xi}$ in the Poisson model
[…]
$\Delta_{q,N}(X)$ $={\boldsymbol{\mathbb{E}}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{{\boldsymbol{\xi}}}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{{\boldsymbol{\xi}}^{\prime}}(\beta_{\star}),c\bigr>I\left(|{\cal E}({\boldsymbol{\xi}})\cap{\cal E}({\boldsymbol{\xi}}^{\prime})|=q\right)|X\right]$
$=\frac{1}{N^{2}}\sum_{\xi,\xi^{\prime}}{\mathbb{E}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>I\left(|{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=q\right)|X\right]$
where $\boldsymbol{\xi}$} and $\boldsymbol{\xi}$$^{\prime}$ are iid uniformly distributed over $\Xi$.
The proof of the consistency result uses the convexity of the loss function as a key ingredient. It is based on a slight modification of Theorem 2.7 of newey1994large that requires to revisit some classical results in convex analysis (such as those from Chapter 10 in rockafellar-1970a) as well as some results in asymptotic statistics from newey1994large and Andersen1982. All these revisited versions of these classical results may be found in Section C.3. Theorem 1 is a consequence of Lemma A.1 below whose proof may be found also in Section C.3.
Lemma A.1.
Let $\Theta$ be a non-empty open and convex set in $\mathbb{R}^{d}$. Let $(\hat{Q}_{n})_{n}$ be a sequence of random convex functions and $(Q_{n})_{n}$ be a sequence of (deterministic) convex functions all defined on $\Theta$. We assume that there exists $\theta_{0}\in\Theta$ and $\varepsilon_{0}>0$ such that $\theta_{0}+3\varepsilon_{0}B_{2}\subset\Theta$ and the following holds:
  1. 1.
    for $n$ large enough, $\hat{Q}_{n}$ is uniquely minimized over $\Theta$ by $\hat{\theta}_{n}$;
  2. 2.
    for all $0<\varepsilon\leq\varepsilon_{0}$ there exists $\eta>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and all $\theta\in\Theta$, if $\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon$ then $Q_{n}(\theta)-Q_{n}(\theta_{0})\geq\eta$;
  3. 3.
    there exists $\theta_{1}\in\theta_{0}+3\varepsilon_{0}B_{2}$ and $L_{1}\in\mathbb{R}$ such that $\inf_{n}Q_{n}(\theta_{1})\geq L_{1}$ and for every $\theta\in\theta_{0}+3\varepsilon_{0}B_{2}$, there exists $L>0$ such that $\sup_{n}Q_{n}(\theta)\leq L$;
  4. 4.
    for all $\theta\in\theta_{0}+3\varepsilon_{0}B_{2}$, as $n$ tends to infinity, $\hat{Q}_{n}(\theta)-Q_{n}(\theta)\overset{p}{\to}0$.
Then $\hat{\theta}_{n}\overset{p}{\to}\theta_{0}$.
They are two main advantages of Lemma A.1: first, as in Theorem 2.7 from newey1994large, we don’t need to assume that $\Theta$ is compact (that will be useful for us since we are minimizing over all $\beta\in\mathbb{R}^{p}$ to define the polyads estimator); second, we don’t need to assume that the sequence of risk functions $(Q_{N})_{N}$ has a limit or that it is the same function for all $N$. The latter assumption is problematic when the fixed effects are considered as parameters (or when they are random variables and we work conditionally to them); the same remark holds for the co-variable vectors $(X_{{\bf i}})_{\bf i}$. This is our main motivation to use Lemma A.1 in place of the classical Theorem 2.7 from newey1994large where a limiting function is assumed to exists.
Proof of Theorem 1.
To prove Theorem 1, we apply Lemma A.1 to $\Theta=\mathbb{R}^{d}$, $n=N$ the number of polyads, for all $\beta\in\mathbb{R}^{d}$, $\hat{Q}_{N}(\beta)=N_{a}^{-1}L(Y|X,\beta)$ and $Q_{N}(\beta)=N_{a}^{-1}{\mathbb{E}}_{\beta_{\star}}[L(Y|X,\beta)|X]$ so that $\theta_{0}=\beta_{\star}$ and $\varepsilon_{0}=1/3$. Under the assumptions from Theorem 1, the first three points from Lemma A.1 are satisfied. It therefore only remains to show the fourth item of Lemma A.1, which is the pointwise convergence in probability of $\hat{Q}_{N}-Q_{N}$ to 0 for $\beta\in\beta_{\star}+B_{2}$.
Let $\beta\in\mathbb{R}^{d}$ be such that $\left\|\beta_{\star}-\beta\right\|_{2}\leq 1$. By Chebyshev’s inequality we only need to show that $N_{a}^{-2}{\mathbb{V}}_{\beta_{\star}}(L(Y|X,\beta)|X)\to 0$ as $N$ tends to infinity. Since, two polyads $\xi$ and $\xi^{\prime}$ that do not have an edge in common are independent we obtain:
${\mathbb{V}}_{\beta_{\star}}(L(Y|X,\beta)|X)=\sum_{\begin{subarray}{c}\xi,\xi^{\prime}\in\Xi\\ {\cal E}(\xi)\cap{\cal E}(\xi^{\prime})\neq\emptyset\end{subarray}}{\mathbb{E}}_{\beta_{\star}}\left[(\ell_{\xi}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi}(Y|X,\beta))(\ell_{\xi^{\prime}}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi^{\prime}}(Y|X,\beta))|X\right].$
It follows from Lemma C.9 that there exists some constant $C_{1}>0$ such that for all $\xi,\xi^{\prime}\in\Xi$, we have
${\mathbb{E}}_{\beta_{\star}}\left[(\ell_{\xi}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi}(Y|X,\beta))(\ell_{\xi^{\prime}}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi^{\prime}}(Y|X,\beta))|X\right]\leq C_{1}\mathbb{P}_{\beta_{\star}}\left[\xi\mbox{ and }\xi^{\prime}\mbox{ are both active}|X\right].$
According to Lemma C.9 we can choose
$C_{1}=c_{D}(L_{N}\left\|\beta\right\|_{2}+1)^{2}\left((\bar{\lambda}+1)^{4}+\frac{\exp(-2^{D-1}e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{2^{D-1}}}\right).$ (A.1)
where $c_{D}$ is a constant depending only on $D$, $L_{N}=\max_{\xi}\left\|\tilde{X}_{\xi}\right\|_{2}$,
[…]
(note that we used $\left\|\beta_{\star}-\beta\right\|_{2}\leq 1$ to define $\bar{\lambda}$ and $\underline{\lambda}$).
As a consequence, we get
${\mathbb{E}}_{\beta_{\star}}\left[(\hat{Q}_{N}(\beta)-Q_{N}(\beta))^{2}|X\right]\leq\frac{C_{1}}{N_{a}^{2}}\sum_{\begin{subarray}{c}\xi,\xi^{\prime}\in\Xi\\ {\cal E}(\xi)\cap{\cal E}(\xi^{\prime})\neq\emptyset\end{subarray}}\mathbb{P}_{\beta_{\star}}\left[\xi\mbox{ and }\xi^{\prime}\mbox{ are both active}|X\right]$
[…]
Appendix B Asymptotic normality of the polyads estimator
We first prove an asymptotic normality result with a theoretical asymptotic covariance matrix that show that the polyads estimator does not suffer an incidental parameter problem (IPP) unlike PPML. We then show an a.n. result with an empirical asymptotic covariance matrix that will be useful to derive data-driven confidence sets for the polyads estimator. We will use these confidence sets in our simulation studies.
B.1 Main theorem for asymptotic normality under convexity assumption
The following result is adapted from Victor-Emmanuel Brunel lecture notes victor who cites MR1026303 and MR1186263 as classical references for asymptotic statistics under convexity. However, all the later results use a loss function which is a sum of independent variables. This is not our case here since our loss function sums over all polyads that are not necessarily independent, moreover, as mentioned previously in the section on the consistency result above, we do not assume the existence of a limiting risk function; we only require the existence of a limiting covariance matrix of the risk in $\beta_{\star}$. We therefore need to adapt the classical proof of asymptotic normality under convexity assumption to our setup. The proof of asymptotic normality of the polyads estimator will rely on the following general result.
Theorem 7.
Let $\Theta$ be a non-empty, open and convex set in $\mathbb{R}^{d}$, let $\theta_{0}$ be in $\Theta$ and denote $\varepsilon_{0}>0$ such that $B_{2}(\theta_{0},3\varepsilon_{0})\subset\Theta$. Let $(\hat{Q}_{n})_{n}$ be a sequence of random functions defined on $\Theta$ and denote $Q_{n}={\mathbb{E}}\hat{Q}_{n}$. We assume that
  1. 1.
    for $n$ large enough, $Q_{n}$ is convex and twice differentiable, it is uniquely minimized at $\theta_{0}$ and as $n$ tends to infinity, $\nabla^{2}Q_{n}(\theta_{0})\to H_{0}$ for some $H_{0}\succ 0$; we also assume that $(\nabla^{2}Q_{n})_{n}$ is equi-continuous at $\theta_{0}$ (i.e. for all $0<\varepsilon\leq\varepsilon_{0}$, there exists $n_{1}$ and $\eta>0$ such that for all $n\geq n_{1}$ and all $\theta_{1}\in\Theta$, if $\left\|\theta_{1}-\theta_{0}\right\|_{2}\leq\eta$ then $\left\|\nabla^{2}Q_{n}(\theta_{0})-\nabla^{2}Q_{n}(\theta_{1})\right\|_{op}\leq\varepsilon$), that there exists $\theta_{1}\in\Theta$ and $L_{1}\in\mathbb{R}$ such that $\inf_{n}Q_{n}(\theta_{1})\geq L_{1}$ and for every $\theta\in\Theta$, there exists $L>0$ such that $\sup_{n}Q_{n}(\theta)\leq L$;
  2. 2.
    for $n$ large enough, $\hat{Q}_{n}$ is convex, twice differentiable and $\inf_{\theta\in\Theta}\hat{Q}_{n}(\theta)$ is achieved at $\hat{\theta}_{n}\in\Theta$;
  3. 3.
    for the compact set $K:=B_{2}(\theta_{0},2\varepsilon_{0})$, we have $\sup_{\theta\in K}\left\|\nabla^{2}\hat{Q}_{n}(\theta)-\nabla^{2}Q_{n}(\theta)\right\|_{op}\overset{p}{\to}\;0$;
  4. 4.
    there exists $V_{0}\succ 0$ such that, as $n$ tends to infinity, $\sqrt{n}\nabla\hat{Q}_{n}(\theta_{0})\overset{d}{\to}{\cal N}(0,V_{0})$.
Then, as $n$ tends to infinity, $\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\overset{d}{\to}{\cal N}(0,\Sigma_{0})$ where $\Sigma_{0}=H_{0}^{-1}V_{0}H_{0}^{-1}$.
Proof.
For all $t\in\mathbb{R}^{d}$ such that $\theta_{0}+t\in\Theta$, we define
$Z_{n}(t):=\hat{Q}_{n}\left(\theta_{0}+\frac{t}{\sqrt{n}}\right)-\hat{Q}_{n}(\theta_{0})-\bigl<\nabla\hat{Q}_{n}(\theta_{0}),\frac{t}{\sqrt{n}}\bigr$>.
We first show that when $n$ tends to infinity
$n\left(Z_{n}(t)-{\mathbb{E}}Z_{n}(t)\right)\overset{p}{\to}0.$ (B.1)
It follows from a second order Taylor expansion of $\hat{Q}_{n}$ at $\theta_{0}$ that there exists $\hat{\theta}_{1}\in[\theta_{0},\theta_{0}+t/\sqrt{n}]$ such that
$n\left|Z_{n}(t)-{\mathbb{E}}Z_{n}(t)\right|=\frac{1}{2}\left|t^{\top}\left(\nabla^{2}\hat{Q}_{n}(\hat{\theta}_{1})-{\mathbb{E}}\nabla^{2}\hat{Q}_{n}(\hat{\theta}_{1})\right)t\right|\leq\frac{\left\|t\right\|_{2}^{2}}{2}\sup_{\theta_{1}\in[\theta_{0},\theta_{0}+t/\sqrt{n}]}\left\|\nabla^{2}Q_{n}(\theta_{1})-\nabla^{2}\hat{Q}_{n}(\theta_{1})\right\|_{op}.$
For $n$ large enough, the segment $\theta_{0},\theta_{0}+t/\sqrt{n}]$ is in $K=B_{2}(\theta_{0},2\varepsilon_{0})$ and so (B.1) follows from the assumption on the uniform convergence over $K$ of the sequence of Hessians.
The latter result holds pointwise (i.e. for a given $t$) but we know that a sequence of convex functions that converges pointwise also converges uniformly over all compact set (see Chapter 10 in rockafellar-1970a). However, we cannot use directly this result in our case since we are in a situation where it is the difference of two sequences of convex functions that converges pointwise to zero in (B.1). Hence, as in the proof of consistency of the polyads estimator from the previous section, we apply the convergence result from Lemma C.5. As a consequence, it follows from the pointwise convergence (B.1) that on the compact set $B_{2}(0,2\varepsilon_{0})$, we have
$n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|Z_{n}(t)-{\mathbb{E}}Z_{n}(t)\right|\overset{p}{\to}0.$ (B.2)
Next, we define for all $t\in\mathbb{R}^{d}$ such that $\theta_{0}+t\in\Theta$, $\hat{F}_{n}(t)=\hat{Q}_{n}(\theta_{0}+t/\sqrt{n})-\hat{Q}_{n}(\theta_{0})$ and
$\hat{H}_{n}(t)=\bigl<\nabla\hat{Q}_{n}(\theta_{0}),\frac{t}{\sqrt{n}}\bigr>+\frac{1}{2}\left(\frac{t}{\sqrt{n}}\right)^{\top}H_{0}\left(\frac{t}{\sqrt{n}}\right).$
It follows from a second order Taylor expansion of $Q_{n}$ at $\theta_{0}$ that
$n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}|\hat{F}_{n}(t)-\hat{H}_{n}(t)|=n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|Z_{n}(t)-\frac{1}{2}\left(\frac{t}{\sqrt{n}}\right)^{\top}H_{0}\left(\frac{t}{\sqrt{n}}\right)\right|$
$\leq n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|Z_{n}(t)-{\mathbb{E}}Z_{n}(t)\right|+n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|{\mathbb{E}}Z_{n}(t)-\frac{1}{2}\frac{t}{\sqrt{n}}\nabla^{2}Q_{n}(\theta_{0})\frac{t}{\sqrt{n}}\right|$
$+\frac{1}{2}\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|t^{\top}\left(\nabla^{2}Q_{n}(\theta_{0})-H_{0}\right)t\right|$
$\leq n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}\left|Z_{n}(t)-{\mathbb{E}}Z_{n}(t)\right|+\sup_{\left\|\theta_{1}-\theta_{0}\right\|_{2}\leq 2\varepsilon_{0}/\sqrt{n}}\left\|\nabla^{2}Q_{n}(\theta_{1})-\nabla^{2}Q_{n}(\theta_{0}))\right\|_{op}+2\varepsilon_{0}^{2}\left\|\nabla^{2}Q_{n}(\theta_{0})-H_{0}\right\|_{op}.$
Hence, it follows from (B.2), the equi-continuity of the sequence $(\nabla^{2}Q_{n})_{n}$ at $\theta_{0}$ and the convergence $\nabla^{2}Q_{n}(\theta_{0})\to H_{0}$ that
$n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}|\hat{F}_{n}(t)-\hat{H}_{n}(t)|\overset{p}{\to}0.$ (B.3)
By definition, $\hat{t}_{n}=\sqrt{n}(\hat{\theta}_{n}-\theta_{0})$ is a minimizer of $\hat{F}_{n}$ and $\tilde{t}_{n}=-H_{0}^{-1}(\sqrt{n}\nabla Q_{n}(\theta_{0}))$ is a minimizer of $\hat{H}_{n}$. It follows from the convergence assumption on the sequence of gradient that $\tilde{t}_{n}\overset{d}{\to}{\cal N}(0,\Sigma_{0})$; hence, by Slutsky’s Lemma, it only remains to show that $\hat{t}_{n}-\tilde{t}_{n}\overset{p}{\to}0$ to get the result.
Let $0<\varepsilon\leq\varepsilon_{0}$. Let us prove that $\mathbb{P}\left[\left\|\hat{t}_{n}-\tilde{t}_{n}\right\|_{2}\geq\varepsilon\right]\to 0$. Let $0<\varepsilon_{1}<1$. Let us show that there exists $n_{0}\in{\mathbb{N}}$ such that for all $n\geq n_{0},\mathbb{P}\left[\left\|\hat{t}_{n}-\tilde{t}_{n}\right\|_{2}\geq\varepsilon\right]\leq\varepsilon_{1}$. We have $\tilde{t}_{n}=\mathcal{O}_{p}(1)$, hence, there exists $n_{1}$ such that $\mathbb{P}\left[\left\|\tilde{t}_{n}\right\|_{2}\leq\varepsilon_{0}\right]\geq 1-\varepsilon_{1}/2$ for all $n\geq n_{1}$. It follows from (B.3) that there exists $n_{2}$ such that for all $n\geq n_{2}$, $\mathbb{P}\left[n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}|\hat{F}_{n}(t)-\hat{H}_{n}(t)|\leq\varepsilon^{2}\lambda_{min}/12\right]\geq 1-\varepsilon_{1}/2$ where $\lambda_{min}$ is the smallest singular value of $H_{0}$ (it is positive by assumption). Next, we check (directly or by doing a second order Taylor expansion of $\hat{H}_{n}$ at $\tilde{t}_{n}$) that for all $t\in\mathbb{R}^{d}$ such that $\theta_{0}+t\in\Theta$,
$\hat{H}_{n}(t)-\hat{H}_{n}(\tilde{t}_{n})=\frac{1}{2n}\left\|H_{0}^{1/2}(t-\tilde{t}_{n})\right\|_{2}^{2}\geq\frac{\lambda_{min}}{2n}\left\|t-\tilde{t}_{n}\right\|_{2}^{2}.$
Denote by $\Omega_{0}$ the event onto which $n\sup_{t:\left\|t\right\|_{2}\leq 2\varepsilon_{0}}|\hat{F}_{n}(t)-\hat{H}_{n}(t)|\leq\varepsilon^{2}\lambda_{min}/12$ and $\left\|\tilde{t}_{n}\right\|_{2}\leq\varepsilon_{0}$. Let $n$ be larger than $\max(n_{1},n_{2})$ and denote by $S_{2}(\tilde{t}_{n},\varepsilon)=\{t:\left\|t-\tilde{t}_{n}\right\|_{2}=\varepsilon\}$. On $\Omega_{0}$, we have $S_{2}(\tilde{t}_{n},\varepsilon)\subset B_{2}(0,2\varepsilon_{0})$ and so for all $t$ such that $\left\|t-\tilde{t}_{n}\right\|_{2}=\varepsilon$,
$\hat{F}_{n}(t)$ $\geq\hat{H}_{n}(t)-\frac{\varepsilon^{2}\lambda_{min}}{12n}>\hat{H}_{n}(\tilde{t}_{n})+\frac{\lambda_{min}}{2n}\left\|t-\tilde{t}_{n}\right\|_{2}^{2}-\frac{\varepsilon^{2}\lambda_{min}}{12n}=\hat{H}_{n}(\tilde{t}_{n})+\frac{5\varepsilon^{2}\lambda_{min}}{12n}$
$\geq\hat{F}_{n}(\tilde{t}_{n})-\frac{\varepsilon^{2}\lambda_{min}}{12n}+\frac{5\varepsilon^{2}\lambda_{min}}{12n}=\hat{F}_{n}(\tilde{t}_{n})+\frac{\varepsilon^{2}\lambda_{min}}{3n}.$
We proved that, on the event $\Omega_{0}$, for all $t$ such that $\left\|t-\tilde{t}_{n}\right\|_{2}=\varepsilon$, we have $\hat{F}_{n}(t)>\hat{F}_{n}(\tilde{t}_{n})$. It follows from the convexity of $\hat{F}_{n}$ that this results extends to all $t$ such that $\left\|t-\tilde{t}_{n}\right\|_{2}\geq\varepsilon$. As a consequence, we conclude that the minimizer $\hat{t}_{n}$ of $\hat{F}_{n}$ necessary lies in $B_{2}(\tilde{t}_{n},\varepsilon)$. In other words, we proved that for all $n\geq\max(n_{1},n_{2})$, $\mathbb{P}\left[\left\|\hat{t}_{n}-\tilde{t}_{n}\right\|_{2}>\varepsilon\right]\leq\mathbb{P}[\Omega_{0}^{c}]\leq\varepsilon_{1}.$
There are situations where only directional CLTs for the gradient of the loss function at $\theta_{0}$ are available and where the Crámer-Wold device does not apply. In that case, we may also prove directional CLTs for $(\hat{\theta}_{n})_{n}$.
Theorem 8.
Let $u\in\mathbb{R}^{d}$. We consider the same setup as Theorem 7 except that point 4. is replaced by:
  • $4^{\prime}$.
    There exists a sequence of PSD matrices $(\hat{V}_{n})_{n}$ such that,
    $\frac{\bigl<\sqrt{n}\nabla\hat{Q}_{n}(\theta_{0}),H_{0}^{-1}u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}\overset{d}{\to}{\cal N}(0,1)$
    and $1=\mathcal{O}_{\mathbb{P}}\left(\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}\right)$.
Then as $n$ tends to infinity,
$\frac{\bigl<\sqrt{n}(\hat{\theta}_{n}-\theta_{0}),u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}\overset{d}{\to}{\cal N}(0,1).$
Proof of Theorem 8.
The proof of Theorem 8 follows the same line as the one of Theorem 7 except that we have
$\frac{\bigl<\sqrt{n}(\hat{\theta}_{n}-\theta_{0}),u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}=\frac{\bigl<\hat{t}_{n}-\tilde{t}_{n},u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}-\frac{\bigl<\sqrt{n}\nabla\hat{Q}_{n}(\theta_{0}),H_{0}^{-1}u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}.$ (B.4)
By the assumption of the directional asymptotic normality of the gradient, we have
$\frac{\bigl<\sqrt{n}\nabla\hat{Q}_{n}(\theta_{0}),H_{0}^{-1}u\bigr>}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}\to{\cal N}(0,1)$
hence by Slutsky it only remains to show that the first term in the right-hand side equality of (B.4) tends to 0 in probability. This can be proved by using that
$\frac{\left|\bigl<\hat{t}_{n}-\tilde{t}_{n},u\bigr>\right|}{\left\|\hat{V}_{n}^{1/2}H_{0}^{-1}u\right\|_{2}}\leq\mathcal{O}_{\mathbb{P}}(1)\left\|\hat{t}_{n}-\tilde{t}_{n}\right\|_{2},$
together with the argument in the proof of Theorem 7 that shows that $\hat{t}_{n}-\tilde{t}_{n}\overset{p}{\to}0$. ∎
B.2 Proof of Theorem 2
To show a.n. of the polyads estimator, we apply Theorem 8. We therefore need to check the assumptions from this theorem for the choice of loss and risk functions from (20).
First, convexity and the ${\cal C}^{2}$-regularity of $\hat{Q}_{N}$ follow from Lemma 1 as well as for the existence of the polyads estimator. It remains to show the properties of the risk functions $Q_{N}$ (given Assumption 3, we only need to check the uniform continuity of the Hessian), the uniform over compact sets convergence in probability of Hessian matrices and the a.n. of the gradient of the loss function at $\beta_{\star}$. The next two sections are devoted to this latter task. It goes through the a.n. of its Hájek projection.
B.2.1 Hájek projection of the gradient of the loss function at $\beta_{\star}$
We use the notation $\beta\to\ell_{\xi}(\beta)=\ell_{\xi}(Y|X,\beta)$ for all $\xi\in\Xi$, the set of all polyads, and
$U_{N}:=\nabla\hat{Q}_{N}(\beta_{\star})=\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}\nabla\ell_{\xi}(\beta_{\star}).$
The main difficulty in proving a CLT for $U_{N}$ is that it is not a sum of independent variables. However, we prove in this section, as in graham2017econometric; jochmans2018semiparametric, that it can be well approximated (in a sense given in Proposition 2 below) by its Hájek projection - which is a sum of independent variables.
We first show that
$U_{N}^{*}:=\sum_{{\bf i}}{\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{\bf i}]$
is the Hájeck projection of $U_{N}$ onto the linear sub-space $\cal S$} of $\mathbb{R}^{p}$-valued random variables in $L^{2}(\mathbb{P}_{\beta^{*}}[\cdot|X])$ defined by
${\cal S}:=\left\{\sum_{{\bf i}}Z_{{\bf i}}:Z_{{\bf i}}{\mbox{ is }}\sigma(X,Y_{\bf i})-{\mbox{measurable and }}{\mathbb{E}}[\left\|Z_{{\bf i}}\right\|_{2}^{2}|X]<\infty\right\}.$
In other word, we want to show that
$\min\left({\mathbb{E}}_{\beta_{\star}}[\left\|U_{N}-Z\right\|_{2}^{2}|X]:Z\in{\cal S}\right)$ (B.5)
is achieved by $U_{N}^{*}$. Note that the $(\sigma(X,Y_{\bf i}))_{{\bf i}}$ are not independent in general, however, they are independent conditionally on $X$ because all edges weights are independent conditionally on $X$ in model assumption (1). That is the reason why we are considering the Hájeck projection conditionally to $X$. However, because of this conditioning, we cannot directly apply Lemma 11.10 from MR1652247 but we can adapt its proof to our context. We therefore follow the proof strategy of Lemma 11.10 from MR1652247 to show that $U_{N}^{*}$ is solution to (B.5).
First, it is clear that $U_{N}^{*}$ belongs to $\cal S$} (note that $U_{N}$ has a second moment because of Lemma 1 and $m_{\xi}(Y)$ has a second moment). Next, we want to show that $U_{N}-U_{N}^{*}$ is orthogonal to $\cal S$} conditionally on $X$. To prove it, we show that for every $\bf i$} and $Z_{{\bf i}}$ which is $\sigma(X,Y_{\bf i})$-measurable with a second moment, we have
${\mathbb{E}}_{\beta_{\star}}[\bigl<Z_{{\bf i}},U_{N}-U_{N}^{*}\bigr>|X]=0.$
Given that ${\mathbb{E}}_{\beta_{\star}}[\bigl<Z_{{\bf i}},U_{N}-U_{N}^{*}\bigr>|X]={\mathbb{E}}_{\beta_{\star}}[{\mathbb{E}}_{\beta_{\star}}[\bigl<Z_{{\bf i}},U_{N}-U_{N}^{*}\bigr>|X,Y_{\bf i}]|X]$ (see Lemma C.12), it is enough to show that ${\mathbb{E}}_{\beta_{\star}}[\bigl<Z_{{\bf i}},{\mathbb{E}}[(U_{N}-U_{N}^{*})|X,Y_{\bf i}]\bigr>|X]=0$. The latter will be true if we show that
${\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{\bf i}]={\mathbb{E}}_{\beta_{\star}}[U_{N}^{*}|X,Y_{\bf i}].$
Given the definition of $U_{N}^{*}$, we only need to prove that ${\mathbb{E}}_{\beta_{\star}}[{\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{\bf i}]|X,Y_{{\bf i}^{\prime}}]=0$ for all $\bf i$$^{\prime}$ different from $\bf i$}. But for all $\bf i$$^{\prime}$ different from $\bf i$}, $Y_{{\bf i}}$ is independent of $Y_{\bf i}$ conditionally on $X$, hence, by Lemma C.13 and Lemma C.12, we have
${\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{{\bf i}}|]|X,Y_{{\bf i}^{\prime}}\right]={\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{{\bf i}}|]|X\right]={\mathbb{E}}_{\beta_{\star}}[U_{N}|X]=0$
by definition of $\beta_{\star}$. As a consequence, we proved that $U_{N}-U^{*}_{N}$ is orthogonal to $\cal S$} conditionally on $X$. This proved that $U_{N}^{*}$ is the Hájeck projection of $U_{N}$.
Theorem 2 is a directional CLT for the polyads estimator. To prove it we will apply Theorem 8 that requires a directional CLT for $U_{N}$. Our approach is to show that such a directional CLT for $U_{N}$ may be derived from a directional CLT for its Hájek projection. To that end, we need to prove an asymptotic equivalence between the directional projections $\bigl<U_{n},c\bigr$> and $\bigl<U_{n}^{*},c\bigr$> for a given $c\in\mathbb{R}^{p}$. The following result show that such a result holds when the variances of $\bigl<U_{n},c\bigr$> and $\bigl<U_{n}^{*},c\bigr$> are asymptotically equivalent.
Proposition 2.
Let $c\in\mathbb{R}^{p}$. For almost all $X$, the following statement holds:
 $\mbox{ if }\frac{{\mathbb{E}}_{\beta_{\star}}\left[\bigl<U_{N},c\bigr>^{2}|X\right]}{{\mathbb{E}}_{\beta_{\star}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]}\to 1\mbox{ then }\frac{\bigl<U_{N},c\bigr>}{\sqrt{{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}}-\frac{\bigl<U_{N}^{*},c\bigr>}{\sqrt{{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}}\overset{\mathbb{P}_{\beta_{\star}}[\cdot|X]}{\longrightarrow}0$
Proof of Proposition 2.
We provide the proof of Proposition 2 for the sake of completeness. We denote ${\mathbb{E}}_{\beta_{\star}}={\mathbb{E}}$. We first start to show that
${\mathbb{E}}[\bigl<U_{N}-U_{N}^{*},c\bigr>^{2}|X]={\mathbb{E}}[\bigl<U_{N},c\bigr>^{2}|X]-{\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X].$ (B.6)
We have
${\mathbb{E}}\left[\bigl<U_{n},c\bigr>\bigl<U_{N}^{*},c\bigr>|X\right]=\sum_{\bf i}{\mathbb{E}}\left[\bigl<U_{N},c\bigr>{\mathbb{E}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]|X\right]=\sum_{\bf i}{\mathbb{E}}\left[\left({\mathbb{E}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]\right)^{2}|X\right]$
and, since $Y_{\bf i}$ and $Y_{\bf j}$ are independent conditionally on $X$ when ${\bf i}\neq{\bf j}$ and ${\mathbb{E}}[U_{N}|X]=0$, we also have
${\mathbb{E}}\left[\bigl<U_{n}^{*},c\bigr>\bigl<U_{N}^{*},c\bigr>|X\right]=\sum_{{\bf i},{\bf j}}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<U_{N},c\bigr>|X,Y_{\bf j}]{\mathbb{E}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]|X\right]=\sum_{{\bf i}}{\mathbb{E}}\left[\left({\mathbb{E}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]\right)^{2}|X\right].$
Hence, the latter two quantities are equal and so (B.6) follows. We conclude the proof by using a second order Chebyshev’s inequality and by controlling the second order moment with
$\frac{{\mathbb{E}}[\bigl<U_{N}-U_{N}^{*},c\bigr>^{2}|X]}{{\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]}=\frac{{\mathbb{E}}[\bigl<U_{N},c\bigr>^{2}|X]-{\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}{{\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}=1-\frac{{\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}{{\mathbb{E}}[\bigl<U_{N},c\bigr>^{2}|X]}.$
Following Proposition 2, our next step is to find asymptotic equivalents for ${\mathbb{E}}[\bigl<U_{N},c\bigr>^{2}|X]$ and ${\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]$ and show that they are the same. This will show that the condition ${\mathbb{E}}[\bigl<U_{N},c\bigr>^{2}|X]/{\mathbb{E}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]\to 1$ of Proposition 2 holds and so we will be able to apply Proposition 2. We have
${\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>^{2}|X]$ $=\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}{\mathbb{E}}_{\beta_{\star}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X]$
$=\frac{1}{N_{a}^{2}}\sum_{q=0}^{2^{D}}\sum_{\begin{subarray}{c}\xi,\xi^{\prime}\\ |{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=q\end{subarray}}{\mathbb{E}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X\right].$
If $\xi$ and $\xi^{\prime}$ are two polyads with no edge in common then $\nabla\ell_{\xi}(\beta_{\star})$ and $\nabla\ell_{\xi^{\prime}}(\beta_{\star})$ are independent conditionally on $X$ and so we have ${\mathbb{E}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X\right]=0$ in that case because ${\mathbb{E}}_{\beta_{\star}}[\nabla\ell_{\xi}(\beta_{\star})|X]=0$ by definition of $\beta_{\star}$. Moreover, the total number of edges in common to two given polyads lies in $\{0,1,2,2^{2},\cdots,2^{D}\}$. Hence, we have
${\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>^{2}|X]=\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}\Big($ ${\mathbb{E}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X\right]I\left(|{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=1\right)$
$+\sum_{r=1}^{D}{\mathbb{E}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X\right]I\left(|{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=2^{r}\right)\Big).$
Our aim is now to show that, under the condition of Theorem 7, the dominating term in the sum above is the one with all couples of polyads sharing exactly one edge in common. To that end, we introduce the following notation: for all $q\in[2^{D}]$,
$\Delta_{q,N}(X):={\boldsymbol{\mathbb{E}}}_{\beta_{\star}}\left[\bigl<\nabla\ell_{{\boldsymbol{\xi}}}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{{\boldsymbol{\xi}}^{\prime}}(\beta_{\star}),c\bigr>I\left(|{\cal E}({\boldsymbol{\xi}})\cap{\cal E}({\boldsymbol{\xi}}^{\prime})|=q\right)|X\right]$ (B.7)
where $\boldsymbol{\xi}$} and $\boldsymbol{\xi}$$^{\prime}$ are two independent random variables with values in the set of all polyads and uniformly distributed over this set. Using these notation, we have
${\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>^{2}|X]=\frac{N^{2}}{N_{a}^{2}}\sum_{q=1}^{2^{D}}\Delta_{q,N}(X)=T_{1}+\frac{N^{2}}{N_{a}^{2}}\sum_{q=2}^{2^{D}}\Delta_{q,N}(X)$
where
$T_{1}:=\frac{N^{2}}{N_{a}^{2}}\Delta_{1,N}(X)$ (B.8)
is the term that we are proving now to be the dominant term in the variance of $\bigl<U_{N},c\bigr$> as $n$ grows. It follows from Lemma C.11 and Assumption 4 that as $n$ grows to $\infty$,
$\frac{N^{2}}{N_{a}^{2}}\sum_{q=2}^{2^{D}}\Delta_{q,N}(X)\lesssim$
Then, regarding the variance of the Hájek projection in direction $c$, we have
${\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]=\sum_{{\bf i},{\bf i}^{\prime}}{\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{{\bf i}^{\prime}}]\big|X\right].$
Since $Y_{\bf i}$ and $Y_{{\bf i}^{\prime}}$ are independent conditionally to $X$ when ${\bf i}\neq{\bf i}^{\prime}$, we deduce from Lemma C.13 and Lemma C.12 that
${\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}^{\prime}]\big|X\right]={\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]\big|X\right]{\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}^{\prime}]\big|X\right]$
$=\left({\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X]\right)^{2}=0$
and so
${\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]=\sum_{{\bf i}}{\mathbb{E}}_{\beta_{\star}}\left[\left({\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]\right)^{2}\big|X\right].$
For all $\bf i$$\in{\cal I}$, we have
$\left({\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N},c\bigr>|X,Y_{\bf i}]\right)^{2}=\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]$
and so
${\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]=\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}\sum_{{\bf i}}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]|X\right].$
We first note that if $\bf i$} is not an edge of $\xi$ then ${\mathbb{E}}[\nabla\ell_{\xi}(\beta_{\star})|X,Y_{\bf i}]=0$ because in that case $\nabla\ell_{\xi}(\beta_{\star})$ and $Y_{\bf i}$ are independent conditionally to $X$ and we always have ${\mathbb{E}}_{\beta_{\star}}[\nabla\ell_{\xi}(\beta_{\star})|X]=0$ by definition of $\beta_{\star}$. We therefore obtain that
${\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]=\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}\left[\sum_{{\bf i}\in{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]|X\right]\right]$
$=\frac{1}{N_{a}^{2}}\sum_{q=1}^{2^{D}}\sum_{\xi,\xi^{\prime}}\sum_{{\bf i}\in{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]|X\right]I\left(|{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=q\right).$ (B.9)
Moreover, if $\xi$ and $\xi^{\prime}$ share exactly one edge, denoted by $\bf i$}, in common then $\nabla\ell_{\xi}(\beta_{\star})$ and $\nabla\ell_{\xi^{\prime}}(\beta_{\star})$ are independent conditionally to $\sigma(X,Y_{\bf i})$, hence
${\mathbb{E}}_{\beta_{\star}}\left[{\mathbb{E}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}\right]{\mathbb{E}}\left[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}\right]|X\right]={\mathbb{E}}\left[{\mathbb{E}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}\right]|X\right]$
$={\mathbb{E}}\left[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X\right]$
and so for the ’$q=1$’ term in (B.2.1), we get
$\frac{1}{N_{a}^{2}}\sum_{\xi,\xi^{\prime}}\sum_{{\bf i}\in{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]|X\right]I\left(|{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|=1\right)=\frac{N^{2}}{N_{a}^{2}}\Delta_{1,N}(X)=T_{1}$
where $\Delta_{1,N}(X)$ has been introduced in (B.7) and $T_{1}$ in (B.8). This shows that in the variance decomposition of both $\bigl<U_{N},c\bigr$> and $\bigl<U_{N}^{*},c\bigr$> the term for $q=1$ (i.e. the term coming from all couples of polyads sharing exactly one edge in common) is the same given by $T_{1}$. The next step is to show, under Assumption 4, that this term is also the dominant term in ${\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]$ as $n$ grows. It follows from Lemma C.11 that
$\frac{1}{N_{a}^{2}}\sum_{\begin{subarray}{c}\xi,\xi^{\prime}\\ |{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})|\geq 2\end{subarray}}\sum_{{\bf i}\in{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})}{\mathbb{E}}\left[{\mathbb{E}}[\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]{\mathbb{E}}[\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>|X,Y_{\bf i}]|X\right]$
[…]
and so we conclude that $T_{1}$ is the dominating term in the variance of $\bigl<U_{N}^{*},c\bigr$> thanks to Assumption 4.
We conclude that both ${\mathbb{E}}\left[\bigl<U_{N},c\bigr>^{2}|X\right]$ and ${\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]$ are asymptoticaly equivalent to $T_{1}$ and so their ratio tends to $1$ as $n$ grows. Then, it follows from Proposition 2 that, as $n$ tends to infinity,
$\frac{\bigl<U_{N},c\bigr>}{\sqrt{{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}}-\frac{\bigl<U_{N}^{*},c\bigr>}{\sqrt{{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}}\overset{\mathbb{P}_{\beta_{\star}}[\cdot|X]}{\longrightarrow}0.$ (B.10)
B.2.2 Asymptotic normality of the Hájek projection of the gradient
We recall that $U^{*}_{N}$ is the Hájek projection of $U_{N}$ where
$U_{N}^{*}:=\sum_{{\bf i}}{\mathbb{E}}_{\beta_{\star}}[U_{N}|X,Y_{\bf i}]\mbox{ and }U_{N}=\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}\nabla\ell_{\xi}(\beta_{\star}).$
Our aim, in this section, is to show a directional CLT for $U_{N}^{*}$ conditionally to $X$. Together with the asymptotic equivalence proved in (B.10) this will prove a similar directional CLT for $U_{N}$.
We know that if $\bf i$} is not an edge of $\xi$ then ${\mathbb{E}}[\nabla\ell_{\xi}(\beta_{\star})|X,Y_{\bf i}]=0$ by definition of $\beta_{\star}$. As a consequence, we have
$U_{N}^{*}=\frac{1}{{N_{a}}}\sum_{{\bf i}}\sum_{\xi:{\bf i}\in{\cal E}(\xi)}{\mathbb{E}}[\nabla\ell_{\xi}(\beta_{\star})|X,Y_{\bf i}]=\frac{1}{{N_{a}}}\sum_{{\bf i}}\bar{s}_{{\bf i}}$ (B.11)
where, we recall that for all $\bf i$$\in{\cal I}$
$\bar{s}_{{\bf i}}=\sum_{\xi:{\bf i}\in{\cal E}(\xi)}{\mathbb{E}}[\nabla\ell_{\xi}(\beta_{\star})|X,Y_{\bf i}].$
All the point of working on $U_{N}^{*}$ and note directly on $U_{N}$ is that, conditionally on $X$, the $\bar{s}_{{\bf i}}$’s are independent because $\bar{s}_{\bf i}$ is $\sigma(X,Y_{\bf i})$-measurable and the edge weights $Y_{{\bf i}}$’s are independent conditionally on $X$. To prove a.n. for $U_{N}^{*}$ we apply Chatterjee’s result MR2294976 as in graham2017econometric.
To simplify the exposition we denote by $\mathbb{E}$} the expectation under $\mathbb{P}_{\beta_{\star},\theta}$, i.e. ${\mathbb{E}}_{\beta_{\star}}$ and $\alpha_{1,N}=\alpha_{1,N}(X)$. Let $c\in\mathbb{R}^{p}$. We first compute the variance of the Hájek projection using the notation from (B.11): since $\bar{s}_{{\bf i}}$ and $\bar{s}_{{\bf i}^{\prime}}$ are independent conditionally on $X$ when ${\bf i}\neq{\bf i}^{\prime}$ and ${\mathbb{E}}[\bar{s}_{\bf i}|X]=0$, we have
${\mathbb{E}}\left[\bigl<U_{N}^{*},c\bigr>^{2}|X\right]=\frac{1}{N_{a}^{2}}\sum_{{\bf i},{\bf i}^{\prime}}c^{\top}{\mathbb{E}}\left[\bar{s}_{{\bf i}}\bar{s}_{{\bf i}^{\prime}}^{\top}|X\right]c=\frac{1}{{N_{a}}^{2}}\sum_{{\bf i}}c^{\top}{\mathbb{E}}\left[\bar{s}_{{\bf i}}\bar{s}_{{\bf i}}^{\top}|X\right]c=\frac{n\alpha_{1,N}}{{N_{a}}^{2}}c^{\top}\tilde{\Delta}_{N}c$ (B.12)
where $n=\prod_{d}n_{d}$ and
$\tilde{\Delta}_{N}=\frac{1}{n}\sum_{{\bf i}}\tilde{\Delta}_{{\bf i}}\mbox{ for }\tilde{\Delta}_{{\bf i}}=\alpha_{1,N}^{-1}{\mathbb{E}}\left[\bar{s}_{{\bf i}}\bar{s}_{{\bf i}}^{\top}|X\right].$
Therefore, we want to show that for almost all $X$, conditionally on $X$,
$\frac{{N_{a}}}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}\bigl<\frac{1}{\sqrt{n\alpha_{1,N}}}U_{N}^{*},c\bigr>=\frac{1}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}\bigl<\frac{1}{\sqrt{n\alpha_{1,N}}}\sum_{{\bf i}}\bar{s}_{{\bf i}},c\bigr>\overset{d}{\to}{\cal N}(0,1).$
For all $\bf i$}, we set
$R_{\bf i}=\frac{1}{\sqrt{\alpha_{1,N}}}\frac{\bigl<\bar{s}_{\bf i},c\bigr>}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}.$
Our aim is to show that for almost all $X$, conditionally on $X$, $n^{-1/2}\sum_{{\bf i}}R_{\bf i}\overset{d}{\to}{\cal N}(0,1)$ as $n\to+\infty$.
Let us first start with the computation of the conditional expectation and variance of $R_{\bf i}$. Since $\tilde{\Delta}_{N}$ is $X$-measurable, we have
${\mathbb{E}}[R_{{\bf i}}|X]=\frac{1}{\sqrt{\alpha_{1,N}}}\frac{\bigl<{\mathbb{E}}[\bar{s}_{\bf i}|X],c\bigr>}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}=0$
because ${\mathbb{E}}[\bar{s}_{\bf i}|X]=0$ since ${\mathbb{E}}[\nabla\ell_{\xi}(\beta_{\star})|X]=0$ for all polyads $\xi$. To compute the variance of $R_{\bf i}$ conditionally on $X$, we compute its second moment and since $\tilde{\Delta}_{N}$ is $\sigma(X)$-measurable, it only depends on the second moment of $\bigl<\bar{s}_{\bf i},c\bigr$> conditionally on $X$, which is the quantity appearing in the asymptotic covariance matrix from Theorem 2:
${\mathbb{E}}\left[R_{\bf i}^{2}|X\right]=\frac{c^{\top}\tilde{\Delta}_{\bf i}c}{c^{\top}\tilde{\Delta}_{N}c}:=\sigma_{\bf i}^{2}.$
Now, as in MR2294976 and graham2017econometric, we consider independent Gaussian variables with the same first and second moments as the $R_{\bf i}$’s: let $(G_{\bf i})_{\bf i}$ be independent Gaussian variable with mean 0 and variances $(\sigma_{\bf i}^{2})_{\bf i}$. In particular, we observe that, conditionally on $X$, $n^{-1/2}\sum_{\bf i}G_{\bf i}$ is a ${\cal N}(0,1)$ random variable since $n^{-1}\sum_{\bf i}\sigma_{\bf i}^{2}=1$. As a consequence, in order to show that $n^{-1/2}\sum_{{\bf i}}R_{\bf i}\overset{d}{\to}{\cal N}(0,1)$ conditionally on $X$, it is enough to show that for all functions $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$, we have
${\mathbb{E}}\left[f\left(\frac{1}{\sqrt{n}}\sum_{{\bf i}}R_{\bf i}\right)-f\left(\frac{1}{\sqrt{n}}\sum_{{\bf i}}G_{\bf i}\right)\bigg|X\right]\to 0.$ (B.13)
Let $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$. Our aim is to show that (B.13) holds. As in MR2294976 and graham2017econometric, we consider
$Z_{\bf i}=(R_{1},R_{2},\ldots,R_{\bf i},G_{{\bf i}+1},\ldots,G_{n})\mbox{ and }Z_{\bf i}=(R_{1},R_{2},\ldots,R_{{\bf i}-1},0,G_{{\bf i}+1},\ldots,G_{n})$
where $n=\prod_{d}n_{d}=|{\cal I}|$ is the total number of possible edges. Using a second order Taylor approximation of $f$, we get, as $n$ tends to infinity,
[…]
$=\sum_{{\bf i}=1}^{n}f\left(\frac{1}{\sqrt{n}}\bigl<Z_{\bf i}^{0},\mathbb{I}\bigr>+\frac{R_{\bf i}}{\sqrt{n}}\right)-f\left(\frac{1}{\sqrt{n}}\bigl<Z_{{\bf i}}^{0},\mathbb{I}\bigr>+\frac{G_{\bf i}}{\sqrt{n}}\right)$
$=\sum_{{\bf i}=1}^{n}\left(\frac{R_{\bf i}}{\sqrt{n}}-\frac{G_{\bf i}}{\sqrt{n}}\right)f^{\prime}\left(\frac{1}{\sqrt{n}}\bigl<Z_{{\bf i}}^{0},\mathbb{I}\bigr>\right)+\frac{1}{2}\left(\frac{R_{\bf i}^{2}}{n}-\frac{G_{\bf i}^{2}}{n}\right)f^{\prime\prime}\left(\frac{1}{\sqrt{n}}\bigl<Z_{{\bf i}}^{0},\mathbb{I}\bigr>\right)+\mathcal{O}\left(\frac{|R_{\bf i}|^{3}}{n^{3/2}}+\frac{|G_{\bf i}|^{3}}{n^{3/2}}\right)$
where $\mathbb{I}=(1)_{1}^{n}$. Next, we observe that $Z_{\bf i}^{0}$ is independent of $R_{\bf i}$ and $G_{\bf i}$ conditionally on $X$ and since $R_{\bf i}$ and $G_{\bf i}$ have the same first and second moments conditionally on $X$, we obtain that, as $n$ tends to infinity
$\left|{\mathbb{E}}\left[f\left(\frac{1}{\sqrt{n}}\sum_{{\bf i}}R_{\bf i}\right)-f\left(\frac{1}{\sqrt{n}}\sum_{{\bf i}}G_{\bf i}\right)\bigg|X\right]\right|=\mathcal{O}\left(\frac{1}{n^{3/2}}\sum_{\bf i}{\mathbb{E}}\left[|R_{\bf i}|^{3}+|G_{\bf i}|^{3}|X\right]\right).$ (B.14)
Next, we show that the right-hand side from (B.14) goes to 0 as $n$ goes to infinity. First, we observe that for all $\bf i$}, ${\mathbb{E}}[|G_{\bf i}|^{3}|X]\lesssim\left({\mathbb{E}}[|G_{\bf i}|^{2}|X]\right)^{3/2}=\left({\mathbb{E}}[R_{\bf i}^{2}|X]\right)^{3/2}\leq{\mathbb{E}}[|R_{\bf i}|^{3}|X]$. Hence, we only need to show that $(1/n)\sum_{\bf i}{\mathbb{E}}\left[|R_{\bf i}|^{3}|X\right]=o(n^{1/2})$. We have
$\frac{1}{n}\sum_{{\bf i}}{\mathbb{E}}[|R_{\bf i}|^{3}|X]={\mathbb{E}}\left[\frac{\frac{1}{n}\sum_{\bf i}|Z_{\bf i}|^{3}}{\left(\frac{1}{n}\sum_{\bf i}|Z_{\bf i}|^{2}\right)^{3/2}}\bigg|X\right]\mbox{ where }Z_{\bf i}=\bigl<\bar{s}_{\bf i},c\bigr$>.
We note that $Z_{\bf i}$’s are independent conditionally to $X$ because $Z_{\bf i}$ is $\sigma(X,Y_{\bf i})$-measurable and the $Y_{\bf i}$’s are independent conditionally on $X$. Then, we apply Proposition 5 to $X_{{\bf i},n}=Z_{\bf i}$ — thanks to Assumption 6, the condition (C.12) required to apply Proposition 5 is satisfied. It then follows from Proposition 5, for $S_{k,n}=\frac{1}{n}\sum_{\bf i}{\mathbb{E}}[|Z_{\bf i}|^{k}|X]$, that when $n\to+\infty$,
$\frac{1}{n}\sum_{\bf i}{\mathbb{E}}\left[|R_{\bf i}|^{3}|X\right]={\mathbb{E}}\left[\frac{\frac{1}{n}\sum_{\bf i}|Z_{\bf i}|^{3}}{\left(\frac{1}{n}\sum_{\bf i}|Z_{\bf i}|^{2}\right)^{3/2}}\bigg|X\right]=o(\sqrt{n}).$
In the end, we showed that as $n$ goes to infinity,
${\mathbb{E}}\left[f\left(\frac{1}{\sqrt{n}}\sum_{{\bf i}}R_{\bf i}\right)\bigg|X\right]\to{\mathbb{E}}f(g)$ (B.15)
where $g\sim{\cal N}(0,1)$ for all $f\in{\cal C}^{3}$ so that $\left\|f^{(r)}\right\|_{\infty}\leq L$ for all $r=0,1,2,3$. We conclude that for almost all $X$, conditionally on $X$, as $n$ tends to infinity,
$\frac{\bigl<U_{N}^{*},c\bigr>}{\sqrt{{\mathbb{E}}_{\beta_{\star}}[\bigl<U_{N}^{*},c\bigr>^{2}|X]}}=\frac{{N_{a}}}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}\bigl<\frac{1}{\sqrt{n\alpha_{1,N}}}U_{N}^{*},c\bigr>=\frac{1}{\sqrt{n}}\sum_{{\bf i}}R_{\bf i}\overset{d}{\to}{\cal N}(0,1)$
and so, it follows from (B.10) and Slutsky’s lemma that the same result holds for the gradient itself:
$\frac{{N_{a}}}{\sqrt{c^{\top}\tilde{\Delta}_{N}c}}\bigl<\frac{1}{\sqrt{n\alpha_{1,N}}}U_{N},c\bigr>\overset{d}{\to}{\cal N}(0,1)$ (B.16)
where this convergence holds w.r.t. $\mathbb{P}_{\beta_{\star}}[\cdot|X]$.
B.2.3 Uniform convergence of the Hessian over the compact set $B_{2}(\beta_{\star},1)$
The final ingredient needed to apply Theorem 8 is the uniform convergence in probability over the compact set $K:=B_{2}(\beta_{\star},1)=\{\beta:\left\|\beta-\beta_{\star}\right\|_{2}\leq 1\}$ of the Hessian matrices of the loss functions (it is item 3. from Theorem 7 for $\varepsilon_{0}=1/2$). It is the aim of this section to check the uniform convergence (conditionally on $X$) $\sup_{\beta\in K}\left\|\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)\right\|_{op}\overset{p}{\to}\;0$ where, we recall that for all $\beta\in\mathbb{R}^{p}$,
$\hat{Q}_{N}(\beta)=\frac{1}{{N_{a}}}L(Y|X,\beta)\mbox{ and }Q_{N}(\beta)={\mathbb{E}}_{\beta_{\star}}[\hat{Q}_{N}(\beta)|X].$
In graham2017econometric; jochmans2018semiparametric, the authors apply Lemma 2.9 from newey1994large to prove the uniform convergence of the sequence of Hessian matrices of the loss function. However, the latter result requires the existence of a limit risk function that we don’t have. We therefore apply Corollary 2.2 from newey91. We recall this result adapted to our setup (we work with functions with values in $\mathbb{R}^{p\times p}$ unlike newey91).
Proposition 3 (Corollary 2.2 in newey91).
Let $(\hat{H}_{n})_{n}$ (resp. $(H_{n})_{n}$) be a sequence of random (resp. deterministic) functions with values in $\mathbb{R}^{p\times p}$ defined on a compact set $K$ of $\mathbb{R}^{p}$. We assume that:
  • for all $\beta\in K$, $\hat{H}_{n}(\beta)-H_{n}(\beta)\overset{p}{\to}0$;
  • there exists $(B_{n})_{n}$ such that $B_{n}=\mathcal{O}_{p}(1)$ and for all $\beta_{0},\beta_{1}\in K,\left\|\hat{H}_{n}(\beta_{0})-\hat{H}_{n}(\beta_{1})\right\|_{op}\leq B_{n}\left\|\beta_{0}-\beta_{1}\right\|_{2}$;
  • $(H_{n})_{n}$ is equicontinuous.
Then, we have $\sup_{\beta\in K}\left\|\hat{H}_{n}(\beta)-H_{n}(\beta)\right\|_{op}\overset{p}{\to}0$.
Proof.
The proof is a straightforward application of the one-dimensional result given in Corollary 2.2 in newey91 since one can write
[…]
where $B$ is the dual (compact) ball of the operator norm in $\mathbb{R}^{p\times p}$. ∎
The equicontinuity of the family of Hessian matrix of the risk functions is assumed in Assumption 5. We therefore, only have to show the Lipshitz property of the Hessian matrices of the loss functions and their point wise convergence in probability, i.e.:
  • (a)
    for all $\beta_{0},\beta_{1}\in K$, $\left\|\nabla^{2}\hat{Q}_{N}(\beta_{0})-\nabla^{2}\hat{Q}_{N}(\beta_{1})\right\|_{op}=O_{p}(1)\left\|\beta_{1}-\beta_{0}\right\|_{2}$
  • (b)
    for all $\beta\in K$, $\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)\overset{p}{\to}0$
We start with the point-wise convergence in probability over $K$. To prove this result it is enough to show that all $p^{2}$ entries of the matrix $\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)$ tend to zero in probability. Let $\beta\in\mathbb{R}^{p}$ and $(i,j)\in[p]^{2}$. It follows from Lemma 1 that
$\left(\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)\right)_{ij}$ $=\frac{1}{{N_{a}}}\sum_{\xi}\left(\mathbb{V}_{\beta}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]-{\mathbb{E}}_{\beta_{\star}}\left[\mathbb{V}_{\beta}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]\right]\right)(\tilde{X}_{\xi})_{i}(\tilde{X}_{\xi}^{\top})_{j}$
[…]
${\mathbb{E}}_{\beta_{\star}}\left[\left(\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)\right)_{ij}^{2}\bigg|X\right]$
${\mathbb{E}}_{\beta_{\star}}\left[\left(\frac{1}{{N_{a}}}\sum_{\xi}\left(\mathbb{V}_{\beta}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]-{\mathbb{E}}_{\beta_{\star}}\left[\mathbb{V}_{\beta}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]\right]\right)(\tilde{X}_{\xi})_{i}(\tilde{X}_{\xi}^{\top})_{j}\right)^{2}\bigg|X\right]$
$=\frac{1}{{N_{a}}^{2}}\sum_{\begin{subarray}{c}\xi,\xi^{\prime}\in\Xi\\ {\cal E}(\xi)\cap{\cal E}(\xi^{\prime})\neq\emptyset\end{subarray}}{\mathbb{E}}_{\beta_{\star}}\left[\left(V_{\xi}(Y)-{\mathbb{E}}_{\beta_{\star}}V_{\xi}(Y)\right)\left(V_{\xi^{\prime}}(Y)-{\mathbb{E}}_{\beta_{\star}}V_{\xi^{\prime}}(Y)\right)|X\right](\tilde{X}_{\xi})_{i}(\tilde{X}_{\xi}^{\top})_{j}(\tilde{X}_{\xi^{\prime}})_{i}(\tilde{X}_{\xi^{\prime}}^{\top})_{j}$
where we denote for all $y$, $V_{\xi}(y)=\mathbb{V}_{\beta}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(y)\right]$. Let us now take a closer look at the correlation terms in the sum above: let $\xi$ and $\xi^{\prime}$ be two polyads, we have
${\mathbb{E}}_{\beta_{\star}}\left[\left(V_{\xi}(Y)-{\mathbb{E}}_{\beta_{\star}}V_{\xi}(Y)\right)\left(V_{\xi^{\prime}}(Y)-{\mathbb{E}}_{\beta_{\star}}V_{\xi^{\prime}}(Y)\right)|X\right]\leq{\mathbb{E}}_{\beta_{\star}}\left[V_{\xi}(Y)V_{\xi^{\prime}}(Y)|X\right]$
because $V_{\xi}(Y)\geq 0$ a.s.. Next, we have $m_{\xi}(y^{\prime})\leq m_{\xi}(Y)+M_{\xi}(Y)$ for all $y^{\prime}\in\mathcal{O}_{\xi}(Y)$ and so $V_{\xi}(Y)\leq m_{\xi}(Y)+M_{\xi}(Y)$. Hence, it follows from Lemma C.10 that
${\mathbb{E}}_{\beta_{\star}}\left[V_{\xi}(Y)V_{\xi^{\prime}}(Y)|X\right]\leq{\mathbb{E}}_{\beta_{\star}}\left[(m_{\xi}(Y)+M_{\xi}(Y))(m_{\xi^{\prime}}(Y)+M_{\xi^{\prime}}(Y))|X\right]\leq c_{0}\mathbb{P}_{\beta_{\star}}\left[\xi\mbox{ and }\xi^{\prime}\mbox{ are active}|X\right]$ (B.17)
and so under Assumption 2, we have $\nabla^{2}\hat{Q}_{N}(\beta)-\nabla^{2}Q_{N}(\beta)\overset{p}{\to}0$.
Next, we move to the Lipschitz property over $K=B_{2}(\beta_{\star},1)$ of the Hessian matrices of the loss functions from point (a) above. Let $\beta_{0},\beta_{1}\in K$. We compute the third order derivative of the loss functions to prove the Lipschitz property of the Hessian. For all $s\in{\mathbb{N}}^{*},y=(y_{\bf i})_{\bf i}$ and $\beta\in\mathbb{R}^{p}$, we let
$\kappa_{s,\xi}(\beta,y):={\mathbb{E}}_{\beta}[(m_{\xi}(Y^{\prime})-m_{\xi}(y))^{s}|X,Y^{\prime}\in\mathcal{O}_{\xi}(y)]=\frac{\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}r^{s}\exp(a_{r}+r\bigl<\widetilde{X}_{\xi},\beta\bigr>)}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(a_{R}+R\bigl<\widetilde{X}_{\xi},\beta\bigr>)}$
where the last inequality follows from the change of variable $m=r+m_{\xi}(y)$ and the result on the distribution of $m_{\xi}(Y)$ given $X$ and $Y\in\mathcal{O}_{\xi}(y)$ from Lemma 1. Next, it follows from (C.4) that for all $\beta$:
$\nabla^{3}_{\beta}\ell_{\xi}(y|X,\beta)=\left(\kappa_{3}(\beta,y)-3\kappa_{2}(\beta,y)\kappa_{1}(\beta,y)+2\kappa_{1}(\beta,y)^{3}\right)\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}\widetilde{X}_{\xi}^{\top}$
where for any vector $\in\mathbb{R}^{p}, uu$$u^{\top}u^{\top}$ is the linear operator $h\in\mathbb{R}^{p}\to\bigl<u,h\bigr>uu^{\top}\in\mathbb{R}^{p\times p}$. As a consequence, if we denote by $B_{2}^{p}=\{x\in\mathbb{R}^{p}:\left\|x\right\|_{2}=1\}$, we have for all $\beta_{0},\beta_{1}\in\mathbb{R}^{p}$ that
[…]
$\leq\sup_{x\in B_{2}^{p},\bar{\beta}\in[\beta_{0},\beta_{1}]}\frac{1}{{N_{a}}}\sum_{\xi}|\bigl<xx^{\top},\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}\bigr>|\left|\kappa_{3,\xi}(\bar{\beta},Y)-3\kappa_{2,\xi}(\bar{\beta},Y)\kappa_{1,\xi}(\bar{\beta},Y)+2\kappa_{1,\xi}^{3}(\bar{\beta},Y)\right|\left\|\widetilde{X}_{\xi}\right\|_{2}\left\|\beta_{0}-\beta_{1}\right\|_{2}.$
Next, we use that $\bigl<xx^{\top},\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}\bigr>=\bigl<x,\widetilde{X}_{\xi}\bigr>^{2}\leq\left\|\widetilde{X}_{\xi}\right\|_{2}^{2}$ to get
$\left\|\nabla^{2}\hat{Q}_{N}(\beta_{0})-\nabla^{2}\hat{Q}_{N}(\beta_{1})\right\|_{op}\leq\max_{\xi\in\Xi}\left\|\widetilde{X}_{\xi}\right\|_{2}^{3}\max_{\beta\in K}\left[\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}\left|\kappa_{3,\xi}(\beta,Y)-3\kappa_{2,\xi}(\beta,Y)\kappa_{1,\xi}(\beta,Y)+2\kappa_{1,\xi}^{3}(\beta,Y)\right|\right]\left\|\beta_{0}-\beta_{1}\right\|_{2}.$
Next, we show that for all $k=1,2,3$, we have
$\sup_{\beta\in K}\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}|\kappa_{k,\xi}(\beta,Y)|=\mathcal{O}_{\mathbb{P}_{\beta_{\star}}[\cdot|X]}(1).$ (B.18)
It follows from Lemma C.3 that if $Y^{\prime}\in\mathcal{O}_{\xi}(Y)$ then $|m_{\xi}(Y^{\prime})-m_{\xi}(Y)|\leq m_{\xi}(Y)+M_{\xi}(Y)=|\mathcal{O}_{\xi}(Y)|$. As a consequence, we have for all $\beta$ and $\xi$,
$|\kappa_{k,\xi}(\beta,Y)|\leq{\mathbb{E}}_{\beta}[|m_{\xi}(Y^{\prime})-m_{\xi}(Y)|^{k}|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y),Y]\leq|\mathcal{O}_{\xi}(Y)|^{k}$
and so we have
$\sup_{\beta\in K}\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}|\kappa_{k,\xi}(\beta,Y)|\leq\frac{1}{{N_{a}}}\sum_{\xi\in\Xi}|\mathcal{O}_{\xi}(Y)|^{k}$
By Chebyshev, we only need to show that for all $k=1,2,3,\sum_{\xi\in\Xi}{\mathbb{E}}_{\beta_{\star}}[|\mathcal{O}_{\xi}(Y)|^{k}|X]=\mathcal{O}(N_{a})$ as $n$ tends to $\infty$. This result follows from Lemma C.10 since we have
${\mathbb{E}}_{\beta_{\star}}[|\mathcal{O}_{\xi}(Y)|^{k}|X]\leq 2^{k-1}{\mathbb{E}}_{\beta_{\star}}\left[|m_{\xi}(Y)|^{k}+|M_{\xi}(Y)|^{k}|X\right]$
$\leq 2^{k-1}C_{2^{D},k}\left(\mathbb{P}_{\beta_{\star}}[m_{\xi}(Y)\geq 1|X]+\mathbb{P}_{\beta_{\star}}[M_{\xi}(Y)\geq 1|X]\right)\leq 2^{k-1}C_{2^{D},k}\mathbb{P}_{\beta_{\star}}[\xi\mbox{ is active}|X]$
where $C_{2^{D},k}$ is the constant appearing in Lemma C.10. Next, since we assumed the fixed effects and the covariates are uniformly bounded,
$\max_{\xi}\left\|\widetilde{X}_{\xi}\right\|_{2}\leq C_{0}\mbox{ and }c_{0}\leq\lambda_{\bf i}\leq C_{0},\forall{\bf i}\in{\cal I}$ (B.19)
we obtain the Lipschitz property of the Hessian matrices of the loss functions.
B.2.4 Final step to the proof of Theorem 2: apply Theorem 8
The statements in this section hold for almost all $X$, contionally on $X$. Both convergence in distribution and in probability are therefore given w.r.t. to the probability distribution $\mathbb{P}_{\beta_{\star}}$ conditionally on $X$ that we denote by $\mathbb{P}_{\beta_{\star}}[\cdot|X]$. Let us now gather all the results we obtained previously and apply Theorem 8. Let $c\in\mathbb{R}^{p}$.
We obtained in (B.16) (applied to $c=\Gamma_{0}^{-1}c$) that
$\frac{{N_{a}}}{\sqrt{n\alpha_{1,N}}}\frac{\bigl<\nabla\hat{Q}_{N}(\beta_{\star}),\Gamma_{0}^{-1}c\bigr>}{\sqrt{c^{\top}\Gamma_{0}^{-1}\tilde{\Delta}_{N}\Gamma_{0}^{-1}c}}\overset{d}{\to}{\cal N}(0,1).$ (B.20)
As a consequence, the directional CLT for the gradient granted in item $4^{\prime}$ of Theorem 8 is satisfied with $\hat{V}_{n}$ defined as the deterministic matrix
$\hat{V}_{n}:=\frac{n\alpha_{1,N}}{{N_{a}}}\tilde{\Delta}_{N}$
and where $N_{a}$ plays the role of the number of data (called $n$ in Theorem 8).
In order to apply Theorem 8, we need to check that there exists some absolute constant $c_{0}>0$ and $n_{0}$ such that for all $n\geq n_{0}$,
$\left\|\hat{V}_{n}^{1/2}\Gamma_{0}^{-1}c\right\|_{2}=\sqrt{\frac{n\alpha_{1,N}}{{N_{a}}}}\left\|\tilde{\Delta}_{N}^{1/2}\Gamma_{0}^{-1}c\right\|_{2}\geq c_{0}.$ (B.21)
In (B.12), we proved that
${\mathbb{E}}\left[\bigl<U_{N}^{*},\Gamma_{0}^{-1}c\bigr>^{2}|X\right]=\frac{n\alpha_{1,N}}{{N_{a}}^{2}}\left\|\tilde{\Delta}_{N}^{1/2}\Gamma_{0}^{-1}c\right\|_{2}^{2}=\frac{1}{N_{a}}\left\|\hat{V}_{n}^{1/2}\Gamma_{0}^{-1}c\right\|_{2}^{2}$
and we showed in Section B.2.1 that ${\mathbb{E}}\left[\bigl<U_{N}^{*},\Gamma_{0}^{-1}c\bigr>^{2}|X\right]$ is asymptotically equivalent to $T_{1}$ defined in (B.8) as $T_{1}=N^{2}\Delta_{1,N}(X)/N_{a}^{2}$ where $\Delta_{1,N}$ is defined in (B.7) where $c$ is taken equal to $\Gamma_{0}^{-1}c$ in this equation. Finally, under Assumption 4, we have $N_{a}T_{1}$ which is lower bounded by an absolute constant for $n$ large enough and so (B.21) follows.
We apply Theorem 8 to the convex (random) loss functions $\hat{Q}_{N}:\beta\in\mathbb{R}^{p}\to(1/{N_{a}})L(Y|X,\beta)$ and its associated risk function $\beta\to Q_{N}(\beta)={\mathbb{E}}_{\beta_{\star}}[\hat{Q}_{N}(\beta)|X]$.
It follows from Lemma 1 that for $V_{\xi}(y)={\mathbb{V}}_{\beta_{\star}}\left(m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(y)\right)$, the Hessian of the risk function is
$\nabla^{2}Q_{N}(\beta)=\frac{1}{N_{a}}\sum_{\xi\in\Xi}{\mathbb{E}}_{\beta_{\star}}\left[V_{\xi}(Y)|X\right]\tilde{X}_{\xi}\tilde{X}_{\xi}^{\top}.$
Hence, the risk function is convex and twice differentiable. It follows from Assumption 3 that $\beta_{\star}$ is the unique minimum of $Q_{N}$. Therefore, Assumption 3 together with Assumption 5 shows that the first item of Theorem 7 is satisfied. The second item of Theorem 7 follows from Lemma 1 and Assumption 3. The third item of Theorem 7 has been proved in Section B.2.3. Finally item $4^{\prime}$ from Theorem 8 was proved right above. It follows from Theorem 8 that
$\frac{N_{a}}{\sqrt{n\alpha_{1,N}}}\frac{\bigl<\hat{\beta}_{\Xi}-\beta_{\star},c\bigr>}{\sqrt{c^{\top}\Gamma_{0}^{-1}\tilde{\Delta}_{N}\Gamma_{0}^{-1}c}}=\frac{\bigl<\sqrt{N_{a}}\left(\hat{\beta}_{\Xi}-\beta_{\star}\right),c\bigr>}{\left\|V_{n}^{1/2}\Gamma_{0}^{-1}c\right\|_{2}}\overset{d}{\to}{\cal N}(0,1).$
Appendix C Appendix
In this section, we collect several technical tools used to prove the consistency result from Theorem 1 and the asymptotic normality of Theorem 2.
C.1 Characterization of degree-preserving transformations
The aim of this subsection is to proof the counterpart of Proposition 1, which assumes $\mathcal{G}=\mathcal{G}^{\max}$. We need to show that if two graphs $y,y^{\prime}\in\mathbb{Z}^{\mathcal{I}}$ satisfy $\delta(y)=\delta(y^{\prime})$, then it is possible to find a finite number $m$, polyads $\xi_{1},\dots,\xi_{m}$ and integers $r_{1},\dots,r_{m}$ such that
$y^{\prime}=T_{\xi_{1}}^{r_{1}}\circ\cdots\circ T_{\xi_{m}}^{r_{m}}(y).$
We will prove by induction on both the dimension $D$ and the component sizes $n_{1},\dots,n_{D}$. We start with the following lemma, that is the base case:
Lemma C.1.
If two graphs $y\neq y^{\prime}\in\ \mathbb{Z}^{\mathcal{I}}$ with $n_{1}=n_{2}=\dots=n_{D}=2$ satisfy $\delta(y)=\delta(y^{\prime})$, then there exist a polyad $\xi$ and an integer $r$ such that $y^{\prime}=T_{\xi}^{r}(y)$.
Proof.
Let $r=y^{\prime}_{1\dots 1}-y_{1\dots 1}$. In a graph with $n_{1}=n_{2}=\dots=n_{D}=2$ there are only $2^{D}$ polyads. Take
$\xi=\begin{pmatrix}1&1&\cdots&1\\ 2&2&\cdots&2\\ \end{pmatrix}$.
We now check that $y^{\prime}=T_{\xi}^{r}(y)$. First, notice that $T_{\xi}^{r}(y)_{1\dots 1}=y_{1\dots 1}+s_{\xi}(1\dots 1)r=y_{1\dots 1}+r=y^{\prime}_{1\dots 1}$. Now we look at ${\bf i}=1\dots 12$ (i.e., the last index is $2$). We know that $T_{\xi}^{r}$ preserves degrees. Looking at the degree associated with $i_{1}=\dots=i_{D-1}=1$ we have
$y^{\prime}_{1\dots 11}+y^{\prime}_{1\dots 12}=y_{1\dots 11}+y_{1\dots 12}=T_{\xi}^{r}(y)_{1\dots 11}+T_{\xi}^{r}(y)_{1\dots 12}=y^{\prime}_{1\dots 11}+T_{\xi}^{r}(y)_{1\dots 12}$
and so $T_{\xi}^{r}(y)_{1\dots 12}=y^{\prime}_{1\dots 12}$. The same argument holds to proof the equalities for all $\bf i$}’s that differ from $1\dots 1$ in one index. Once those are proved they can be used to ensure equality for the $\bf i$}’s that differ in two indices from $1\dots 1$, but only by one index from the already shown. And so on, until we equality is proved for $2\dots 2$. ∎
Now we show that the proposition holds for any sizes $n_{1},n_{2}\geq 2$ when $D=2$. We start noticing that if $n_{1}=n_{2}=2$ it is proven by the previous lemma. By induction we assume it is proven for $n_{1},n_{2}\geq 2$ and we show for $n_{1}+1,n_{2}$ (the case $n_{1},n_{2}+1$ follows permuting the order of the indices). The strategy will be to make $y_{\bf i}=y^{\prime}_{\bf i}$ for all ${\bf i}=(i_{1},i_{2})$ such that $i_{1}=n_{1}+1$, once this is done we can get back to the induction assumption since the degrees of the $n_{1}\times n_{1}$ subgraphs obtained removing the indices with $i_{1}=n_{1}+1$ from $y^{\prime}$ and from the transformed $y$ must be the same. Let $i_{2}\in[n_{2}]$, $i_{2}\neq 1$, define the polyad
$\xi=\begin{pmatrix}n_{1}+1&i_{2}\\ 1&1\\ \end{pmatrix}$,
notice that it can be applied $y^{\prime}_{n_{1}+1,i_{2}}-y_{n_{1}+1,i_{2}}$ times to make the index ${\bf i}=(n_{1}+1,i_{2})$ of the transformed $y$ match the one of $y^{\prime}$. Let $T$ be the composition of these operations for $i_{2}=2,\dots,n_{2}$. It remains to show that $y^{\prime}_{n_{1}+1,1}=T(y)_{n_{1}+1,1}$. The degree equality associated with $i_{1}=n_{1}+1$ yields
$y^{\prime}_{n_{1}+1,1}+\sum_{i_{2}=1}^{n_{2}}y^{\prime}_{n_{1}+1,i_{2}}=y_{n_{1}+1,1}+\sum_{i_{2}=1}^{n_{2}}y_{n_{1}+1,i_{2}}=T(y)_{n_{1}+1,1}+\sum_{i_{2}=1}^{n_{2}}T(y)_{n_{1}+1,i_{2}}=T(y)_{n_{1}+1,1}+\sum_{i_{2}=1}^{n_{2}}y^{\prime}_{n_{1}+1,i_{2}}$
and so $y^{\prime}_{n_{1}+1,1}=T(y)_{n_{1}+1,1}$. We have thus proven the induction step for $D=2$.
Finally we move to $D>2$. The base case with $n_{1}=\cdots=n_{D}=2$ already given by the Lemma. By induction assume that it holds for $n_{1},\dots,n_{D}\geq 2$ and for dimension D1D-1 too. We now show it holds for $n_{1}+1,n_{2},\dots,n_{D}$. Notice that $\{y_{\bf i}:i_{1}=n_{1}+1\}$ and $\{y^{\prime}_{\bf i}:i_{1}=n_{1}+1\}$ are $n_{2}\times\dots\times n_{D}$ graphs and, since the degrees of $y$ are the same as the ones of $y^{\prime}$, they have the same degrees. For example, the degree associated with $(i_{2},\dots,i_{D-1})$ in the subgraphs is the one previously associated with $(n_{1}+1,i_{2},\dots,i_{D-1})$. Therefore, there are D1D-1-dimensional polyads that can turn one subgraph into the other. To obtain $D$-dimensional polyads from these we can simply concatenate with the entry $i_{1}=n_{1}+1$ and $i_{1}^{\prime}=1$. This finishes the proof.
C.2 General results on orbits, $m_{\xi}$ and $M_{\xi}$ and the proof of Lemma 1
We start with two simple observations on the orbits that justifies its name: an orbit is an equivalent class.
Lemma C.2.
Let $\xi$ be a polyad. For all $y,z\in{\mathbb{N}}^{{\cal I}}$, the following are equivalent:
  • (a)
    $y\in\mathcal{O}_{\xi}(z)$
  • (b)
    $\mathcal{O}_{\xi}(y)=\mathcal{O}_{\xi}(z)$.
Lemma C.3.
Let $y,y^{\prime}\in{\mathbb{N}}^{{\cal I}}$ and assume that $y^{\prime}\in\mathcal{O}_{\xi}(y)$. Let $m_{\xi}(y)\leq r\leq M_{\xi}(y)$ then the following are equivalent:
  • (i)
    $y^{\prime}=y^{r}$
  • (ii)
    $m_{\xi}(y^{\prime})=m_{\xi}(y)+r$
  • (iii)
    $M_{\xi}(y^{r})=M_{\xi}(y)-r$.
Proposition 4.
(the diff-in-diff property of $s_{\xi}$) Let $g$ be an ordered tuple of numbers in $[D]$ such that $g\neq\emptyset$ and $g\neq(1,2,\ldots,D)$. We have for all $(i_{d})_{d\in g}\in{\cal I}_{g}$,
$\sum_{{\bf i}=(i_{d})_{d=1}^{D}:(i_{d})_{d\in\bar{g}}\in{\cal I}_{\bar{g}}}s_{\xi}({\bf i})=0$
where ${\cal I}_{g}=\prod_{d\in g}[n_{d}]$ and ${\cal I}_{\bar{g}}=\prod_{d\notin g}[n_{d}]$ [^11] Here, we identify the tuple $g$ with the set of elements made of the coordinates in $g$.. In particular, for all $\theta=\left(\theta_{g({\bf i})}^{g}\in\mathbb{R}:{\bf i}\in\mathcal{I},g\in\mathcal{G}\right)$ and $g\in{\cal G}$, we have
$\sum_{{\bf i}}s_{\xi}({\bf i})\theta_{g({\bf i})}^{g}=\sum_{{\bf i}:s_{\xi}({\bf i})\neq 0}s_{\xi}({\bf i})\theta_{g({\bf i})}^{g}=0.$ (C.1)
Proof.
To verify the first equality, let $\bf i$$^{\prime}$ be fixed such that $s_{\xi}({\bf i}^{\prime})\neq 0$. There are $2^{D-|g|}$ choices of $\bf i$} such that $g({\bf i})=g({\bf i}^{\prime})$ and $s_{\xi}({\bf i})=0$, half of these choices have $s_{\xi}({\bf i})=1$ and the other half $s_{\xi}({\bf i})=-1$ since it suffices to flip one index $d\not\in g$ to obtain a bijection between positive and negative signs. The second inequality is a direct consequence of the first property. ∎
Remark 5.
The diff-in-diff property also holds if we define polyads with equal nodes for some axis $d\in[D]$, i.e. we may not assume that $j_{d}\neq j_{d}^{\prime}$ for all $d\in[D]$ - the minimal requirement is to have at least $2$ axis with two different nodes – i.e. a tetrads. This may be useful when the number of polyads is too large regarding computational cost.
Next we prove Lemma 1.
Proof.
Let $\xi$ be a polyad, $\beta\in\mathbb{R}^{p}$ and $y=(y_{\bf i})_{{\bf i}}$. Let $r\in\{-m_{\xi}(y),\ldots,M_{\xi}(y)\}$. It follows from the diff-in-diff property of the sign function from Proposition 4 that
$\mathbb{P}_{\beta}[Y=y^{r}|X,Y\in\mathcal{O}_{\xi}(y)]=\frac{\mathbb{P}[Y=y^{r}|X]}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\mathbb{P}[Y=y^{R}|X]}=\frac{\Pi_{{\bf i}\in{\cal E}(\xi)}\exp(-\lambda_{\bf i})\lambda_{\bf i}^{y_{\bf i}^{r}}/y_{\bf i}^{r}!}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\Pi_{{\bf i}\in{\cal E}(\xi)}\exp(-\lambda_{\bf i})\lambda_{\bf i}^{y_{\bf i}^{R}}/y_{\bf i}^{R}!}$
$=\left[\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\Pi_{{\bf i}\in{\cal E}(\xi)}\frac{y_{\bf i}^{r}!}{y_{\bf i}^{R}!}\lambda_{\bf i}^{y_{\bf i}^{R}-y_{\bf i}^{r}}\right]^{-1}=\left[\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(-a_{r})\exp(a_{R})\exp[(R-r)\bigl<\tilde{X}_{\xi},\beta\bigr>]\right]^{-1}$
$=\frac{\exp(a_{r}+r\bigl<\widetilde{X}_{\xi},\beta\bigr>)}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(a_{R}+R\bigl<\widetilde{X}_{\xi},\beta\bigr>)}$ (C.2)
where $a_{R}=\ln(\Pi_{{\bf i}\in{\cal E}(\xi)}y_{\bf i}!/y_{\bf i}^{R}!)$ for all $R\in\{-m_{\xi}(y),\ldots,M_{\xi}(y)\}$. Next, let $m\in\{0,\ldots,m_{\xi}(y)+M_{\xi}(y)\}$. We observe that if $Y\in\mathcal{O}_{\xi}(y)$ then $m_{\xi}(Y)=m$ iff $Y=y^{m-m_{\xi}(y)}$. As a consequence, it follows from (C.2) that
$\ln\mathbb{P}_{\beta}(m_{\xi}(Y)=m|X,Y\in\mathcal{O}_{\xi}(y))=\ln\mathbb{P}_{\beta}(Y=y^{m-m_{\xi}(y)}|X,Y\in\mathcal{O}_{\xi}(y))=a_{m-m_{\xi}(y)}+(m-m_{\xi}(y))\bigl<\widetilde{X}_{\xi},\beta\bigr>-\ell_{\xi}(y|X,\beta)$
where we used that
[…]
This proves the last result of Lemma 1 regarding the probability distribution of $m_{\xi}(Y)$ given $Y\in\mathcal{O}_{\xi}(y)$ and $X$.
The gradient of $\beta\to\ell_{\xi}(y|X,\beta)$ can be derived from (C.3) and (C.2):
$\nabla_{\beta}\ell_{\xi}(y|X,\beta)=\frac{\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}r\exp(a_{r}+r\bigl<\widetilde{X}_{\xi},\beta\bigr>)}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(a_{R}+R\bigl<\widetilde{X}_{\xi},\beta\bigr>)}\widetilde{X}_{\xi}=\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}r\mathbb{P}_{\beta}[Y=y^{r}|X,Y\in\mathcal{O}_{\xi}(y)]\widetilde{X}_{\xi}$
$=\sum_{m=0}^{m_{\xi}(y)+M_{\xi}(y)}(m-m_{\xi}(y))\mathbb{P}_{\beta}[m_{\xi}(Y)=m|X,Y\in\mathcal{O}_{\xi}(y)]\widetilde{X}_{\xi}=\left(\operatorname{\mathbb{E}}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]-m_{\xi}(y)\right)\widetilde{X}_{\xi}$
where we used the fact that if $Y\in\mathcal{O}_{\xi}(y)$ then $Y=y^{r}$ iff $m_{\xi}(Y)=r+m_{\xi}(y)$. Finally, we derive the formula for the Hessian of $\beta\to\ell_{\xi}(y|X,\beta)$:
$\nabla^{2}_{\beta}\ell_{\xi}(y|X,\beta)=\left[\frac{\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}r^{2}\exp(a_{r}+r\bigl<\widetilde{X}_{\xi},\beta\bigr>)}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(a_{R}+R\bigl<\widetilde{X}_{\xi},\beta\bigr>)}-\left(\frac{\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}r\exp(a_{r}+r\bigl<\widetilde{X}_{\xi},\beta\bigr>)}{\sum_{R=-m_{\xi}(y)}^{M_{\xi}(y)}\exp(a_{R}+R\bigl<\widetilde{X}_{\xi},\beta\bigr>)}\right)^{2}\right]\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{\top}$ (C.4)
$=\mathbb{V}_{\beta}\left[m_{\xi}(Y)|X,Y\in\mathcal{O}_{\xi}(y)\right]\widetilde{X}_{\xi}\widetilde{X}_{\xi}^{T}$
where we used again that if $Y\in\mathcal{O}_{\xi}(y)$ then $Y=y^{r}$ iff $m_{\xi}(Y)=r+m_{\xi}(y)$ and that $m_{\xi}(Y)$ has the same variance as $m_{\xi}(Y)+m_{\xi}(y)$.
To prove similar results for the gradient and the Hessian in terms of $M_{\xi}(Y)$ we also rely on Lemma C.3 and use the same arguments as above. ∎
C.3 Proof of Lemma A.1
Proof of Lemma A.1:.
Denote by $B_{2}=\{\theta\in\mathbb{R}^{d}:\left\|\theta\right\|_{2}\leq 1\}$ the unit ball with respect to (w.r.t.) the $\ell_{2}^{d}$-norm. Let $\varepsilon>0$ be such that $\theta_{0}+3\varepsilon B_{2}\subset\Theta$. We want to show that with probability approaching $1$ (w.p.a.1), $\hat{\theta}_{n}\in\theta_{0}+\varepsilon B_{2}$. Let $\tilde{\theta}_{n}\in\operatorname*{arg\,min}_{\theta\in\theta_{0}+2\varepsilon B_{2}}\hat{Q}_{n}(\theta)$. Since $\theta_{0}+2\varepsilon B_{2}$ is a compact set, it follows from Lemma C.5 (applied to $\Theta=\theta_{0}+3\varepsilon\overset{\circ}{B_{2}}\;$ and $K=\theta_{0}+2\varepsilon B_{2}$) that the following uniform convergence result holds
$\sup_{\theta\in\theta_{0}+2\varepsilon B_{2}}|\hat{Q}_{n}(\theta)-Q_{n}(\theta)|\overset{p}{\to}0$
and so, by Lemma C.4 (applied to $\Theta=\theta_{0}+2\varepsilon B_{2}$), that $\tilde{\theta}_{n}\overset{p}{\to}\theta_{0}$. In particular, wpa1, $\tilde{\theta}_{n}\in\theta_{0}+\varepsilon B_{2}$. Let us now place ourselves on the event $\tilde{\theta}_{n}\in\theta_{0}+\varepsilon B_{2}$ and let $\theta\notin\theta_{0}+2\varepsilon B_{2}$. There exists $\theta_{1}\in\theta_{0}+2\varepsilon S_{2}$ - where $S_{2}=\{\theta\in\mathbb{R}^{d}:\left\|\theta\right\|_{1}=1\}$ is the unit sphere of the $\ell_{2}^{d}$-norm - and $\lambda\geq 1$ such that $\theta=\tilde{\theta}_{n}+\lambda(\theta_{1}-\tilde{\theta}_{n})$. By convexity of $\hat{Q}_{n}$, we have
$\hat{Q}_{n}(\theta)-\hat{Q}_{n}(\tilde{\theta}_{n})\geq\lambda(\hat{Q}_{n}(\theta_{1})-\hat{Q}_{n}(\tilde{\theta}_{n}))\geq 0$
where the last inequality is due to the fact that $\theta_{1}\in\theta_{0}+2\varepsilon S_{2}\subset\theta_{0}+2\varepsilon B_{2}$ and $\tilde{\theta}_{n}$ minimizes $\hat{Q}_{n}$ over $\theta_{0}+2\varepsilon B_{2}$. As a consequence, $\tilde{\theta}_{n}$ minimizes $\hat{Q}_{n}$ over $\Theta$ and by the uniqueness of $\hat{\theta}_{n}$, we have $\hat{\theta}_{n}=\tilde{\theta}_{n}$. This concludes the proof since $\tilde{\theta}_{n}$ is a consistent estimator of $\theta_{0}$. ∎
The proof of Lemma A.1 provided above is based on three ingredients: 1) convexity of $\hat{Q}_{n}$, 2) the uniform convergence in probability result over the compact set $\theta_{0}+2\varepsilon B_{2}$ derived from Lemma C.5 and, 3) the consistency result over the compact model $\theta_{0}+2\varepsilon B_{2}$ that follows from Lemma C.4. We start with the proof of the latter result which is an adaptation of Theorem 2.1 from newey1994large.
Lemma C.4.
Let $\Theta$ be a non empty and convex set in $\mathbb{R}^{d}$. Let $(\hat{Q}_{n})_{n}$ be a sequence of random functions and $(Q_{n})_{n}$ be a sequence of (deterministic) convex functions all defined on $\Theta$. We assume that:
  • (a)
    with probability approaching $1$, there exists $\hat{\theta}_{n}\in\Theta$ minimizing $\hat{Q}_{n}$ over $\Theta$,
  • (b)
    there exists $\theta_{0}\in\Theta$ and $\varepsilon_{0}>0$ such that $\theta_{0}+\varepsilon_{0}B_{2}\subset\Theta$ and for all $0<\varepsilon\leq\varepsilon_{0}$ there exists $\eta>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and all $\theta\in\Theta$, if $\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon$ then $Q_{n}(\theta)-Q_{n}(\theta_{0})\geq\eta$,
  • (c)
    $\sup_{\theta\in\Theta}|\hat{Q}_{n}(\theta)-Q_{n}(\theta)|\overset{p}{\to}0$.
Then $\hat{\theta}_{n}\overset{p}{\to}\theta_{0}$.
Proof of Lemma C.4.
Let $0<\varepsilon\leq\varepsilon_{0}$ and denote by $\bar{B}=\theta_{0}+\varepsilon B_{2}$ the closed $\ell_{2}^{d}$-ball centered at $\theta_{0}$ with radius $\varepsilon$. We want to show that with probability approaching $1$, $\hat{\theta}_{n}\in\bar{B}$. To that end, it is enough to show that for $n$ large enough we have $Q_{n}(\hat{\theta}_{n})<\inf_{\theta\notin\bar{B}}Q_{n}(\theta)$.
We first show that thanks to the convexity of the $Q_{n}$’s we have $\inf_{\theta\notin\bar{B}}Q_{n}(\theta)=\min_{\theta:\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon}Q_{n}(\theta)$. First, it follows from convexity and (b) that $\theta_{0}$ is the unique global minimizer of $Q_{n}$ over $\Theta$ for all $n\geq n_{0}$. Second, if $\theta\notin\bar{B}$, there exists $\lambda\geq 1$ and $\theta_{1}\in\theta_{0}+\varepsilon S_{2}$ (where $S_{2}$ is the unit $\ell_{2}^{d}$-sphere) such that $\theta=\theta_{0}+\lambda(\theta_{1}-\theta_{0})$ and, from the convexity of $Q_{n}$, we have
$Q_{n}(\theta)-Q_{n}(\theta_{0})\geq\lambda(Q_{n}(\theta_{1})-Q_{n}(\theta_{0}))\geq Q_{n}(\theta_{1})-Q_{n}(\theta_{0})$
where the last inequality follows because $\theta_{0}$ is a global minimizer of $Q_{n}$ and $\lambda\geq 1$. As a consequence, $Q_{n}(\theta)\geq Q_{n}(\theta_{1})$ and so, by continuity of $Q_{n}$, we obtain that $\inf_{\theta\notin\bar{B}}Q_{n}(\theta)=\min_{\theta:\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon}Q_{n}(\theta)$.
It follows from the property (b) of $(Q_{n})_{n}$ that there exists $n_{0}$ such that $\eta>0$ where
$\eta:=\min_{n\geq n_{0}}\min_{\theta:\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon}Q_{n}(\theta)-Q_{n}(\theta_{0}).$
As a consequence, with probability approaching $1$, the following holds from uniform convergence:
$Q_{n}(\hat{\theta}_{n})\leq\hat{Q}_{n}(\hat{\theta}_{n})+\frac{\eta}{3}<\hat{Q}_{n}(\theta_{0})+\frac{2\eta}{3}\leq Q_{n}(\theta_{0})+\eta\geq\min_{\theta:\left\|\theta-\theta_{0}\right\|_{2}=\varepsilon}Q_{n}(\theta)=\inf_{\theta\notin\bar{B}}Q_{n}(\theta).$
The next result shows that uniform convergence in probability over a compact follows from pointwise convergence in probability thanks to the convexity assumption.
Lemma C.5.
Let $\Theta$ be a non empty, open and convex set in $\mathbb{R}^{d}$. Let $(\hat{Q}_{n})_{n}$ be a sequence of convex random functions and $(Q_{n})_{n}$ be a sequence of (deterministic) convex functions all defined on $\Theta$. We assume that:
  1. 1.
    for every $\theta\in\Theta$ there exists $L>0$ such that $\sup_{n}Q_{n}(\theta)\leq L$,
  2. 2.
    there exists $L_{0}\in\mathbb{R}$ and $\theta_{1}\in\Theta$ such that $\inf_{n}Q_{n}(\theta_{1})\geq L_{0}$.
  3. 3.
    for all $\theta\in\Theta$, as $n$ tends to infinity, $\hat{Q}_{n}(\theta)-Q_{n}(\theta)\overset{p}{\to}0$.
Then, for any compact set $K$ in $\Theta$,
$\sup_{\theta\in K}|\hat{Q}_{n}(\theta)-Q_{n}(\theta)|\overset{p}{\to}0.$
Proof of Lemma C.5.
Let $K$ be a compact set in $\Theta$. To show the uniform convergence in probability over the compact set $K$, it is enough to show that for any increasing sequence $(\phi_{n})_{n}$ of integers we can extract a sub-sequence $(\psi_{n})_{n}\subset(\phi_{n})_{n}$ along which $\sup_{\theta\in K}|\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)|$ tends to 0 almost surely.
Let $\Theta^{\prime}$ be a countable dense subset of $\Theta$. According to the diagonalization argument from Andersen1982 (recalled in Lemma C.8 below), there exists a sub-sequence $(\psi_{n})_{n}$ of $(\phi_{n})_{n}$ such that for all $\theta\in\Theta^{\prime},|\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)|\overset{a.s.}{\to}0.$
Next it follows from Lemma C.6 that almost surely for $n$ large enough, $\hat{Q}_{\psi_{n}}-Q_{\psi_{n}}$ is Lipchitz on the compact set $K$. Then, we conclude with Lemma C.7 that $\sup_{\theta\in K}|\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)|$ tends to 0 almost surely. This is true for any increasing sequence $(\phi_{n})_{n}$ and so this shows the uniform over $\Theta$ convergence in probability. ∎
The proof of Lemma A.1 requires to revisit classical results from convex analysis and asymptotic statistics. We first start with Theorem 10.6 and Theorem 10.8 from rockafellar-1970a that need to be adapted to our setup.
Lemma C.6.
(Adapted from Theorem 10.6 in rockafellar-1970a) Let $\Theta$ be a non empty, open and convex set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of convex functions defined on $\Theta$. We assume that there exists $\Theta^{\prime}\subset\Theta$ such that $\Theta\subset{\rm conv}(\bar{\Theta}^{\prime})$ and
  • for all $\theta\in\Theta^{\prime}$, $(f_{n}-g_{n})(\theta)\to 0$,
  • for all $\theta\in\Theta^{\prime}$, there exists $L>0$ such that $\sup_{n}g_{n}(\theta)\leq L$,
  • there exists $L_{0}\in\mathbb{R}$ and $\theta_{1}\in\Theta$ such that $\inf_{n}g_{n}(\theta_{1})\geq L_{0}$.
Then, for any non empty compact set $K$ in $\Theta$, there exists $L>0$ such that for $n$ large enough, we have for all $x,y\in K$,
$|(f_{n}-g_{n})(x)-(f_{n}-g_{n})(y)|\leq L\left\|x-y\right\|_{2}.$
Proof.
Since $(g_{n})_{n}$ is a sequence of convex functions satisfying (a) and (b) from Theorem 10.6 in rockafellar-1970a, it follows from the latter theorem that $(g_{n})_{n}$ is equi-Lipschitzian relative to $K$. Next, since $(f_{n}-g_{n})(\theta)\to 0$ for all $\theta\in\Theta^{\prime}$ and $(g_{n}(\theta))_{n}$ is uniformly (in $n$) bounded from above for all $\theta\in\Theta^{\prime}$ and there is some $\theta_{1}$ for which $(g_{n}(\theta_{1}))$ is bounded from below, then for all $\theta\in\Theta^{\prime}$, $(f_{n}(\theta))_{n}$ is also bounded from above and $(g_{n}(\theta_{1}))$ is bounded from below. Hence, it follows from Theorem 10.6 in rockafellar-1970a that $(f_{n})_{n}$ is equi-Lipschitzian relative to $K$. We conclude that the sequence of differences $(f_{n}-g_{n})_{n}$ is also equi-Lipschitzian relative to $K$. ∎
Lemma C.7.
(adapted from Theorem 10.8 from rockafellar-1970a) Let $\Theta$ be a non empty set in $\mathbb{R}^{d}$. Let $(f_{n})_{n}$ and $(g_{n})_{n}$ be two sequences of functions defined on $\Theta$. We assume that for all compact set $K$ in $\Theta$, there exists $L>0$ such that for all $n$ large enough and all $x,y\in K$,
$|(f_{n}-g_{n})(x)-(f_{n}-g_{n})(y)|\leq L\left\|x-y\right\|_{2}.$
If $(f_{n}-g_{n})_{n}$ tends to 0 pointwise over a dense subset of $\Theta$ then, for any compact set $K$ in $\Theta$, $(f_{n}-g_{n})_{n}$ converges uniformly over $K$ to 0.
Proof of Lemma C.7.
For all integer $n$ we let $h_{n}=f_{n}-g_{n}$ and we denote by $\Theta_{0}$ a dense subset in $\Theta$ onto which $(h_{n})_{n}$ converges pointwise to 0. Let $K$ be a compact subset in $\Theta$. Let $L>0$ and $n_{1}$ be such that for all $n\geq n_{1}$ and all $x,y\in K$,
$|h_{n}(x)-h_{n}(y)|\leq L\left\|x-y\right\|_{2}.$ (C.5)
Let $\varepsilon>0$. Denote by $\Theta_{1}$ an $(\varepsilon/(2L))$-net for $K$ in $K\cap\Theta_{0}$ with respect to the $\ell_{2}^{d}$-norm. Since $K$ is compact we can choose $\Theta_{1}$ to be finite. Since $\Theta_{1}$ is finite and $(h_{n}(\theta))_{n}$ tends to 0 for all $\theta\in\Theta_{1}$, there exists $n_{2}\geq n_{1}$ such that for all $\theta\in\Theta_{1}$, $|h_{n}(\theta)|\leq\varepsilon/2$. Let $\theta\in K$. Let $\theta_{1}\in\Theta_{1}$ be such that $\left\|\theta-\theta_{1}\right\|_{2}\leq\varepsilon/2$. Then, for all $n\geq n_{2}$, we have
$|h_{n}(\theta)|\leq|h_{n}(\theta)-h_{n}(\theta_{1})|+|h_{n}(\theta_{1})|\leq L\left\|\theta-\theta_{1}\right\|_{2}+\frac{\varepsilon}{2}\leq\varepsilon.$
This concludes the proof since the latter holds for all $\theta\in K$. ∎
The next result is the diagonalization method used to prove Theorem II.1 in Andersen1982 that we reproduce here in our setup for the sake of completeness.
Lemma C.8.
(Theorem II.1 from Andersen1982)Let $\Theta$ be a non empty set. Let $(\hat{Q}_{n})_{n}$ be a sequence of random functions and $(Q_{n})_{n}$ be a sequence of (deterministic) functions, all defined on $\Theta$. We assume that for all $\theta\in\Theta$, $\hat{Q}_{n}(\theta)-Q_{n}(\theta)\overset{p}{\to}0$ as $n\to+\infty$. Then, for any countable subset $\Theta^{\prime}$ in $\Theta$ there exists an increasing sequence $(\psi_{n})_{n}$ of integers such that for all $\theta\in\Theta^{\prime}$,
$\hat{Q}_{\psi_{n}}(\theta)-Q_{\psi_{n}}(\theta)\overset{a.s.}{\to}0.$
Proof of Lemma C.8: Denote for all $n$, $\hat{H}_{n}=\hat{Q}_{n}-Q_{n}$. Denote by $(x_{j})_{j\in{\mathbb{N}}}$ the sequence of all elements in $\Theta^{\prime}$. Since $(\hat{H}_{n}(x_{1}))_{n}$ tends in probability to 0, we can extract a sub-sequence $(\phi_{n,1})_{n}$ along which the converge is almost sure. Next, since $(\hat{H}_{\phi_{n,1}}(x_{2}))_{n}$ tends in probability to 0, it is possible to further construct a sub-sequence $(\phi_{n,2})_{n}$ along which the converge is almost sure. We repeat the argument and get that for all integers $k$, it is possible to construct a sequence $(\phi_{n,k})_{n}$ such that for all $i\in[k]$, $(\hat{H}_{\phi_{n,k}}(x_{i}))_{n}$ tends almost surely to 0. Once, this construction is done, we move to the diagonalization argument used in the proof of Theorem II.1 from Andersen1982: we construct a new sequence of integers, denoted by $(\psi_{n})_{n}$, by setting $\psi_{1}$ to be the first element in $(\phi_{n,1})_{n}$, then $\psi_{2}$ to be the second one in $(\phi_{n,2})_{n}$, etc.. Along this new sequence $(\psi_{n})_{n}$ we have for all $i\in{\mathbb{N}}$, $(\hat{H}_{\psi_{n}}(x_{i}))_{n}$ converges almost surely to 0. $\square$
C.4 Some properties of the loss functions, its differential and Hessian
We first start with a result that we use to prove the consistency result from Theorem 1.
Lemma C.9.
Let $\beta\in\mathbb{R}^{p}$. There exists a constant $C_{1}$ such that for all $\xi,\xi^{\prime}\in\Xi$, we have
${\mathbb{E}}_{\beta_{\star}}\left[(\ell_{\xi}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi}(Y|X,\beta))(\ell_{\xi^{\prime}}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi^{\prime}}(Y|X,\beta))|X\right]\leq C_{1}\mathbb{P}_{\beta_{\star}}\left[\xi\mbox{ and }\xi^{\prime}\mbox{ are both active}|X\right].$
One can choose $C_{1}$ to be
$C_{1}=c_{D}(L_{N}\left\|\beta\right\|_{2}+1)^{2}\left((\bar{\lambda}+1)^{4}+\frac{\exp(-2^{D-1}e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{2^{D-1}}}\right).$
where $c_{D}$ is a constant depending only on $D$, $L_{N}=\max_{\xi}\left\|\tilde{X}_{\xi}\right\|_{2}$ and any $\bar{\lambda}$ and $\underline{\lambda}$ such that $\bar{\lambda}\geq\max\left(\lambda_{\bf i}(\beta,\theta):{\bf i}\in{\cal I}\right)$ and $\underline{\lambda}\leq\min\left(\lambda_{\bf i}(\beta,\theta):{\bf i}\in{\cal I}\right)$.
Proof.
Since $\ell_{\xi}(Y|X,\beta)\geq 0$ - as the negative log of a probability - the correlation between $\ell_{\xi}$ and $\ell_{\xi^{\prime}}$ is less than the expectation of their product:
${\mathbb{E}}_{\beta_{\star}}\left[(\ell_{\xi}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi}(Y|X,\beta))(\ell_{\xi^{\prime}}(Y|X,\beta)-{\mathbb{E}}_{\beta_{\star}}\ell_{\xi^{\prime}}(Y|X,\beta))|X\right]\leq{\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi^{\prime}}(Y|X,\beta)|X\right].$ (C.6)
Then, we observe that for all $y\in{\mathbb{N}}^{\cal I}$, when $\xi$ is not active with respect to some $y$ then $\mathcal{O}_{\xi}(y)$ contains only one element which is $y$ and so $\mathbb{P}_{\beta_{\star}}[Y=y|X,Y\in\mathcal{O}_{\xi}(y)]=1$, hence $\ell_{\xi}(y|X,\beta)=0$. As a consequence, for all $\xi$ and $\xi^{\prime}$, we have
${\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi^{\prime}}(Y|X,\beta)|X\right]={\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi^{\prime}}(Y|X,\beta)I\left(\xi\mbox{ and }\xi^{\prime}\mbox{ are both active}\right)|X\right].$
However, the loss functions $\ell_{\xi}$ are not almost surely bounded, we therefore need to deal carefuly with this issue. We recall that for every polyads $\xi=\begin{pmatrix}j_{1}&j_{2}&\cdots&j_{D}\\ j_{1}^{\prime}&j_{2}^{\prime}&\cdots&j_{D}^{\prime}\\ \end{pmatrix}$ we denote ${\cal I}_{\xi}=\{j_{1},j_{1}^{\prime}\}\times\cdots\{j_{D},j_{D}^{\prime}\}$. We have
${\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi}(Y|X,\beta)|X\right]=\sum_{y=(y_{\bf i})_{\bf i}}\mathbb{P}_{\beta_{\star}}[Y=y|X]\left(\ln\mathbb{P}_{\beta}\left(Y=y|X,Y\in\mathcal{O}_{\xi}(y)\right)\right)\left(\ln\mathbb{P}_{\beta}\left(Y=y|X,Y\in\mathcal{O}_{\xi^{\prime}}(y)\right)\right)$
$=\sum_{y=(y_{\bf i})_{\bf i}}\mathbb{P}_{\beta_{\star}}[Y=y|X]\left(\ln\frac{\mathbb{P}_{\beta}[Y=y|X]}{\sum_{y^{\prime}\in\mathcal{O}_{\xi}(y)}\mathbb{P}_{\beta}[Y=y^{\prime}|X]}\right)\left(\ln\frac{\mathbb{P}_{\beta}[Y=y|X]}{\sum_{y^{\prime}\in\mathcal{O}_{\xi^{\prime}}(y)}\mathbb{P}_{\beta}[Y=y^{\prime}|X]}\right)$
$=\sum_{y=(y_{\bf i})_{\bf i}}\mathbb{P}_{\beta_{\star}}[Y=y|X]\left(\ln\sum_{y^{\prime}\in\mathcal{O}_{\xi}(y)}\frac{\mathbb{P}_{\beta}[Y=y^{\prime}|X]}{\mathbb{P}_{\beta}[Y=y|X]}\right)\left(\ln\sum_{y^{\prime}\in\mathcal{O}_{\xi^{\prime}}(y)}\frac{\mathbb{P}_{\beta}[Y=y^{\prime}|X]}{\mathbb{P}_{\beta}[Y=y|X]}\right)$
$=\sum_{(y_{\bf i})_{{\bf i}\in{\cal I}_{\xi}\cup{\cal I}_{\xi^{\prime}}}}\mathbb{P}_{\beta_{\star}}[Y_{{\bf i}}=y_{\bf i},\forall{\bf i}\in{\cal I}_{\xi}\cup{\cal I}_{\xi^{\prime}}|X]\left(\ln\left(\sum_{r=-m_{\xi}(y)}^{M_{\xi}(y)}\frac{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i}^{r},\forall{\bf i}\in{\cal I}_{\xi}|X]}{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i},\forall{\bf i}\in{\cal I}_{\xi}|X]}\right)\right)\left(\ln\left(\sum_{r=-m_{\xi^{\prime}}(y)}^{M_{\xi^{\prime}}(y)}\frac{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i}^{r},\forall{\bf i}\in{\cal I}_{\xi^{\prime}}|X]}{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i},\forall{\bf i}\in{\cal I}_{\xi^{\prime}}|X]}\right)\right)$
where we recall that for all polyads $\xi$, all $\bf i$$\in{\cal I}$ and all $r\in\{-m_{\xi}(y),\ldots,M_{\xi}(y)\},y_{\bf i}^{r}=y_{\bf i}+rs_{\xi}({\bf i})$. Next, we use the independence of the $Y_{\bf i}$’s conditionally on $X$ to get that for all $r\in\{-m_{\xi}(y),\ldots,M_{\xi}(y)\}$,
$\frac{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i}^{r},\forall{\bf i}\in{\cal I}_{\xi}|X]}{\mathbb{P}_{\beta}[Y_{\bf i}=y_{\bf i},\forall{\bf i}\in{\cal I}_{\xi}|X]}=\prod_{{\bf i}\in{\cal I}_{\xi}}\frac{\lambda_{\bf i}^{y_{\bf i}^{r}}y_{\bf i}!}{\lambda_{\bf i}^{y_{\bf i}}y_{\bf i}^{r}!}=\prod_{{\bf i}\in{\cal I}_{\xi}}\lambda_{\bf i}^{rs_{\xi}({\bf i})}\frac{y_{\bf i}!}{y_{\bf i}^{r}!}$
where $\lambda_{\bf i}=\exp\left(\bigl<X_{\bf i},\beta\bigr>+\sum_{g\in{\cal G}}\theta_{g({\bf i})}^{g})\right)$. By the diff-in-diff property of the sign function $s_{\xi}$ from Proposition 4, we have
$\prod_{{\bf i}\in{\cal I}_{\xi}}\lambda_{\bf i}^{rs_{\xi}({\bf i})}=\exp(r\bigl<\tilde{X}_{\xi},\beta\bigr>):=w_{\xi,r}$
where $\tilde{X}_{\xi}$ is the polyads feature defined in (12). Furthermore, we define for all $r\in\{-m_{\xi}(y),\ldots,M_{\xi}(y)\}$,
$v_{\xi,r}(y):=\prod_{{\bf i}\in{\cal I}_{\xi}}\frac{y_{\bf i}!}{y_{\bf i}^{r}!}=\left\{\begin{array}[]{cc}\frac{\prod_{{\bf i}:s_{\xi}({\bf i})=-1}y_{\bf i}(y_{\bf i}-1)\cdots(y_{\bf i}-r+1)}{\prod_{{\bf i}:s_{\xi}({\bf i})=1}(y_{\bf i}+r)\cdots(y_{\bf i}+1)}&\mbox{ if }r>0\\ \frac{\prod_{{\bf i}:s_{\xi}({\bf i})=1}y_{\bf i}(y_{\bf i}-1)\cdots(y_{\bf i}+r+1)}{\prod_{{\bf i}:s_{\xi}({\bf i})=-1}(y_{\bf i}-r)\cdots(y_{\bf i}+1)}&\mbox{ if }r<0\\ 1&\mbox{ if }r=0.\end{array}\right$.
We have, for $M_{\xi}=M_{\xi}(Y)$, $m_{\xi}=m_{\xi}(Y)$ and $V_{\xi,r}:=v_{\xi,r}(Y)$ for all $r\in\{-m_{\xi},\cdots,M_{\xi}\}$,
${\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi}(Y|X,\beta)|X\right]={\mathbb{E}}_{\beta_{\star}}\left[\ln\left(\sum_{r=-m_{\xi}}^{M_{\xi}}w_{\xi,r}V_{\xi,r}\right)\ln\left(\sum_{r=-m_{\xi^{\prime}}}^{M_{\xi^{\prime}}}w_{\xi^{\prime},r}V_{\xi^{\prime},r}\right)|X\right].$
[…]
$w_{\zeta,r},w_{\zeta,r}^{-1}\leq\exp\left[(M_{\zeta}+m_{\zeta})|\bigl<\tilde{X}_{\zeta},\beta\bigr>|\right]\mbox{ and }V_{\zeta,r},V_{\zeta,r}^{-1}\leq\prod_{{\bf i}\in{\cal I}_{\zeta}}(2Y_{\bf i}+1)^{M_{\zeta}+m_{\zeta}}$
to get the following bound:
${\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi^{\prime}}(Y|X,\beta)|X\right]$
$\leq{\mathbb{E}}_{\beta_{\star}}\left[\prod_{\zeta\in\{\xi,\xi^{\prime}\}}\left((M_{\zeta}+m_{\zeta})|\bigl<\tilde{X}_{\zeta},\beta\bigr>|+\ln\left((M_{\zeta}+m_{\zeta})\prod_{{\bf i}\in{\cal I}_{\xi}}(2Y_{\bf i}+1)^{M_{\zeta}+m_{\zeta}}\right)I(M_{\zeta}+m_{\zeta}\geq 1)\right)|X\right]$
$\leq{\mathbb{E}}_{\beta_{\star}}\left[\prod_{\zeta\in\{\xi,\xi^{\prime}\}}\left((M_{\zeta}+m_{\zeta})|\left(\bigl<\tilde{X}_{\zeta},\beta\bigr>|+1\right)+(M_{\zeta}+m_{\zeta})\ln\left(\prod_{{\bf i}\in{\cal I}_{\xi}}(2Y_{\bf i}+1)\right)\right)|X\right]$
$\leq{\mathbb{E}}_{\beta_{\star}}\left[\prod_{\zeta\in\{\xi,\xi^{\prime}\}}\left((M_{\zeta}+m_{\zeta})|\left(\bigl<\tilde{X}_{\zeta},\beta\bigr>|+1\right)+2^{D}(M_{\zeta}+m_{\zeta})\left(\sum_{{\bf i}\in{\cal I}_{\xi}}\ln(Y_{\bf i}+1)\right)\right)|X\right]$
$\leq{\mathbb{E}}_{\beta_{\star}}\left[(M_{\xi}+m_{\xi})(M_{\xi^{\prime}}+m_{\xi^{\prime}})\prod_{\zeta\in\{\xi,\xi^{\prime}\}}\left(L_{N}\left\|\beta\right\|_{2}+1+2^{D}\sum_{{\bf i}\in{\cal I}_{\zeta}}\ln\left(Y_{\bf i}+1\right)\right)|X\right]$ (C.7)
where $L_{N}=\max_{\xi}\left\|\tilde{X}_{\xi}\right\|_{2}$. Next, we remark that all of the four terms in the product $(M_{\xi}+m_{\xi})(M_{\xi^{\prime}}+m_{\xi^{\prime}})$ are product of two minimum of Poisson variables. To handle (C.4), we rely on the next lemma.
Lemma C.10.
Let $(U_{i})_{i\in{\mathbb{N}}}$ be a sequence of independent Poisson variables with intensities $(\lambda_{i})_{i\in{\mathbb{N}}}$. Let $I_{1},I_{2},I_{3},I_{4}\subset{\mathbb{N}}$ be such that $I_{1}\cap(I_{2}\cup I_{4})=\emptyset$ and $I_{3}\cap(I_{2}\cup I_{4})=\emptyset$ (but possibly $I_{1}\cap I_{3}\neq\emptyset$ and $I_{2}\cap I_{4}\neq\emptyset$). Let $\underline{\lambda}$ and $\bar{\lambda}$ be such that $\underline{\lambda}\leq\lambda_{i}\leq\bar{\lambda}$ for all $i\in\cup_{k=1,2,3,4}I_{k}$. We have for all $i,j\in\cup_{k=1,2,3,4}I_{k}$
${\mathbb{E}}\left[(m_{1}+m_{2})(m_{3}+m_{4})\ln(U_{i}+1)\ln(U_{j}+1)\right]\leq C_{0}\mathbb{P}\left[m_{1}+m_{2}\geq 1\mbox{ and }m_{3}+m_{4}\geq 1\right]$
where $m_{k}=\wedge_{i\in I_{k}}U_{i}$ and $C_{0}$ is given by
$C_{0}=4\max_{\begin{subarray}{c}r=1,2\\ s=3,4\end{subarray}}\left(2e^{8}(\bar{\lambda}+1)^{4}+\frac{e\exp(-|I_{r}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{r}|}}+\frac{e\exp(-|I_{s}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{s}|}}\right).$
For all interger $r\geq 1$, we also have
${\mathbb{E}}[m_{1}^{r}]\leq C_{|I_{1}|,r}\mathbb{P}[m_{1}\geq 1]\mbox{ where for all }n,C_{n,r}=(e^{2}\bar{\lambda})^{r}+\frac{e\exp(-ne^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{n}}.$
Proof.
Let $I\subset\cup_{k=1,2,3,4}I_{k}$ and $i\in\cup_{k=1,2,3,4}I_{k}$. Let $M=\wedge_{i\in I}U_{i}$. We first establish some preliminary results on $M$. It follows from a Chernoff bound that for all $i$ and all $k\geq e^{2}\lambda_{i}$, $\mathbb{P}[U_{i}\geq k]\leq\exp(-k)$. As a consequence, we have for all $r\in{\mathbb{N}}^{*}$,
${\mathbb{E}}[M^{r}]$ $=\sum_{k=1}^{\infty}\mathbb{P}[M\geq k^{1/r}]=\sum_{k=1}^{(e^{2}\bar{\lambda})^{r}-1}\mathbb{P}[M\geq k^{1/r}]+\sum_{k=(e^{2}\bar{\lambda})^{r}}^{\infty}\mathbb{P}[M\geq k^{1/r}]$
$\leq(e^{2}\bar{\lambda})^{r}\mathbb{P}[M\geq 1]+\sum_{k=(e^{2}\bar{\lambda})^{r}}^{\infty}\exp(-k)\leq(e^{2}\bar{\lambda})^{r}\mathbb{P}[M\geq 1]+e\exp(-ne^{2}\bar{\lambda})$
where $n=|I|$. We also have $\mathbb{P}[M\geq 1]=\prod_{i}(1-\mathbb{P}[U_{i}=0])\geq(1-\exp(-\underline{\lambda}))^{n}$. As a consequence,
${\mathbb{E}}[M^{r}]\leq C_{n,r}(\underline{\lambda},\bar{\lambda})\mathbb{P}[M\geq 1]\mbox{ where }C_{n,r}=C_{n,r}(\underline{\lambda},\bar{\lambda})=(e^{2}\bar{\lambda})^{r}+\frac{e\exp(-ne^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{n}}.$ (C.8)
If $i\notin I$ then by independence,
${\mathbb{E}}[M\ln(U_{i}+1)]={\mathbb{E}}[M]{\mathbb{E}}[\ln(U_{i}+1)]\leq\bar{\lambda}C_{n,1}(\underline{\lambda},\bar{\lambda})\mathbb{P}[M\geq 1]$
If $i\in I$ then we use that
$M\ln(U_{i}+1)\leq M^{2}+\ln^{2}(U_{i}+1)I(M\geq 1)=M^{2}+\ln^{2}(U_{i}+1)I(U_{i}\geq 1)I(U_{j}\geq 1,\forall j\neq i)\leq M^{2}+U_{i}I(U_{j}\geq 1,\forall j\neq i),$
that ${\mathbb{E}}[U_{i}]=C(\lambda_{i})\mathbb{P}[U_{i}\geq 1]$ where $C(\lambda_{i})=\lambda_{i}(1-\exp(-\lambda_{i}))^{-1}$ and $\ln^{2}(x+1)\leq x,\forall x\geq 1$ to get
${\mathbb{E}}[M\ln(U_{i}+1)]$ $\leq{\mathbb{E}}[M^{2}]+{\mathbb{E}}[U_{i}]\mathbb{P}[U_{j}\geq 1,\forall j\neq i]\leq C_{n,2}(\underline{\lambda},\bar{\lambda})\mathbb{P}[M\geq 1]+C(\lambda_{i})\prod_{i}\mathbb{P}[U_{i}\geq 1]$
$\leq\left(C_{n,2}(\underline{\lambda},\bar{\lambda})+C(\lambda_{i})\right)\mathbb{P}[M\geq 1].$
We have
${\mathbb{E}}\left[(m_{1}+m_{2})(m_{3}+m_{4})\ln(U_{i}+e)\ln(U_{j}+e)\right]=\sum_{r=1,2;s=3,4}{\mathbb{E}}\left[m_{r}m_{s}\ln(U_{i}+1)\ln(U_{j}+1)\right]$
The four terms in the right-hand side of the equality above are of the same ’type’, the only difference we need to take care of is the relation between $I_{s}$, $I_{r}$, $i$ and $j$. Let us consider the hardest situation when $I_{s}$ and $I_{r}$ intersect and $i$ and $j$ are different and both in this intersection. The other situations are easier using independence and can be handled using similar technics as below. We are therefore considering now the situation where $i\neq j\in I:=I_{r}\cap I_{s}$. We use the following decomposition:
$m_{r}m_{s}\ln(U_{i}+1)\ln(U_{j}+1)$ $\leq m_{r}^{2}m_{s}^{2}+\ln^{2}(U_{i}+1)\ln^{2}(U_{j}+1)I(m_{r}m_{s}\geq 1)$
$\leq m_{r}^{2}m_{s}^{2}+U_{i}U_{j}I(\forall k\in I_{r}\cup I_{s}\backslash\{i,j\},U_{k}\geq 1).$
Let us handle the two terms in the right-hand side separately. We have for
${\mathbb{E}}[m_{r}^{2}m_{s}^{2}]=\sum_{k=1}^{\infty}\mathbb{P}[m_{r}^{2}m_{s}^{2}\geq k]\leq\sum_{k=1}^{\infty}\mathbb{P}[m_{r}^{2}\geq\sqrt{k}\mbox{ and }m_{s}\geq 1]+\mathbb{P}[m_{s}^{2}\geq\sqrt{k}\mbox{ and }m_{r}\geq 1]$
$\leq 2(e^{2}\bar{\lambda})^{4}\mathbb{P}[m_{s}\geq 1\mbox{ and }m_{r}\geq 1]+\sum_{k\geq(e^{2}\bar{\lambda})^{4}}\mathbb{P}[m_{s}\geq k^{1/4}\mbox{ and }m_{r}\geq 1]+\mathbb{P}[m_{r}\geq k^{1/4}\mbox{ and }m_{s}\geq 1]$
and using independence and the Chernoff bound above we have for all $k\geq(e^{2}\bar{\lambda})^{4}$,
$\mathbb{P}[m_{r}\geq k^{1/4}\mbox{ and }m_{s}\geq 1]$ $=\prod_{i\in I_{r}}\mathbb{P}[U_{i}\geq k^{1/4}]\prod_{i\in I_{s}\backslash I}\mathbb{P}[U_{i}\geq 1]\leq\exp(-|I_{r}|k^{1/4})\prod_{i\in I_{s}\backslash I}\mathbb{P}[U_{i}\geq 1]$
and so
$\sum_{k\geq(e^{2}\bar{\lambda})^{4}}\mathbb{P}[m_{r}\geq k^{1/4}\mbox{ and }m_{r}\geq 1]\leq\sum_{k\geq(e^{2}\bar{\lambda})^{4}}\exp(-|I_{r}|k^{1/4})\prod_{i\in I_{s}\backslash I}\mathbb{P}[U_{i}\geq 1]\leq e\exp(-|I_{r}|e^{2}\bar{\lambda})\prod_{i\in I_{s}\backslash I}\mathbb{P}[U_{i}\geq 1]$
$\leq\frac{e\exp(-|I_{r}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{r}|}}\mathbb{P}[\forall i\in I_{r},U_{i}\geq 1]\prod_{i\in I_{s}\backslash I}\mathbb{P}[U_{i}\geq 1]=\frac{e\exp(-|I_{r}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{r}|}}\mathbb{P}[m_{r}\geq 1\mbox{ and }m_{s}\geq 1].$
Gathering the results above, we obtain
${\mathbb{E}}[m_{r}^{2}m_{s}^{2}]\leq\left(2(e^{2}\bar{\lambda})^{4}+\frac{e\exp(-|I_{r}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{r}|}}+\frac{e\exp(-|I_{s}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{s}|}}\right)\mathbb{P}[m_{r}\geq 1\mbox{ and }m_{s}\geq 1].$
Next, using that ${\mathbb{E}}[U_{i}^{2}]=2\lambda_{i}(1-\exp(-\lambda_{i}))^{-1}\mathbb{P}[U_{i}\geq 1]\leq 4(\bar{\lambda}+1)\mathbb{P}[U_{i}\geq 1]$, we have
[…]
$\leq 8(\bar{\lambda}+1)\mathbb{P}[m_{r}\geq 1\mbox{ and }m_{s}\geq 1].$
Collecting, the results above, we conclude that
${\mathbb{E}}[m_{r}m_{s}\ln(U_{i}+1)\ln(U_{j}+1)]\leq\left(2(e^{2}\bar{\lambda})^{4}+8(\bar{\lambda}+1)+\frac{e\exp(-|I_{r}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{r}|}}+\frac{e\exp(-|I_{s}|e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{|I_{s}|}}\right)\mathbb{P}[m_{r}\geq 1\mbox{ and }m_{s}\geq 1].$
Finally, the result follows since we have for all $r\in\{1,2\}$ and $s\in\{3,4\}$ that
$\mathbb{P}[m_{r}\geq 1\mbox{ and }m_{s}\geq 1]\leq\mathbb{P}\left[m_{1}+m_{2}\geq 1\mbox{ and }m_{3}+m_{4}\geq 1\right]$
Using Lemma C.10 in (C.4), we get
${\mathbb{E}}_{\beta_{\star}}\left[\ell_{\xi}(Y|X,\beta)\ell_{\xi^{\prime}}(Y|X,\beta)|X\right]\leq C_{D}(L_{N}\left\|\beta\right\|_{2}+1)^{2}C_{1}\mathbb{P}[M_{\xi}+m_{\xi}\geq 1\mbox{ and }M_{\xi^{\prime}}+m_{\xi^{\prime}}\geq 1].$
where $C_{D}$ is a constant depending only on $D$ and $C_{1}$ is the constant coming from Lemma C.10 adapted to our setup:
$C_{1}=4\left(2e^{8}(\bar{\lambda}+1)^{4}+\frac{2e\exp(-2^{D-1}e^{2}\bar{\lambda})}{(1-\exp(-\underline{\lambda}))^{2^{D-1}}}\right).$
Next we move to results that will be useful for both consistency and a.n. regarding quantities involving random polyads and/or random edges. We recall that $n=|{\cal I}|$ is the total number of edges and $N=|\Xi|$ is the total number of polyads. We also recall that two polyads can either share no edge in common and have a total number of edges in common in $\{2^{r}:r\in\{0,1,\cdots,D\}\}$. That is why in the lemmas below we only consider the cases where $q\in\{2^{r}:r\in\{0,1,\cdots,D\}\}$.
Lemma C.11.
There exists an absolute $c_{0}$ such that the following holds. Let $\xi$ and $\xi^{\prime}$ be two polyads. Let $c\in\mathbb{R}^{d}$. We have
[…]
and
[…]
Proof.
It follows from Lemma 1 that
$\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>=\left({\mathbb{E}}_{\beta_{\star}}\left[m_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]-m_{\xi}(Y)\right)\bigl<\widetilde{X}_{\xi},c\bigr$>
where $Y,Y^{\prime}$ are iid distributed according to $\mathbb{P}_{\beta_{\star},\theta}$. It follows from Lemma C.3 that, given $Y^{\prime}\in\mathcal{O}_{\xi}(Y)$, we have $m_{\xi}(Y^{\prime})\leq m_{\xi}(Y)+M_{\xi}(Y)$. As a consequence,
[…]
and so
${\mathbb{E}}_{\beta_{\star}}\left[|\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>\bigl<\nabla\ell_{\xi^{\prime}}(\beta_{\star}),c\bigr>||X\right]\leq{\mathbb{E}}_{\beta_{\star}}\left[M_{\xi}(Y)M_{\xi^{\prime}}(Y)|X\right]|\bigl<\widetilde{X}_{\xi},c\bigr>||\bigl<\widetilde{X}_{\xi^{\prime}},c\bigr$>
and the result follows from Lemma C.10 since
${\mathbb{E}}_{\beta_{\star}}\left[M_{\xi}(Y)M_{\xi^{\prime}}(Y)|X\right]\leq C_{0}\mathbb{P}_{\beta_{\star}}[M_{\xi}(Y)\geq 1\mbox{ and }M_{\xi^{\prime}}(Y)\geq 1|X]\leq C_{0}\mathbb{P}_{\beta_{\star}}[\xi\mbox{ and }\xi^{\prime}\mbox{ are active}|X].$
Note that it follows from Lemma 1 that we also have
$\bigl<\nabla\ell_{\xi}(\beta_{\star}),c\bigr>=\left(M_{\xi}(Y)-{\mathbb{E}}_{\beta_{\star}}\left[M_{\xi}(Y^{\prime})|X,Y^{\prime}\in\mathcal{O}_{\xi}(Y)\right]\right)\bigl<\widetilde{X}_{\xi},c\bigr$>
and from Lemma C.3 that, given $Y^{\prime}\in\mathcal{O}_{\xi}(Y)$, we have $M_{\xi}(Y^{\prime})\leq m_{\xi}(Y)+M_{\xi}(Y)$. As a consequence,
[…]
Let ${\bf i}\in{\cal E}(\xi)\cap{\cal E}(\xi^{\prime})$. We have $s_{\xi}({\bf i}),s_{\xi^{\prime}}({\bf i})\in\{-1,1\}$ and so $Y_{\bf i}$ is independent of $m_{\xi}(Y)$ or $M_{\xi}(Y)$ and of $m_{\xi}(Y^{\prime})$ or $M_{\xi}(Y^{\prime})$. For all cases, one can choose one of the upper bounds in (C.9) or (C.10) to get an upper bound independent of $Y_{\bf i}$ for both $\xi$ and $\xi^{\prime}$. For instance, if $Y_{\bf i}$ is independent of $m_{\xi}(Y)$ and of $M_{\xi^{\prime}}(Y)$ (the three other cases can be handled using similar arguments) then we get
[…]
where the last but one inequality follows from Lemma C.10. ∎
C.5 A result on the expectation of the empirical $L_{3}/L_{2}$ ratio and other probability results
In this section, we provide the tools we used in (B.14) to show that as $n\to\infty$
$\frac{1}{n}\sum_{{\bf i}}{\mathbb{E}}\left[|R_{\bf i}|^{3}|X\right]=o(\sqrt{n}).$ (C.11)
Since we can write the mean above as the expectation of the ratio to the power three of the third empirical moment over the second one of independent variables, we provide upper bound for such a quantity in this section.
Proposition 5.
Let $((X_{i,n})_{i\in[n]})_{n}$ be a triangular array of real-valued random variables. We assume that for all $n\in{\mathbb{N}}^{*}$, $(X_{i,n})_{i\in[n]}$ are independent. For all $n\in{\mathbb{N}}^{*}$, we consider for $k=2,3$,
$\hat{S}_{k,n}=\frac{1}{n}\sum_{i=1}^{n}|X_{i,n}|^{k},S_{k,n}={\mathbb{E}}\hat{S}_{k,n}\mbox{ and }V_{k,n}:=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{V}}\left(|X_{i,n}|^{k}\right).$
We assume that
$\frac{S_{3,n}+u_{n}\sqrt{V_{3,n}}}{\left(S_{2,n}-u_{n}\sqrt{V_{2,n}}\right)^{3/2}}=o\left(\sqrt{n}\right)$ (C.12)
for some $u_{n}\sim a_{n}/\sqrt{n}$ and $a_{n}\to+\infty$ when $n\to\infty$. Under this assumption, as $n$ tends to $\infty$, we have
${\mathbb{E}}\left[\frac{\hat{S}_{3,n}}{\left(\hat{S}_{2,n}\right)^{3/2}}\right]=o\left(\sqrt{n}\right).$
Proof.
Let $n\in{\mathbb{N}}^{*}$ and denot $\hat{S}_{k}=\hat{S}_{k,n}$, $S_{k}=S_{k,n}$ and $V_{k}=V_{k,n}$. We consider the following decomposition:
${\mathbb{E}}\left[\frac{\hat{S}_{3}}{\left(\hat{S}_{2}\right)^{3/2}}\right]={\mathbb{E}}\left[\frac{\hat{S}_{3}}{\left(\hat{S}_{2}\right)^{3/2}}\left(I_{\Omega_{0}}+I_{\Omega_{0}^{c}}\right)\right]$
according to the event
$\Omega_{0}:=\left\{|\hat{S}_{2}-S_{2}|\leq u_{n}\sqrt{V_{2}}\mbox{ and }|\hat{S}_{3}-S_{3}|\leq u_{n}\sqrt{V_{3}}\right\}.$
It follows from Markov’s inequality that $\mathbb{P}[\Omega_{0}^{c}]=o(1)$ and since we have $\hat{S}_{3}\leq\sqrt{n}\left(\hat{S}_{2}\right)^{3/2}$ a.s., we get
${\mathbb{E}}\left[\frac{\hat{S}_{3}}{\left(\hat{S}_{2}\right)^{3/2}}I_{\Omega_{0}^{c}}\right]=o(\sqrt{n}).$
On the complementary event, we use Assumption C.12 to get
${\mathbb{E}}\left[\frac{\hat{S}_{3}}{\left(\hat{S}_{2}\right)^{3/2}}I_{\Omega_{0}}\right]\leq{\mathbb{E}}\left[\frac{S_{3}+|S_{3}-\hat{S}_{3}|}{\left(S_{2}-|\hat{S}_{2}-S_{2}|\right)^{3/2}}I_{\Omega_{0}}\right]\leq\frac{S_{3}+u_{n}\sqrt{V_{3}}}{\left(S_{2}-u_{n}\sqrt{V_{2}}\right)^{3/2}}=o\left(\sqrt{n}\right).$
Assumption C.12 is a pretty weak assumption. For instance, when the $X_{i,n}$’s have all the same order $3$ and $2$ moments $\mu_{3}$ and $\mu_{2}$ as well as the same variance terms $v_{3}$ and $v_{2}$ then Assumption C.12 is equivalent to
$\frac{\mu_{3}+a_{n}\sqrt{v_{3}/n}}{\left(\mu_{2}-a_{n}\sqrt{v_{2}/n}\right)^{3/2}}=o\left(\sqrt{n}\right)$ (C.13)
where $(a_{n})_{n}$ is any sequence such that $a_{n}\to\infty$, for instance $a_{n}=\log n$. In that case, (C.13) is trivially satisfied.
Lemma C.12 (Example 11.9 from MR1652247).
Let $X,Y,Z$ be random variables. We have ${\mathbb{E}}[{\mathbb{E}}[X|Y,Z]|Y]={\mathbb{E}}[X|Y]$.
Lemma C.13.
Let $X,Y,Z$ be random variables such that $Y$ and $Z$ are independent conditionally on $X$. Then for all measurable functions $g$ and $f$, we have
  • ${\mathbb{E}}[f(X,Y)g(X,Z)|X]={\mathbb{E}}[f(X,Y)|X]{\mathbb{E}}[g(X,Z)|X]$, in other words, conditionally on $X$, $(X,Y)$ and $(X,Z)$ are independent,
  • ${\mathbb{E}}[f(X,Y)|X,Z]={\mathbb{E}}[f(X,Y)|X]$.
Proof.
For the first item, we check the result for all tensor functions $g=g_{1}\otimes g_{2}$ and $f=f_{1}\otimes f_{2}$ for which the result is trivial. For the second item, it is enough to show that $\mathbb{P}^{(X,Y)|(X,Z)}=\mathbb{P}^{(X,Y)|X}$. we have $\mathbb{P}^{(X,Y)|(X,Z)}=\mathbb{P}^{X|(X,Z)}\otimes\mathbb{P}^{Y|(X,Z)}$, $\mathbb{P}^{X|(X,Z)}=\delta_{X}$ and for all measurable sets $A,B,C$, by conditional independence,
$\mathbb{P}[Y\in A|X\in B,Z\in C]=\frac{\mathbb{P}[Y\in A,Z\in C|X\in B]}{\mathbb{P}[X\in B,Z\in C]}=\mathbb{P}[Y\in A|X\in B]$
so that $\mathbb{P}^{Y|(X,Z)}=\mathbb{P}^{Y|X}$. Hence, $\mathbb{P}^{(X,Y)|(X,Z)}=\delta_{X}\otimes\mathbb{P}^{Y|X}=\mathbb{P}^{(X,Y)|X}$. ∎
Appendix D Health data application
Based on claims data issued from the French comprehensive Système National des Données de Santé (SNDS) database that contain information on each encounter between patients and physicians, we aggregate visits at the doctor level for each month between January 2016 and October 2018.121212Our observation period runs from 2016 to 2018, but we exclude November and December 2018 because of right-censoring. Indeed, the processing time for claim files may last up to 40 days, and the date of record may well fall after the end of 2018. April and May 2017 are also removed from the sample to neutralize anticipation effects. We restrict our attention to doctors aged 30 to 79 practicing in mainland France (at the exclusion of overseas and Corsica).
Table D.1: Effect of the fee increase on physician activity
Control group Direct access specialists GPs in unregulated sector 2
Specification Linear PPML Linear PPML
Post $\times$ Treatment 0.0964∗∗∗ 0.104∗∗∗ 0.0878∗∗∗ 0.131∗∗∗
(0.00572) (0.00654) (0.00488) (0.00671)
Physician FE Yes Yes Yes Yes
Month-year FE Yes Yes Yes Yes
Observations 1806756 2159008 1742229 2070592
(Pseudo) R2 0.884 0.754 0.873 0.740
Source: French SNDS claims data aggregated at the doctor-month level.
Sample: GPs and direct access specialists, mainland France.
Period: January 2016-October 2018, at the exclusion of April-May 2017.
Treatment group: GPs in regulated sector 1.
Dependent variable: Number of visits per GP and per month.
Linear model (columns 1 and 3): dependent variable in logarithm.
PPML (columns 2 and 4): includes 0 activity at physician-month level, pseudo $R^{2}$ reported.
Clustered standard errors at the physician level.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
We perform a difference-in-differences exercise in which we consider GPs in the regulated sector (so-called sector 1) as a treated group. We consider two control groups that were unaffected by the reform: direct access specialists (gynecologists, ophthalmologists, and stomatologists [^13] In section 6.2, we exclude gynecologists because we seek to assess how gender homophily changed after the fee increase, and because male patients do not visit gynecologists. Patients aged less than 18 are also excluded from the analysis conducted in section 6.2.), on the one hand, and GPs in the unregulated sector (so-called sector 2), on the other hand. The outcome $Y_{jt}$ is the number of medical visits by doctor $j$ on month $t$ and we assume that it follows a Poisson model: $Y_{jt}\sim\mathcal{P}(\lambda_{jt})$ with
$\ln\lambda_{jt}=\beta\text{Post}_{t}\times T_{j}+\alpha_{j}+\gamma_{t}+\varepsilon_{jt},$ (D.1)
where Post is a dummy equal to 1 after May 2017, $T_{j}$ is a binary variable that accounts for the treatment group status, and $(\alpha,\gamma)$ are physician and month-year fixed effects. $\varepsilon_{jt}$ is an idiosyncratic shock, and standard errors are clustered at the physician level. We also estimate a linear version of the model in which the dependent variable is $\ln Y_{jt}$, which mechanically eliminates doctor-month observations with a zero activity.
Results reported in Table D.1 show that this reform led physicians to increase their activity by between 9 and 14%, depending on the specification and the control group considered. wilner2025physician further show that the effect is more pronounced for young doctors, and they also find suggestive evidence that access to healthcare improved in the sense that doctors actually admitted new patients following the fee increase.